### v2

# Bilinear forms with trace functions

Abstract.
We obtain non-trivial bounds for bilinear sums of trace functions below the Pólya-Vinogradov range assuming only that the geometric monodromy group of the underlying $\ell$-adic sheaf satisfies certain simple structural properties, in contrast to previous works which handled only special cases of Kloosterman and hypergeometric sheaves.
Our approach builds on a general “soft” stratification theorem for sums of products of trace functions, based on an idea of Junyan Xu, combined with a new robust version of the Goursat–Kolchin–Ribet criterion.
Dedicated to Nick Katz
“Too old to break and too young to tame”
1. Introduction
1.1. Bilinear sums of trace functions
A central part of analytic number theory deals with the issue of estimating the “correlation”
$\sum_{m}f(m)\overline{g(m)}$
between arithmetic sequences of various types. In many important instances, this difficult problem can be transformed into a slightly more tractable question of bounding non-trivially some bilinear forms of the type
$\sum_{m,n}\alpha_{m}\beta_{n}k(m,n),$
where the “kernel” $k$ is supposed to be relatively well-understood, whereas not much is assumed to be known of the sequences $(\alpha_{m})$ and $(\beta_{n})$, except for their average size for instance. We recall that the point of the bilinear form is that, using only such basic knowledge of the sequences $(\alpha_{m})$ and $(\beta_{n})$, one can obtain highly non-trivial results due to the oscillations of $k(m,n)$.
The prototypical example of this transformation is found in the decompositions of the von Mangoldt function in terms of bilinear (or multilinear) expressions, first discovered by Vinogradov. In this case, as in a number of others, the kernel is of the form $k(m,n)=K(mn)$ for a suitable function $K$. Moreover, for non-zero integers $b$ and $c$, the “monomial” kernels $k_{b,c}(m,n)=K(m^{b}n^{c})$ are also of particular interest, and in fact appear naturally in our motivating application.
In these, as explained in Section 1.2, we consider the further restriction that the function $K$ is a trace function modulo a prime number $q$, and we seek estimates when $m$, $n$ run over intervals
$1\leqslant M\leqslant m<2M\leqslant q,\quad\quad 1\leqslant N\leqslant n<2N\leqslant q,$
with the key goal (motivated by our applications, and many others) of obtaining non-trivial bounds for $M$ and $N$ as small as possible compared to $q$.
This is the same context as in various previous papers (see [FI] by Friedlander and Iwaniec, [FMAnn] by Fouvry and Michel and [KMSAnn], [Pisa] by Kowalski, Michel and Sawin). The key achievement of the present work is to obtain estimates as good as those of these papers, but in much greater generality. More precisely, whereas it was previously always assumed that $K$ was of a specific type (essentially some exponential sum of specific type, such as hyper-Kloosterman sums), we are now able to succeed with qualitative conditions on $K$, and these conditions are, on the one hand, known to hold for a very wide variety of trace functions and, on the other hand, extremely robust under a variety of transformations.
In order to state a simplified version of the main result, we recall that a trace function $K$ modulo $q$ is associated to a certain algebraic object. We assume that $K$ is the trace function of a constructible $\ell$-adic sheaf $\mathscr{F}$ on the affine line over the finite field $\mathbf{F}_{q}$, where $\ell$ is a prime $\not=q$, which is a middle-extension sheaf pure of weight 0. We moreover identify $\overline{\mathbf{Q}}_{\ell}$ with $\mathbf{C}$, so we can view $K$ as a complex-valued function. There are intrinsic “symmetry groups” attached to $\mathscr{F}$, called the geometric and arithmetic monodromy groups, which have been explicitly computed by Katz in many instances; we denote by $G$ the geometric monodromy group, which can be viewed as a closed subgroup of $\mathbf{GL}_{r}(\mathbf{C})$ for some non-negative integer $r$. Moreover, there is a numerical complexity invariant $c(\mathscr{F})$ of $\mathscr{F}$, which is a non-negative integer. A simplified version of our main result is:
Theorem 1.1.
Assume that $G$ acts irreducibly on $\mathbf{C}^{r}$ and that one of the following conditions holds:
  1. (1)
    the connected component of the identity of $G$ is a simple algebraic group;
  2. (2)
    or the group $G$ is finite and quasisimple.111 Recall that this means that $G$ is equal to its commutator subgroup and the quotient of $G$ by its center is a simple group, which is then necessarily non-abelian; examples include the alternating group $A_{n}$ for $n\geqslant 5$ and $\mathbf{SL}_{n}(\mathbf{F}_{p})$ for $n\geqslant 3$ and $p\geqslant 3$ prime.
Let $b$, $c$ be non-zero integers. Let $\delta>0$ be a real number and let $1\leqslant M,N\leqslant q/2$ be integers such that
$q^{\delta}\leqslant M,\quad\quad MN\geqslant q^{3/4+\delta}.$
There exists $\eta>0$, depending only on $\delta$, such for that for any sequences of complex numbers222 We recall that the notation $n\sim N$ means that $N\leqslant n<2N$, and similarly for $m\sim M$, etc.
$(\alpha_{m})_{m\sim M},\quad(\beta_{n})_{n\sim N}$
we have
$\sum_{m\sim M}\sum_{n\sim N}\alpha_{m}\beta_{n}K(m^{b}n^{c})\ll\Bigl(\sum_{m}|\alpha_{m}|^{2}\Bigr)^{1/2}\Bigl(\sum_{n}|\beta_{n}|^{2}\Bigr)^{1/2}(MN)^{1/2-\eta},$
where $\eta$ and the implicit constant depend only on $b$, $c$, $\delta$, and on the complexity of $\mathscr{F}$.
Example 1.1.
The theorem applies for instance if the connected component of the geometric monodromy group of $\mathscr{F}$ is either the special linear group $\mathbf{SL}_{r}$ or the symplectic group $\mathbf{Sp}_{r}$ for some integer $r\geqslant 2$. In particular, this means that Theorem 1.1 recovers the main results of [Pisa] (to see this, combine Theorem 1.2, Definition 2.1 and Theorem 6.2 of loc. cit., together with the fact that the complexity of the relevant sheaves are bounded in terms of $r$ only).
We refer to Section 1.2 below for the detailed statements of our most general results, and to Section 9 for a wide variety of concrete examples of trace functions where the conditions of this result (or of the more general variant in Definition 1.2) are satisfied. In Section 1.4, we present our motivating application.
Although the strategy we follow remains similar to that of previous works, there are three key new ingredients which are essential to obtain a bound in the generality in which we work:
  • the general Quantitative Sheaf Theory of Sawin (as expounded by Forey, Fresán, Kowalski and Sawin [qst]) allows us to perform complicated algebraic transformations on the sheaves we work with while keeping control of their complexity, whereas one previously had to do this in an ad-hoc way, which was essentially only possible given a “formula” for the trace function (see, e.g., [KMSAnn]*Prop. 4.24 for an instance of this);
  • we establish new flexible and robust variants of the Goursat–Kolchin–Ribet framework of Katz [ESDE, § 1.8], which allow us in practice to obtain cancellation in “sums of products” situation (in the sense of [sumproducts]) under qualitative conditions on the monodromy (such as those in Theorem 1.1), and not simply for specific groups (see Section 3, and in particular Corollary 3.14, which is an estimate in the spirit of [sumproducts] with independent interest). One particular interesting point is that these new results allow us to treat instances of sheaves with finite monodromy groups (see Section 9.8 for examples);
  • we make essential use of a beautiful idea of Xu [Xu], which roughly speaking states that bounds for the moments of a trace function imply stratification results, from which one can deduce further analytic properties which turn out to be significantly stronger than those which could be obtained directly from the moment bounds (see Theorem 2.4). Note that our version of this idea relies also in an essential way on Quantitative Sheaf Theory.
Remark 1.2.
Our method is in one respect less efficient than that of [Pisa]: whereas the latter provides estimates where the dependency on the complexity (when it applies) is polynomial, this is not the case in this paper, due to the use of qualitative results in Quantitative Sheaf Theory.
Remark 1.3.
The possibility of applying the basic analytic idea (the +uv+uv shift) to monomial kernels $K(m^{a}n^{b})$ first appeared in the work of Nunes [nunesarx], who studied the distribution of squarefree numbers in large arithmetic progressions (here the relevant monomial was $mn^{2}$); we refer also to [nunes] for a stronger result of Nunes using a result of Pierce.
1.2. General statements
We will state precisely our most general results in this section. This requires more terminology and background material in algebraic geometry; Section 2 gives precise definitions and references.
We first define formally the class of sheaves for which our results apply. We recall again that a finite group $G$ is called quasisimple if it is perfect and if the quotient of $G$ by its center $Z(G)$ is simple (in which case it is elementary that $G/Z(G)$ is non-abelian; see, e.g., [isaacs, Lemma 9.1]).
Definition 1.2 (Gallant groups and sheaves).
(1) Let $E$ be an algebraically closed field of characteristic 0, let $r\geqslant 0$ be an integer and let $G\subset\mathbf{GL}_{r,E}$ be a linear algebraic subgroup of $\mathbf{GL}_{r}$ over $E$. The group $G$ is said to be gallant if the action of $G$ on $E^{r}$ is irreducible and if moreover one of the following conditions is satisfied:
  1. (1)
    the identity component $G^{0}$ of $G$ is a simple algebraic group (in particular, the integer $r$ is at least $2$);
  2. (2)
    or the group $G$ is finite and contains a quasisimple normal subgroup $N$ acting irreducibly on $E^{r}$.
In the first case, we define $N=G^{0}$, and in all cases we say that $N$ is the core subgroup of $G$.
(2) Let $k$ be a finite field and $\ell$ a prime number invertible in $k$. Let $\mathscr{F}$ be an $\ell$-adic sheaf on $\mathbf{A}^{1}_{k}$ with generic rank $r\geqslant 0$. We say that $\mathscr{F}$ is gallant if the geometric monodromy group of $\mathscr{F}$ is gallant as a subgroup of $\mathbf{GL}_{r}(\overline{\mathbf{Q}}_{\ell})$.
We say that $\mathscr{F}$ is gallant and light if $\mathscr{F}$ is gallant and moreover is mixed of integral weights $\leqslant 0$ and its restriction to some open dense subset is pure of weight 0.
We note that a gallant sheaf is geometrically irreducible by definition. Moreover, again by definition, the sheaves appearing in Theorem 1.1 are gallant.
In addition to general bilinear forms (classically called “type II” sums), we will consider special bilinear forms (called “type I”) of the form
$\sum_{m}\sum_{n}\alpha_{m}K(m^{b}n^{c}),$
for which one can usually prove slightly stronger results with slightly easier proofs.
Our main estimates are stated in the following theorem:
Theorem 1.3 (Type I and II estimates).
Let $b$, $c$ be non-zero integers. Let $q$ be a prime and let $\mathscr{F}$ be a gallant and light $\ell$-adic sheaf over $\mathbf{F}_{q}$ for some prime $\ell\not=q$.
Let $K$ be the trace function of $\mathscr{F}$. Let $l\geqslant 2$ be an integer and let $M$, $N\geqslant 1$ be integers.
  1. (1)
    Suppose that
    $M\leqslant q,\quad\quad 10q^{1/l}\leqslant N\leqslant q^{1/2+1/(2l)}.$
    For any $\varepsilon>0$ and for any family $\text{${\alpha}$}=(\alpha_{m})_{m\sim M}$ of complex numbers, the estimate
    (1.1) $\sum_{m}\sum_{n}\alpha_{m}K(m^{b}n^{c})\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}M^{1/2}N\Bigl(\frac{q^{1+3/(2l)}}{MN^{2}}\Bigr)^{1/(2l)}$
    holds, where the implicit constant depends at most on $\varepsilon$, $b$, $c$, $l$ and on the complexity of $\mathscr{F}$.
  2. (2)
    Suppose that
    (1.2) $M\leqslant q,\quad\quad 10q^{3/(2l)}\leqslant N\leqslant q^{1/2+3/(4l)}.$
    For any $\varepsilon>0$ and for any families $\text{${\alpha}$}=(\alpha_{m})_{m\sim M}$ and $\text{${\beta}$}=(\beta_{n})_{n\sim N}$ of complex numbers, the estimate
    (1.3) $\sum_{m}\sum_{n}\alpha_{m}\beta_{n}K(m^{b}n^{c})\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}(MN)^{1/2}\Bigl(\frac{1}{M}+\Bigl(\frac{q^{\tfrac{3}{4}+\tfrac{7}{4l}}}{MN}\Bigr)^{1/l}\Bigr)^{1/2}$
    holds, where the implicit constant depends at most on $\varepsilon$, $b$, $c$, $l$ and on the complexity of $\mathscr{F}$.
Remark 1.4.
(1) The “trivial” bound for (general) type I sums is of the form
$\Bigl|\sum_{m}\sum_{n}\alpha_{m}k(m,n)\Bigr|\leqslant\|\text{${\alpha}$}\|_{2}\,M^{1/2}N\,\|k\|_{\infty}.$
Taking $l$ large enough, we see that the bound (1.1) is non-trivial, in the sense of providing an estimate of the form
$\sum_{m}\sum_{n}\alpha_{m}K(m^{b}n^{c})\ll\|\text{${\alpha}$}\|_{2}(M^{1/2}N)^{1-\eta}$
for some $\eta>0$, as long as there exists $\delta>0$ such that $M^{2}N\geqslant q^{1+\delta}$, in which case $\eta$ depends on $\delta$. In particular, we obtain a non-trivial estimate in the case
$M=N=q^{\frac{1}{3}(1+\delta)}$
for any $\delta>0$.
(2) For type II sums, the trivial bound is of the form
$\Bigl|\sum_{m}\sum_{n}\alpha_{m}\beta_{n}k(m,n)\Bigr|\leqslant\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}\,(MN)^{1/2}\,\|k\|_{\infty}.$
Here, taking $l$ large enough, the bound (1.3) gives an estimate of the form
$\sum_{m}\sum_{n}\alpha_{m}\beta_{n}K(m^{b}n^{c})\ll\|\alpha\|_{2}\|\beta\|_{2}(MN)^{1/2-\eta}$
for some $\eta>0$ as soon as
$q^{\delta}\leqslant M\leqslant q,\ MN\geqslant q^{3/4+\delta}$
for some $\delta>0$, with $\eta$ depending on $\delta$. It follows that Theorem 1.3 implies Theorem 1.1, and also that we have a non-trivial estimate in the special case
$M=N=q^{\frac{3}{8}(1+\delta)}$
for $\delta>0$.
(3) Our main target in this paper is to have non-trivial bounds in the largest possible range of uniformity of $M$ and $N$ relative to $q$. There is no doubt that a more sophisticated treatment could be used to improve the actual saving compared to the trivial bound, in some ranges of $M$ and $N$.
Again with some motivations in mind, we will furthermore prove a “trilinear” bound.
Theorem 1.4.
Let $a$, $b$, $c$ be non-zero integers. Let $q$ be a prime and let $\mathscr{F}$ be a gallant and light $\ell$-adic sheaf over $\mathbf{F}_{q}$ for some prime $\ell\not=q$. Let $K$ be the trace function of $\mathscr{F}$.
Let $l\geqslant 2$ be an integer and let $J$, $M$, $N\geqslant 1$ be integers. Suppose that
$J\leqslant 4q,\quad\quad MN\leqslant 4q.$
For any $\varepsilon>0$ and for any families
$\text{${\alpha}$}=(\alpha_{j})_{j\sim J},\quad\text{${\beta}$}=(\beta_{m})_{m\sim M},\quad\text{${\gamma}$}=(\gamma_{n})_{n\sim N}$
of complex numbers of modulus $\leqslant 1$, the estimate
(1.4) $\mathop{\sum\sum\sum}\limits_{j\sim J,m\sim M,n\sim N}\alpha_{j}\beta_{m}\gamma_{n}K(j^{a}m^{b}n^{c})\ll q^{\varepsilon}JMN\Bigl(\frac{q^{1/2}}{MN}+\frac{q}{J^{l}MN}\Bigr)^{1/(2l)}$
holds, where the implicit constant depends on $\varepsilon$, $a$, $b$, $c$, $l$ and on the complexity of $\mathscr{F}$.
Remark 1.5.
The bound (1.4) is non-trivial as long as
$J\geqslant q^{\delta},\quad MN\geqslant q^{1/2+\delta},$
for some $\delta>0$, taking $l$ sufficiently large so that $J^{l}MN\geqslant q^{1+\delta}$.
Our methods enable us to handle other kinds of multilinear sums of gallant and light trace functions with summation variables varying over ranges that may be quite short. For instance, following Xi and Zheng [XZ], one can show that, as long as
$q^{1/6+\delta}\leqslant M\leqslant q^{1/2-\delta},\quad MN^{2}\geqslant q^{1/2+\delta}$
for some $\delta>0$, the “quadrilinear” sum
$\mathop{\sum\sum}\limits_{m_{1},m_{2}\sim M}\alpha_{m_{1}}\alpha_{m_{2}}\mathop{\sum\sum}\limits_{n_{1},n_{2}\sim N}K(\overline{m}_{1}n_{1}-\overline{m}_{2}n_{2}),$
is bounded by $\ll\|\text{${\alpha}$}\|^{2}MN^{2}(MN)^{-\eta}$ for some $\eta>0$, depending on $\delta$. This bound is non-trivial for
$M=N=q^{1/6+\delta}$
and $\delta\in\mathopen{]}0,1/3\mathclose{[}$.
The proof follows [XZ] together with the following useful special case of Proposition 6.2.
Proposition 1.5.
Let $q$ be a prime number and let $\mathscr{F}$ be a gallant and light sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$.
For $l\geqslant 2$ an even integer, $r,s\in\mathbf{F}_{q}$ and $\text{${v}$}\in\mathbf{F}_{q}^{2l}$, let
$\mathbf{K}(r,s,\text{${v}$})=\prod_{i=1}^{l}K(s(r+v_{i}))\overline{K(s(r+v_{i+l}))}.$
and
$\Sigma_{I}(\text{${v}$})=\sum_{(r,s)\in\mathbf{F}_{q}\times\mathbf{F}^{\times}_{q}}\mathbf{K}(r,s,\text{${v}$}).$
There exists an integer $C\geqslant 1$, depending only on $l$ and on the complexity of $\mathscr{F}$ such that, if $q\geqslant C$, there exist algebraic varieties $\mathscr{V}_{1}$ and $\mathscr{V}_{1}^{\Delta}$ over $\mathbf{F}_{q}$ such that
$\mathbf{A}^{2l}_{\mathbf{F}_{q}}\supset\mathscr{V}_{1}\supset\mathscr{V}_{1}^{\Delta}$
$\dim(\mathscr{V}_{1})\leqslant\frac{3}{2}l,\quad\dim(\mathscr{V}^{\Delta}_{1})\leqslant l,\ \max(\deg(\mathscr{V}_{1}),\deg(\mathscr{V}_{1}^{\Delta}))\leqslant C$
and
$|\Sigma_{I}(\text{${v}$})|\leqslant\begin{cases}Cq&\text{ for }\text{${v}$}\not\in\mathscr{V}_{1}(\mathbf{F}_{q}),\\ Cq^{3/2}&\text{ for }\text{${v}$}\not\in\mathscr{V}^{\Delta}_{1}(\mathbf{F}_{q}),\\ Cq^{2}&\text{ for }\text{${v}$}\in\mathscr{V}_{1}^{\Delta}(\mathbf{F}_{q}).\end{cases}$
Combining this with the counting Lemma 2.1, we deduce that for any $V\in[0,q/2\mathclose{[}$, the estimate
$\frac{1}{V^{2l}}\sum_{\text{${v}$}\in[V,2V]^{2l}}\Bigl|\sum_{(r,s)\in\mathbf{F}_{q}\times\mathbf{F}^{\times}_{q}}\mathbf{K}(r,s,\text{${v}$})\Bigr|\ll q\Bigl(1+\frac{q}{V^{l}}\Bigr),$
holds, where the implicit constant depends only on $l$ and on the complexity of $\mathscr{F}$ (this is a special case of Proposition 4.4). This proves that, for any $V\geqslant q^{1/l}$, the sum $\Sigma_{I}(\text{${v}$})$ achieves squareroot cancellation on average over $\text{${v}$}\in[V,2V]^{2l}$.
As another example, one can check, using Proposition 1.5 or suitable variants, that the results of [Xi] extend to trace functions of gallant and light sheaves.
1.3. An exceptional case
A sheaf $\mathscr{F}$ with geometric monodromy group isomorphic to either $\mathbf{O}_{4}$ or $\mathbf{SO}_{4}$ will not be gallant because the algebraic group $\mathbf{SO}_{4}$ is not simple (it is isomorphic to the quotient of $\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ by the diagonally embedded subgroup $\{-1,1\}$). We will say that $\mathscr{F}$ is sulfatic if the geometric monodromy group of $\mathscr{F}$ is isomorphic to $\mathbf{SO}_{4}$ and oxozonic if it is isomorphic to $\mathbf{O}_{4}$ and has the property that the subgroup $\mathbf{SO}_{4}$ acts irreducibly on the underlying representation.
It turns out that a slight extension of our method allows us to deal with the oxozonic case, which is relevant to the application described in the next section
Theorem 1.6.
Let $b$, $c$ be non-zero integers. Let $q$ be a prime and let $\mathscr{F}$ be an $\ell$-adic sheaf over $\mathbf{F}_{q}$ for some prime $\ell\not=q$ which is mixed of weights $\leqslant 0$. We assume that $\mathscr{F}$ is oxozonic.
  1. (1)
    Theorem 1.3 holds for $\mathscr{F}$ if $c$ is odd, and in particular Theorem 1.1 also holds for $\mathscr{F}$ in this case.
  2. (2)
    Theorem 1.4 holds for $\mathscr{F}$.
1.4. Cubic toroidal moments
The original motivation which led to Theorem 1.1 (and required its general form) was the study of what we call cubic toroidal moments333 Also called cubic mixed moments. of special values of Dirichlet $L$-functions at the central point, generalizing the quadratic case considered by Fouvry, Kowalski and Michel in [FKMAA]; we refer to the introduction of this paper for a general discussion.
In the present context, we consider a prime number $q$ and non-zero integers $a$, $b$, $c$, and we define
$M_{a,b,c}(q)=\frac{1}{q-1}\sum_{\chi\,(\mathrm{mod}\,{q})}L(1/2,\chi^{a})L(1/2,\chi^{b})L(1/2,\chi^{c}).$
We wish to evaluate this moment asymptotically as $q\rightarrow+\infty$. After permuting $(a,b,c)$ and applying complex conjugation with $L(1/2,\overline{\chi})=\overline{L(1/2,\chi)}$, we see that is enough to consider the moments $M_{a,b,c}(q)$ and $M_{a,b,-c}(q)$ when $a$, $b$ and $c$ are positive integers, which we will always assume from now on in this section.
It is also useful to decompose $M_{a,b,c}(q)$ along the even and odd characters, so we write
$M_{a,b,c}(q)=M^{+1}_{a,b,c}(q)+M^{-1}_{a,b,c}(q)$
where
$M^{\varepsilon}_{a,b,c}(q)=\frac{1}{q-1}\sum_{\stackrel{{\scriptstyle{\scriptstyle\chi\,(\mathrm{mod}\,{q})}}}{{{\scriptstyle\chi(-1)=\varepsilon}}}}L(1/2,\chi^{a})L(1/2,\chi^{b})L(1/2,\chi^{c}),\quad\quad\varepsilon=+1,\ -1.$
Define exponential sums modulo $q$ by
$K_{a,b,c}(u;q)=\frac{1}{q}\sum_{\begin{subarray}{c}x,y,z\in\mathbf{F}_{q}\\ x^{a}y^{b}z^{c}=u\end{subarray}}e\Bigl(\frac{x+y+z}{q}\Bigr),\quad\quad u\in\mathbf{F}_{q}^{\times}.$
One can show that there exists a natural lisse $\ell$-adic sheaf $\mathscr{K}_{a,b,c}$ on the multiplicative group with trace function $u\mapsto K_{a,b,c}(u)$, whose complexity is bounded in terms of $(a,b,c)$ only. This sheaf is pure of weight 0. In the companion paper [FKMSmoment], we will prove the following result.
Theorem 1.7.
Let $a$, $b$, $c\geqslant 1$ be setwise coprime integers. Let $\varepsilon\in\{-1,1\}$. Assume that $\mathscr{K}_{a,b,\varepsilon c}$ is either gallant or oxozonic in the sense of Definition 1.2 and of Section 1.3.
There exists $\eta>0$ and there exist polynomials $P_{a,b,c}=1$ and $P_{a,b,-c}$ such that
$\deg(P_{a,b,-c})$ $=\delta_{a,c}=\begin{cases}1&\text{ if }a=c,\\ 0&\text{ if }a\not=c,\end{cases}$
$P_{a,b,-c}$ $\text{ has positive leading term},$
such that the estimates
$M^{+1}_{a,b,\varepsilon c}(q)$ $\geqslant P_{a,b,\varepsilon c}(\log q)+O(q^{-\eta}),$
$M_{a,a,\varepsilon c}(q)$ $=P_{a,a,\varepsilon c}(\log q)+O(q^{-\eta}),$
hold for all primes $q$, where the implied constant depends only on $(a,b,c)$.
In order for this result to be applicable, we will also prove that almost all sheaves $\mathscr{K}_{a,b,c}$ are either gallant or oxozonic, using the work of Katz.
Theorem 1.8.
Let $a$, $b$, $c\geqslant 1$ be globally coprime integers.
  1. (1)
    The sheaf $\mathscr{K}_{a,b,c}$ is gallant unless $a+b+c=4$, and is oxozonic otherwise.
  2. (2)
    The sheaf $\mathscr{K}_{a,b,-c}$ is gallant unless one of the following holds:
    • $a=c$ or $b=c$, so that (a,b,c)=(c,b,c)(a,b,-c)=(c,b,-c) or (a,c,c)(a,c,-c);
    • $a+b=c$ and $c\in\{2,3,4\}$;
    • Otherwise, $a+b+c\equiv 0\,(\mathrm{mod}\,{2})$ and
      $\max(a+b,c)-(a,c)-(b,c)+1=4.$
      In that last case, up to switching $a$ and $b$, the triple $(a,b,c)$ is one of the four triples $(1,2,5)$, $(1,4,3)$, $(1,6,3)$ or $(2,3,1)$, and the sheaf $\mathscr{K}_{a,b,-c}$ is then oxozonic with the exception of $\mathscr{K}_{1,2,-5}$, which is sulfatic.
As a corollary we obtain a non-vanishing result:
Corollary 1.9.
Let $a$, $b$, $c$ be three non-zero setwise coprime integers. For $q\geqslant q_{0}$ large enough, the lower bound
$|\{\chi\,(\mathrm{mod}\,{q})\,\mid\,L(1/2,\chi^{a})L(1/2,\chi^{b})L(1/2,\chi^{c})\not=0\}|\gg q/\log^{12}q$
holds, with an implied constant depending only on $a$, $b$, $c$ and $q_{0}$.
In a similar vein, Berta and zur Verth [BV], using the results of this paper, obtained non-vanishing results for products of two Dirichlet $L$-values under angular constraints on the associated Gauss sums (compare with [MAMS]*Thm. 1.8):
Theorem 1.10 (Berta–zur Verth).
[…]
$|\{\chi\,(\mathrm{mod}\,{q}),\ \chi\not=1,\ L(1/2,\chi^{a})L(1/2,\chi^{b}),\ \varepsilon(\chi)\in I\}|\geqslant c(q-2),$
holds, where $\varepsilon(\chi)$ denotes the unitarily normalized Gauss sum of $\chi$.
1.5. Outline of the paper
  • In Section 2, we discuss the relevant facts from algebraic geometry. Of particular importance is our general form of Xu’s idea (see Theorem 2.4).
  • In Section 3, we present our generalized form of Goursat’s Lemma and the Goursat–Kolchin–Ribet criterion, as well as applications.
  • In Section 4, we present in a general abstract form the basic implication of the “shift by +uv+uv” method to study bilinear forms, reducing statements like Theorem 1.3 to estimates for certain “complete” sums of the kernel. This part does not require $K$ to be a trace function and we express it in the form of Proposition 4.1. We also explain at the end of this section how this proposition will be applied for trace functions.
  • Sections 5 and 6 provide the core of the proof: combining Xu’s idea with other tools, we show how to obtain good estimates for the auxiliary complete sums arising from Proposition 4.1.
  • In Section 7, we give the proof of Theorem 1.4.
  • In Section 8, we handle the case of oxozonic sheaves.
  • Finally, Section 9 provides a series of examples of gallant sheaves, relying especially on the work of Katz (see [GKM] and [ESDE] in particular). We include examples with finite monodromy group, especially those arising from hypergeometric sheaves, from the tables of Beukers and Heckman [BH]. This section can be read independently of the proof of the main results, and is meant to show that variety of situations where the results apply.
Acknowledgements
This research was partially supported by the SNF grants $197045$, $219220$, the SNF-ANR “Etiene” grant $10003145$, the NSF grant DMS-2502029, and a Sloan Research Fellowship.
We are grateful to Filippo Berta, Peter Sarnak, Svenja zur Verth and Ping Xi for encouragements and useful comments.
EF thanks EPFL and ETHZ for their hospitality and for providing an environment conductive to a fruitful collaboration.
Notation
For $z\in\mathbf{C}$, we denote $e(z)=e^{2i\pi z}$.
We use the notation $f\ll g$ and $f=O(g)$, for $f$ and $g$ defined on a set $X$, synonymously: either notation means that there exists a constant $C\geqslant 0$ such that $|f(x)|\leqslant Cg(x)$ for all $x\in X$.
We will often have statements that a certain function $f$ is “bounded polynomially” in terms of certain quantities, say $\alpha$, $\beta$; this means that there exists real numbers $c\geqslant 0$ and $A\geqslant 0$ such that for all $x$ in the domain of definition of $f$ and $g$, the inequality
$|f(x)|\leqslant c(\alpha\beta)^{A}$
holds. This will apply in particular to certain implicit constants in bounds $f\ll g$; we then mean that
$|f(x)|\leqslant c(\alpha\beta)^{A}g(x)$
for suitable $c$ and $A$ as above.
2. Algebro-geometric preliminaries
We recall here some facts from algebraic geometry and establish our notation and conventions.
Let $k$ be a field. An algebraic variety $X$ over $k$ is a reduced and separated scheme of finite type over the spectrum of $k$.
By the degree of a subvariety $X$ of projective space, we mean the sum of the degrees of its irreducible components (which may have different dimensions).
For a prime number $\ell$ different from the characteristic of $k$, we denote by ${\mathrm{D}_{c}^{\mathrm{b}}}(X,\overline{\mathbf{Q}}_{\ell})$ the bounded derived category of complexes of $\ell$-adic constructible sheaves on $X$. An object of this category will be called simply an $\ell$-adic complex; by $\ell$-adic sheaf on $X$, we will mean a constructible $\ell$-adic sheaf. We will sometimes just speak of complex or sheaf, if no confusion concerning $\ell$ can arise.
Whenever we consider a field $k$ and prime $\ell$ invertible in $k$, we also fix an isomorphism $\iota\colon\overline{\mathbf{Q}}_{\ell}\to\mathbf{C}$, and we use it implicitly to identify $\ell$-adic numbers with complex numbers. In particular, if $k$ is a finite field, we can speak of the value in $\mathbf{C}$ of the trace function of a complex $M$ on $X$ at a point $x\in X(k)$. We sometimes denote this value by $t_{M}(x;k)$, and correspondingly write $t_{M}(x;k_{n})$ if $k_{n}/k$ is a finite extension of $k$ of degree $n$ and $x\in X(k_{n})$. When we speak of mixed or pure complexes of $\ell$-adic sheaves, we also mean $\iota$-mixed or $\iota$-pure.
If $X$ is a smooth and geometrically connected curve over kk,444 Which, in this paper, will most often be the affine line. a middle-extension sheaf is an $\ell$-adic sheaf $\mathscr{F}$ on $X$ such that for some (equivalently, for any) open dense subset $U\subset X$ with open immersion $j\colon U\to X$, the natural adjunction morphism $\mathscr{F}\to j_{*}j^{*}\mathscr{F}$ is an isomorphism (see [ESDE, (7.3.1)]).
Under the same assumption on $X$, let $\mathscr{F}$ be an $\ell$-adic sheaf. By definition of constructible sheaves, there exists a dense open subset of $X$, say $U$, such that the restriction of $\mathscr{F}$ to $U$ is lisse. Thus $\mathscr{F}|U$ corresponds to a continuous $\ell$-adic representation of the étale fundamental group of $U$, well-defined up to isomorphism. The arithmetic monodromy group of $\mathscr{F}|U$ is classically defined to be the Zariski-closure of the image of this representation, viewed as a $\overline{\mathbf{Q}}_{\ell}$-linear algebraic group. Up to isomorphism, this group is independent of the choice of $U$, and we call it the arithmetic monodromy group of $\mathscr{F}$. Similarly, let $\bar{k}$ be an algebraic closure of $k$; the restriction of the representation to the fundamental group of $U_{\bar{k}}$ defines the geometric monodromy group of $\mathscr{F}|U_{\bar{k}}$. It is independent of $U$, up to isomorphism, and called the geometric monodromy group of $\mathscr{F}$.
The étale fundamental group of $U$ is a quotient of the Galois group $\operatorname{Gal}(\overline{k(X)}/k(X))$ of the function field of $X$, and we will sometimes identify the restriction of $\mathscr{F}$ to $U$ with a finite-dimensional $\ell$-adic representation of this Galois group.
We will also often use the fact that the arithmetic (resp. geometric) monodromy group of $\mathscr{F}$ coincides with that of the sheaf $j_{!}j^{*}\mathscr{F}$. This allows us to assume in many proofs that $\mathscr{F}$ is the extension by zero of a lisse sheaf on some dense open subset of $X$.
The dual of $\mathscr{F}$ is defined to be the sheaf $\mathscr{F}^{\vee}=\underline{\mathrm{Hom}}(\mathscr{F},\overline{\mathbf{Q}}_{\ell})$; it is lisse on $U$ and coincides there with the lisse sheaf associated to the contragredient of the representation corresponding to $\mathscr{F}$. In particular, if $\mathscr{F}$ is pure of weight 0 on $U$, then the trace function of $\mathscr{F}^{\vee}$ on $U$ is (under $\iota$) the complex conjugate of the trace function of $\mathscr{F}$. This property is also true on all of $X$ if  $\mathscr{F}$ is a middle-extension sheaf (by a theorem of Gabber [MMP, proof of Duality Lemma 1.8.1(1)]).
As in previous works, we will use the following standard lemma (see [Xu, Lemma 1.7]).
Lemma 2.1.
Let $l\geqslant 1$ be an integer and let $A>0$. Let $q$ be a prime number and let $X\subset\mathbf{A}^{l}_{\mathbf{F}_{q}}$ be an algebraic variety of dimension $d\geqslant 0$ given by the vanishing of $\leqslant A$ polynomials of degree $\leqslant A$. Let $V$ be an integer with $0\leqslant V<q/2$. Then
$|\{x=(x_{1},\ldots,x_{l})\in\mathbf{Z}^{l}\,\mid\,V\leqslant x_{i}\leqslant 2V\text{ for }1\leqslant i\leqslant l\text{ and }x\in X(\mathbf{F}_{q})\}|\ll V^{d},$
where the implicit constant depends only on $l$ and $A$.
We will use Sawin’s quantitative sheaf theory [qst], and we recall the framework. Let $k$ be a field and $X/k$ a quasi-projective variety defined over $k$, given with a locally-closed embedding $u\colon X\to\mathbf{P}^{N}$ for some integer $N\geqslant 1$. Let $\ell$ be a prime number invertible in $k$. Sawin defined the complexity $c_{u}(M)$ of objects of ${\mathrm{D}_{c}^{\mathrm{b}}}(X,\overline{\mathbf{Q}}_{\ell})$ (see [qst, Def. 6.3]). This is a non-negative integer which controls quantitatively most invariants of $M$.
If $X$ is an open subvariety of $\mathbf{A}^{n}$ for some integer $n\geqslant 1$, we will always use the natural embedding $u\colon X\to\mathbf{P}^{n}$ to compute the complexity of objects of ${\mathrm{D}_{c}^{\mathrm{b}}}(X,\overline{\mathbf{Q}}_{\ell})$, and will use the notation $c(M)$ instead of $c_{u}(M)$. If $n=1$, then for a middle-extension sheaf $\mathscr{F}$ on $\mathbf{A}^{1}_{k}$, viewed as a complex located in some degree, and with $u$ the obvious embedding of $\mathbf{A}^{1}$ in $\mathbf{P}^{1}$, the complexity $c_{u}(\mathscr{F})$ is comparable with the complexity $c_{\mathrm{FKM}}(\mathscr{F})$ defined by Fouvry, Kowalski and Michel [FKM1, Def. 1.13]; more precisely, we have
$c_{\mathrm{FKM}}(\mathscr{F})\leqslant c_{u}(\mathscr{F})\leqslant 3c_{\mathrm{FKM}}(\mathscr{F})^{2}$
by [qst, Cor. 7.4].
Among the properties of the complexity that we will use, we only quote here a version of Deligne’s Riemann Hypothesis.
Theorem 2.2.
Let $k$ be a finite field. Let $\mathscr{F}$ be an $\ell$-adic sheaf on $\mathbf{A}^{1}_{k}$ which is mixed of weights $\leqslant 0$. Assume that
$H^{2}_{c}(\mathbf{A}^{1}_{\bar{k}},\mathscr{F})=0$
or equivalently that the Galois representation associated to $\mathscr{F}$ has no coinvariants with respect to the geometric monodromy group. We then have
$\sum_{x\in k}t_{\mathscr{F}}(x;k)\ll c_{u}(\mathscr{F})|k|^{1/2}$
where the implied constant is absolute and the embedding $u$ is the obvious one.
Proof.
[…]
$\sum_{x\in U(k)}t_{\mathscr{F}_{i}}(x;k)\ll c_{u}(\mathscr{F}_{i})k|^{1/2}$
by loc. cit, and the conclusion then follows after summing over $i$ from [qst, Th. 6.15]. ∎
For a finite field $k$, an $\ell$-adic additive character $\psi$ of $k$ or an $\ell$-adic multiplicative character $\chi$, we use the usual notation $\mathscr{L}_{\psi}$ or $\mathscr{L}_{\chi}$ for the associated Artin–Schreier or Kummer sheaf, with trace functions $x\mapsto\psi(x)$ or $x\mapsto\chi(x)$.
Proposition 2.3.
Let $k$ be a finite field of characteristic $q$ and $\bar{k}$ an algebraic closure of $k$. Let $\mathscr{{F}}$ be a constructible $\ell$-adic sheaf on $\mathbf{A}^{1}$ over $k$. Assume that the restriction of $\mathscr{F}$ to a dense open subset is a lisse geometrically irreducible sheaf.
Let
$\mathbf{B}_{\mathscr{F}}=\Bigl\{\gamma=\begin{pmatrix}a&b\\ 0&d\end{pmatrix}\in\mathbf{PGL}_{2}(\bar{k})\,\mid\,\gamma^{*}(\mathscr{F})\simeq\mathscr{F}\Bigr\},$
where $\simeq$ denotes geometric isomorphism. This is an algebraic subgroup of $\mathbf{PGL}_{2}(\bar{k})$, and at least one of the following properties holds:
  1. (1)
    The sheaf $\mathscr{F}$ is of generic rank $1$.
  2. (2)
    The group $\mathbf{B}_{\mathscr{{F}}}$ is a finite cyclic group of size bounded polynomially in terms of the complexity of $\mathscr{F}$.
  3. (3)
    The complexity of $\mathscr{F}$ is at least $q^{\delta}$ for some absolute constant $\delta>0$.
Proof.
Let $j\colon U\to\mathbf{A}^{1}$ be the open immersion of an open dense subset where $\mathscr{F}$ is lisse. Define $\mathscr{F}^{*}=j_{*}j^{*}\mathscr{F}$; this is a geometrically simple middle-extension sheaf with complexity bounded polynomially in terms of the complexity of $\mathscr{F}$ (this is clear from the definition of $c_{\mathrm{FKM}}$). We have $\mathbf{B}_{\mathscr{F}}\subset\mathbf{B}_{\mathscr{F}^{*}}$: if $\gamma\in\mathbf{B}_{\mathscr{F}}$, then the existence of a geometric isomorphism $\gamma^{*}(\mathscr{F})\simeq\mathscr{F}$ implies that the middle-extension sheaves $\mathscr{F}^{*}$ and $\gamma^{*}\mathscr{F}^{*}$ are geometrically isomorphic on some dense open set of $\mathbf{A}^{1}$, hence are geometrically isomorphic.
If $\mathscr{F}^{*}$ is an Artin–Schreier sheaf, then the first property of our claim holds. Otherwise, the sheaf $\mathscr{F}^{*}$ is a Fourier sheaf in the sense of [ESDE, Def. 7.3.5] and we can apply [FKM2, Th. 6.3, Prop. 6.4] to the “naïve” (inverse) Fourier transform $\mathscr{G}$ of $\mathscr{F}$ (in the sense of [ESDE, 7.3.3]); the group denoted $\mathbf{B}_{\mathscr{G}}$ in [FKM2] coincides with $\mathbf{B}_{\mathscr{F}^{*}}$ by the Fourier inversion formula [ESDE, 7.3.8].
The first alternative in [FKM2, Th. 6.3] is impossible since it would imply that $\mathscr{F}^{*}$ is an Artin–Schreier, which we assumed is not the case. The second would imply that $\mathscr{G}$ is geometrically isomorphic to $\mathscr{L}_{\psi}\otimes\mathscr{L}_{\chi}$ for some additive character $\psi$ (resp. some multiplicative character $\chi$, necessarily non-trivial since otherwise we would be in the case of an Artin–Schreier sheaf). It is straightforward that $\mathscr{F}^{*}$ would then be of generic rank $1$ (this amounts to the computation
$\sum_{x\in k}\chi(x)\psi(x)\psi(xy)=\tau\overline{\chi}(1+y),$
of trace functions, where $\tau$ is a Gauss sum), thus $\mathscr{F}$ would be of generic rank $1$.
The third possibility is that $\mathbf{B}_{\mathscr{F}^{*}}$ is finite of size bounded by $10c_{\mathrm{FKM}}(\mathscr{G})^{2}$; since the complexity of $\mathscr{G}$ is bounded polynomially in terms of that of $\mathscr{F}$, this gives the second possibility in our statement.
Finally, the fourth possibility is that the complexity $c_{\mathrm{FKM}}(\mathscr{G})$ is $\geqslant(q/10)^{1/2}$; as before this implies the third possibility.
To conclude, we simply note that in the second case, the fact that the group $\mathbf{B}_{\mathscr{F}^{*}}$ is cyclic is not stated in [FKM2], but follows from the proof (see [FKM2, p. 1729]); its subgroup $\mathbf{B}_{\mathscr{F}}$ is then also cyclic. ∎
We now state our version of the result of Xu which we already mentioned. Roughly speaking, it states that bounds for the moments of a trace function imply stratification results, which then imply further analytic results about the trace function which are significantly stronger than those that can be obtained directly from the moment bounds. A concrete form of this idea is the following theorem.
Theorem 2.4.
Let $k$ be a finite field and $\ell$ a prime number different from the characteristic of $k$. Let $X$ be a quasiprojective variety with a locally closed embedding in $\mathbf{P}^{d}_{k}$ for some integer $d\geqslant 0$. Let $M$ be an object of ${\mathrm{D}_{c}^{\mathrm{b}}}(X,\overline{\mathbf{Q}}_{\ell})$ which is mixed of integral weights.
Assume that there exists a positive integer $m$, a real number $A$ and a real number $B\geqslant 1$ such that
$\sum_{x\in X(k_{n})}\left\lvert t_{M}(x;k_{n})\right\rvert^{2m}\leqslant B\left\lvert k_{n}\right\rvert^{A}$
for all integers $n\geqslant 1$, where $k_{n}$ denotes the extension of degree $n$ of $k$ in some algebraic closure of $k$.
There exists a stratification of $X$ into closed subschemes $X^{(w)}$, defined for integers $w$, such that
  1. (1)
    We have $X^{(w)}\subset X^{(w-1)}$ for all $w$.
  2. (2)
    We have $\dim X^{(w)}\leqslant\lfloor A\rfloor-mw$ for all $w$.
  3. (3)
    Each $X^{(w)}$ is a union of subvarieties of total degree bounded only in terms of $n$ and $c(M)$.
  4. (4)
[…]
    $t_{M}(x;k_{n})\ll\left\lvert k_{n}\right\rvert^{\frac{w}{2}},$
    where the implicit constant depends only on $d$ and $c(M)$.
Proof.
It follows from [qst, Lemma 6.26] applied to $M$ that there exist subvarieties
$Y_{d+1}\subset Y_{d}\subset\cdots\subset Y_{1}\subset X$
such that
[…]
  • Each $Y_{i}$ is the union of $\leqslant C$ subvarieties of degree $\leqslant C$, for some integer $C$ depending only on $d$ and the complexity of $M$.
[…]
$t_{M}(x;k)=\sum_{i\in\mathbf{Z}}t_{\mathscr{H}^{2i}(M)}(x;k)-\sum_{i\in\mathbf{Z}}t_{\mathscr{H}^{2i+1}(M)}(x;k),$
the difference of sums of trace functions of pure lisse sheaves. We can then apply a result of Xu [Xu, Theorem 3.5]. This shows that either
(2.1) $\limsup_{n\to\infty}\frac{1}{|k|^{n(\dim(Y)+m(w+1))}}\sum_{y\in Y(k_{n})}\left\lvert t_{M}(y;k_{n})\right\rvert^{2m}\geqslant 1,$
or
$|t_{M}(x;k)|\leqslant D|k|^{w/2},$
with
$D=\sum_{i\in\mathbf{Z}}\operatorname{rank}(\mathscr{H}^{i}(M|Y)).$
This last estimate gives the desired conclusion since $D$ is bounded in terms of $n$ and of the complexity of $M$ (see [qst, Th. 6.8(8)] and [qst, Th. 6.15]; note that to apply the latter, it is important that $M$ is lisse on $Y$). So we need only check that (2.1) is not valid. However, using the assumption we find that
$\sum_{y\in Y(k_{n})}\left\lvert t_{M}(y;k_{n})\right\rvert^{2m}\leqslant\sum_{x\in X(k_{n})}\left\lvert t_{M}(x;k_{n})\right\rvert^{2m}\leqslant B|k|^{nA}$
for all $n\geqslant 1$, hence
$\frac{1}{|k|^{n(\dim(Y)+m(w+1))}}\sum_{y\in Y(k_{n})}\left\lvert t_{M}(y;k_{n})\right\rvert^{2m}\leqslant B|k|^{n(A-\dim(Y)-m(w+1)},$
which tends to 0 as $n\to+\infty$, so this alternative possibility does not occur. ∎
3. Around Goursat’s Lemma
3.1. Goursat’s Lemma
We recall a version of the classical lemma of Goursat.
Lemma 3.1 (Goursat’s Lemma).
Let $G_{1}$ and $G_{2}$ be groups and let $G\subset G_{1}\times G_{2}$ be a subgroup such that the restriction to $G$ of the projections $p_{1}\colon G_{1}\times G_{2}\to G_{1}$ and $p_{2}\colon G_{1}\times G_{2}\to G_{2}$ are surjective.
There exist normal subgroups $H_{i}\subset G_{i}$ and an isomorphism $\varphi\colon G_{1}/H_{1}\to G_{2}/H_{2}$ such that
$H_{1}\times H_{2}\subset G\subset\{(x,y)\in G_{1}\times G_{2}\,\mid\,\varphi(xH_{1})=yH_{2}\}.$
We recall the proof for completeness.
Proof.
The kernel of the restriction of $p_{1}$ (resp. $p_{2}$) to $G$ is identified with a subgroup $H_{2}$ of $G_{2}$ (resp. a subgroup $H_{1}$ of $G_{1}$). One then checks that the image of the natural morphism
$G\to G_{1}/H_{1}\times G_{2}/H_{2}$
is the graph of an isomorphism $\varphi$; the data of $H_{1}$, $H_{2}$ and $\varphi$ has the stated property (e.g. for the inclusion $H_{1}\times H_{2}\subset G$, note that if $(x,y)\in H_{1}\times H_{2}$, then by definition we have $(x,1)\in G$ and $(1,y)\in G$, so that $(x,y)\in G$). ∎
Definition 3.2.
Let $G_{1}$ and $G_{2}$ be groups and let $G\subset G_{1}\times G_{2}$ be a subgroup such that the restriction to $G$ of the projections $G_{1}\times G_{2}\to G_{1}$ and $G_{1}\times G_{2}\to G_{2}$ are surjective. The data $(H_{1},H_{2},\varphi)$ from Goursat’s Lemma is called a Goursat datum for $G$.
3.2. A criterion for vanishing of coinvariants
In applications, the intended conclusion is often that $G=G_{1}\times G_{2}$, which one hopes to achieve using extra assumptions. Generalizations of this criterion exist to handle subgroups of a product of finitely many groups, and were used for instance by Ribet [31] to study $\ell$-adic representations associated to classical modular forms. We are interested in another variant: the groups $G_{i}$ are given with some linear representation, and we wish to understand when the space of (co)invariants of $G$ in the tensor product of the representation spaces may be non-zero.
Proposition 3.3.
Let $k$ be a field. Let $G$ be a group, and let $m\geqslant 1$ be an integer.
For $1\leqslant i\leqslant m$, let $\varrho_{i}\colon G\to\mathbf{GL}(V_{i})$ be a linear representation of $G$ on a finite-dimensional $k$-vector space $V_{i}$. Denote by $G_{i}$ the Zariski-closure of the image of $\varrho_{i}$, and let $N_{i}$ be a normal subgroup of $G_{i}$. Assume that $N_{i}$ is perfect and that the space of $N_{i}$-coinvariants of $V_{i}$ is zero.
One of the following properties holds:
  1. (1)
    The space of $G$-coinvariants of the representation $\varrho=\varrho_{1}\otimes\cdots\otimes\varrho_{m}$ is zero.
  2. (2)
    For each $i$, there exists an integer $j\not=i$ with Goursat datum $(H_{i,j},H_{j,i},\varphi_{ij})$ such that $N_{i}$ is not contained in $H_{i,j}$
Proof.
Replacing $G$ by its Zariski-closure in $G_{1}\times\cdots\times G_{m}$, we may assume that $G\subset G_{1}\times\cdots\times G_{m}$ is an algebraic subgroup and that the maps $\varrho_{i}$ are algebraic. In particular, $\varrho_{i}$ maps $G$ surjectively to $G_{i}$.
We assume that (2) is not satisfied and prove (1). We may assume, up to renumbering the representations, that the condition in this statement fails for $i=1$. For any $j\geqslant 2$, let $(H_{1,j},H_{j,1},\varphi_{1j})$ be the Goursat datum for the image $(\varrho_{1}\times\varrho_{j})(G)$ of $G$ in $G_{1}\times G_{j}$ (Lemma 3.1). Our assumption that (2) is not satisfied is then that $N_{1}\subset H_{1,j}$. In particular, we have $N_{1}\times\{1\}\subset(\varrho_{1}\times\varrho_{j})(G)$ since $\varphi_{1j}(N_{1}H_{1,j})=1_{G_{j}/H_{j,1}}$.
For $2\leqslant l\leqslant m$, let
$K_{l}=\bigcap_{2\leqslant j\leqslant l}\ker(\varrho_{j})\subset G.$
We claim that $N_{1}\subset\varrho_{1}(K_{m})$. To see this, we argue by induction on $2\leqslant l\leqslant m$ that $N_{1}\subset\varrho_{1}(K_{l})$.
The case $l=2$ follows from the previous discussion (for $x\in N_{1}$, the pair $(x,1)$ is in $(\varrho_{1}\times\varrho_{2})(G)$, hence there exists $x^{\prime}\in G$ with $\varrho_{1}(x^{\prime})=x$ and $\varrho_{2}(x^{\prime})=1$). Suppose now that $\varrho_{1}(K_{l-1})\supset N_{1}$ and let us prove the same for $K_{l}=K_{l-1}\cap\ker(\varrho_{l})$. Since $N_{1}$ is perfect, hence generated by commutators $[x,y]$ with $(x,y)\in N_{1}^{2}$, it suffices to show that every such commutator is in $\varrho_{1}(K_{l-1}\cap\ker(\varrho_{l}))$.
By the induction hypothesis (mutatis mutandis), we can find $x^{\prime}\in G$ and $y^{\prime}\in G$ such that
$\varrho_{1}(x^{\prime})=x,\quad\varrho_{j}(x^{\prime})=1\text{ for }2\leqslant j\leqslant l-1,$
$\varrho_{1}(y^{\prime})=y,\quad\varrho_{j}(y^{\prime})=1\text{ for }3\leqslant j\leqslant l.$
We obtain
$\varrho_{1}([x^{\prime},y^{\prime}])=[\varrho_{1}(x^{\prime}),\varrho_{1}(y^{\prime})]=[x,y],$
$\varrho_{j}([x^{\prime},y^{\prime}])=[\varrho_{j}(x^{\prime}),\varrho_{j}(y^{\prime})]=1\quad\text{ for }2\leqslant j\leqslant l,$
(since either $\varrho_{j}(x^{\prime})$ or $\varrho_{j}(y^{\prime})$ equals $1$); this completes the induction (compare with the proof of [31, Lemma 3.3], due to Serre).
The restriction $\bar{\varrho}$ of $\varrho=\varrho_{1}\otimes\cdots\otimes\varrho_{m}$ to the subgroup $K_{m}$ is isomorphic to the restriction of $\varrho_{1}\otimes 1\otimes\cdots\otimes 1$ to $K_{m}$; the space of $G$-coinvariants of $\varrho$ is contained in the space of $K_{m}$-coinvariants of $\bar{\varrho}$, and the latter is contained in the space of $N_{1}$-coinvariants of $\varrho_{1}$ (since $\varrho_{1}(K_{m})$ contains $N_{1}$), which vanishes by assumption. This shows that statement (1) holds and concludes the proof. ∎
Remark 3.1.
(1) In our applications, the desired goal will be (1), and achieving this will be reduced to excluding (2). In practice, we will instead show that this second possibility is restricted to “diagonal” situations. Outside of these special cases, the Riemann Hypothesis (Theorem 2.2) will imply square-root cancellation when the sheaves corresponding to the representations $\varrho_{i}$ have weights $\leqslant 0$.
(2) In contrast to the “classical” theory of the Goursat–Kolchin–Ribet criterion, as in [sumproducts], note that we do not obtain here a clear formula for the main term in the special diagonal cases. Thus the result is mostly useful when one does not wish to extract a precise main term from the analysis of sums of trace functions.
3.3. The orthogonal case
In the “classical” version of the Goursat–Kolchin–Ribet criterion, one has essentially to deal with the situation of Proposition 3.3 where each $G_{i}$ is a simple algebraic group of certain specific types (see [ESDE, Lemma 1.8.5] for instance). The special case where $G$ (or some $G_{i}$) is an orthogonal group $\mathbf{O}_{n}$ in an even number of variables is not covered by these earlier statements (and this explains some restrictions in the type of hyper-Kloosterman sums for which the previous paper [Pisa] obtained variants of Theorem 1.3, see [Pisa, Def. 2.1]).
In this section (which may be omitted in a first reading since it is not directly relevant to our main results), we show how Proposition 3.3 provides a good control of this orthogonal situation. This may be useful, e.g., in the study of families of elliptic curves, where orthogonal monodromy groups occur frequently.
Proposition 3.4.
Let $k$ be an algebraically closed field of characteristic different from $2$. Let $n\geqslant 3$ and $m\geqslant 1$ be integers. Let $\mathbf{O}_{n}$ denote the orthogonal group of the split quadratic form in $n$ variables over $k$. Let $G$ be a group.
For $1\leqslant i\leqslant m$, let $\varrho_{i}\colon G\to\mathbf{GL}(V_{i})$ be an irreducible finite-dimensional representation of $G$ on $k$-vector space $V_{i}$ such that the Zariski closure of the image of $\varrho_{i}$ is $\mathbf{O}_{n}$, and, viewed as a representation of $\mathbf{O}_{n}$, each $V_{i}$ is isomorphic to the same fixed irreducible faithful representation of $\mathbf{O}_{n}$, for example the standard representation.
Then one of the following properties holds:
  1. (1)
    The space of $G$-coinvariants of the representation $\varrho=\varrho_{1}\otimes\cdots\otimes\varrho_{m}$ is zero.
  2. (2)
    For each $i$ with $1\leqslant i\leqslant m$, there exists an integer $j\not=i$ with $1\leqslant j\leqslant m$ and a one-dimensional representation $\chi_{ij}$ of $G$ of order at most $2$ such that $\varrho_{j}$ is isomorphic to $\varrho_{i}\otimes\chi_{ij}$.
We will use the following properties of orthogonal groups, which are probably well-known but for which we do not know of a convenient reference.
Lemma 3.5.
Let $k$ be an algebraically closed field of characteristic different from $2$. Let $n\geqslant 3$ be an integer.
  1. (1)
    Any normal subgroup of $\mathbf{O}_{n}$ which does not contain $\mathbf{SO}_{n}$ is contained in the center $Z_{n}=\{-\mathrm{Id},\mathrm{Id}\}$ of $\mathbf{O}_{n}$.
  2. (2)
    Any automorphism of the quotient group $\mathbf{O}_{n}/Z_{n}$ is inner.
Proof.
We recall that a finite group $H$ normalized by a connected algebraic group $G$ is in fact centralized by this group: indeed, for any $x\in H$, the map $y\mapsto yxy^{-1}$ is a morphism from $H$ to $G$, hence must be constant, equal to the value $x$ taken at $y=1$.
(1) Let $H$ be a normal subgroup of $\mathbf{O}_{n}$. If $n\not=4$, then the subgroup $\mathbf{SO}_{n}$ is a simple algebraic group, hence $H\cap\mathbf{SO}_{n}$ is either equal to $\mathbf{SO}_{n}$ (so that $H$ contains $\mathbf{SO}_{n}$) or contained in the center of $\mathbf{SO}_{n}$; in the second case, it follows that $H$ is finite, and from the fact above, it commutes with $\mathbf{SO}_{n}$, hence is contained in the center by Schur’s Lemma, and then in the center of $\mathbf{O}_{n}$.
Suppose that $n=4$. There exists an isomorphism
$\varphi\colon(\mathbf{SL}_{2}\times\mathbf{SL}_{2})/D\to\mathbf{SO}_{4}$
[…]
In the first case, $H\cap\mathbf{SO}_{4}$ is contained in the kernel of $\pi_{1}\times\pi_{2}$, which is the center of $\mathbf{SO}_{4}$, and it follows again that $H$ is finite, thus contained in the centralizer of $\mathbf{SO}_{4}$ (by the preliminary remark), which is the center of $\mathbf{O}_{4}$.
In the second case, we claim that $H$ contains $\mathbf{SO}_{4}$. To prove this, it suffices to prove that $H$ contains $\varphi(\mathbf{SL}_{2}\times\{1\})$ (since by symmetry $H$ will also contain $\varphi(\{1\}\times\mathbf{SL}_{2})$, and these two subgroups generate $\mathbf{SO}_{4}$). Since $\mathbf{SL}_{2}$ is generated by commutators, it suffices to prove that $H$ contains $\varphi(([x,y],1)D)$ for all $x$ and $y$ in $\mathbf{SL}_{2}$. Since $x\in\pi_{1}(H\cap\mathbf{SO}_{4})$, we must have $\varphi((x,z)D)\in H$ for some $z\in\mathbf{SL}_{2}$. We have
$\varphi(([x,y],1)D)=[\varphi((x,z)D),\varphi((y,1)D)]$
and since $H$ is normal and $\varphi((x,z)D)\in H$, the commutator $\varphi((x,z)D),\varphi((y,1)D)]$ belongs to $H$ as well.
(2) Let $\sigma$ be an automorphism of $\mathbf{O}_{n}/Z_{n}$. Let $\beta\colon\mathbf{O}_{n}/Z_{n}\to\operatorname{Out}(\operatorname{Lie}(\mathbf{O}_{n}))$ be the morphism induced by the conjugation action, and let $\tau=\beta(\sigma)$.
[…]
Proof of Proposition 3.4.
We will apply Proposition 3.3 with $N_{i}=\mathbf{SO}_{n}$ for all $i$, which is perfect since $n>2$. Since $V_{i}$ is an irreducible faithful representation of $\mathbf{O}_{n}$ , we indeed have the space of $N_{i}$-coinvariants of $V_{i}$ zero.
Suppose that the coinvariant space of $\varrho$ is non-zero. We are then in the second case of Proposition 3.3. For each $i$, there exists an integer $j\not=i$ with Goursat datum $(H_{i,j},H_{j,i},\varphi_{ij})$, where $\varphi_{ij}$ is an isomorphism
(3.1) $\varphi_{ij}\colon G_{i}/H_{i,j}\to G_{j}/H_{j,i}$
and $N_{i}=\mathbf{SO}_{n}$ is not contained in $H_{i,j}$.
By Lemma 3.5 (1), we have $H_{i,j}\subset Z_{n}$, so that $H_{j,i}\subset Z_{n}$ also (otherwise $N_{j}=\mathbf{SO}_{n}$, so there could be no isomorphism (3.1)).
Since the group $\mathbf{O}_{n}/Z_{n}$ is not isomorphic to $\mathbf{O}_{n}$ (e.g. because one has trivial center, and the other not), the isomorphism (3.1) is induced either by an automorphism $\mathbf{O}_{n}/Z_{n}\to\mathbf{O}_{n}/Z_{n}$ or by an automorphism $\mathbf{O}_{n}\to\mathbf{O}_{n}$. In either case, $\varphi_{i,j}$ induces an automorphism $\bar{\varphi}_{i,j}\colon\mathbf{O}_{n}/Z_{n}\to\mathbf{O}_{n}/Z_{n}$, and the diagram
$G_{i}$$_{j}GG$$\mathbf{O}_{n}/Z_{n}$$\mathbf{O}_{n}/Z_{n}$$\varrho_{i}$$\varrho_{j}$$\overline{\varphi}_{ij}$
commutes.
We denote by $[x]$ the image of some element $x\in\mathbf{O}_{n}$ by the projection $\mathbf{O}_{n}\to\mathbf{O}_{n}/Z_{n}$. Since all automorphisms of $\mathbf{O}_{n}/Z_{n}$ are inner (Lemma 3.5, (2)), we deduce that the composed maps $G\to G_{i}\to\mathbf{O}_{n}/Z_{n}$ and $G\to G_{j}\to\mathbf{O}_{n}/Z_{n}$ are conjugate, i.e., that there exists $x_{0}\in\mathbf{O}_{n}/Z_{n}$ such that $\varrho_{i}(g)]=x_{0}[\varrho_{j}(g)]x_{0}^{-1}$ for all $g\in G$. We lift $x_{0}$ arbitrarily to an element of $\mathbf{O}_{n}$, without changing notation, and see that $\varrho_{i}(g)=x_{0}\varrho_{j}(g)x_{0}^{-1}\chi(g)$ for some unique element $\chi(g)\in Z_{n}$. The uniqueness implies that $\chi\colon G\to Z_{n}$ is a group morphism (this follows from
$x_{0}\varrho_{j}(g)x_{0}^{-1}\chi(g_{1}g_{2})=\varrho_{i}(g_{1}g_{2})=\varrho_{i}(g_{1})\varrho_{i}(g_{2})=x_{0}\varrho_{j}(g_{1})x_{0}^{-1}\chi(g_{1})x_{0}\varrho_{j}(g_{2})x_{0}^{-1}\chi(g_{2})\\ =x_{0}\varrho_{j}(g_{1})x_{0}^{-1}x_{0}\varrho_{j}(g_{2})x_{0}^{-1}\chi(g_{1})\chi(g_{2})=x_{0}\varrho_{j}(g_{1}g_{2})x_{0}^{-1}\chi(g_{1})\chi(g_{2})$
for $(g_{1},g_{2})\in G^{2}$). Then conjugation by $x_{0}$ is an isomorphism between $\varrho_{i}$ and $\varrho_{j}\otimes\chi$. ∎
3.4. A property of gallant sheaves
The following proposition will be crucial in the proof of our main result. It will be used to prove that the first statement of Proposition 3.3 holds “in most cases” in certain families of sheaves.
We consider in this section a finite field $k$ with an algebraic closure $\bar{k}$, a prime $\ell$ invertible in $k$, and a non-zero integer $c$.
Given $(r,s)\in\bar{k}\times\bar{k}^{\times}$, we define morphisms $\gamma_{r,s}$ and $\gamma_{r,s,c}$ from $\mathbf{A}^{1}_{\bar{k}}$ to $\mathbf{A}^{1}_{\bar{k}}$ by
$\gamma_{r,s}(v)=s(v+r),\quad\quad\gamma_{r,s,c}(v)=s(v+r)^{c}.$
Note that $\gamma_{r,s}$ is an element of the group $\mathrm{Aff}(\bar{k})$ of affine transformations of $\mathbf{A}^{1}_{\bar{k}}$.
For any $\ell$-adic sheaf $\mathscr{F}$, we denote
$\mathscr{F}_{r,s}=\gamma_{r,s}^{*}\mathscr{F},\quad\quad\mathscr{F}_{r,s,c}=\gamma_{r,s,c}^{*}\mathscr{F},$
which are again $\ell$-adic sheaves on $\mathbf{A}^{1}_{\bar{k}}$. These satisfy obvious properties which we will use below, such as
$(\mathscr{F}_{1}\oplus\mathscr{F}_{2})_{r,s,c}=(\mathscr{F}_{1})_{r,s,c}\oplus(\mathscr{F}_{2})_{r,s,c},$
and similarly for tensor products.
Let $(r,s)\in\bar{k}\times\bar{k}^{\times}$. Let $\sigma\in\bar{k}$ be any element such that $\sigma^{c}=s$. Then we have
(3.2) $\mathscr{F}_{r,s,c}=[v\mapsto s(v+r)^{c}]^{*}\mathscr{F}=[u\mapsto\sigma(v+r)]^{*}[v\mapsto v^{c}]^{*}\mathscr{F}=\gamma_{r,\sigma}^{*}\mathscr{F}_{0,1,c}.$
In particular, the arithmetic (resp. geometric) monodromy group of $\mathscr{F}_{r,s,c}$ is (up to isomorphism) independent of $(r,s)$, and coincides with the arithmetic (resp. geometric) monodromy group of $\mathscr{F}_{0,1,c}$. The geometric monodromy group $G$ of $\mathscr{F}_{0,1,c}$ can be identified with a normal subgroup of the geometric monodromy group group of $\mathscr{F}$, such that the corresponding quotient is cyclic of order dividing $c$. In particular, $G$ is finite if and only if the geometric monodromy group of $\mathscr{F}$ is finite, and if $G$ is gallant, then $G$ contains (in all cases) the core subgroup $N$, since the latter is a perfect normal subgroup of the geometric monodromy group of $\mathscr{F}$.
Proposition 3.6.
Let $k$ be a finite field. Let $c$ be a non-zero integer invertible in $k$. Let $\mathscr{F}$ be a gallant $\ell$-adic sheaf over $k$. Denote by $N$ the core subgroup of the geometric monodromy group of $\mathscr{F}$.
If the characteristic $q$ of $k$ is large enough, in terms of $c$ and the complexity of $\mathscr{F}$, then there exists a finite cyclic subgroup $T$ of $\mathrm{Aff}(\bar{k})$ such that the following properties hold:
  1. (1)
    the order of $T$ is bounded in terms of $c$ and the complexity of $\mathscr{F}$, with the bound polynomial if the monodromy group of $\mathscr{F}$ is infinite;
  2. (2)
    for all $(r_{1},s_{1})$ and $(r_{2},s_{2})$ in $\bar{k}\times\bar{k}^{\times}$, the geometric monodromy group of the sheaf
    $\mathscr{F}_{r_{1},s_{1},c}\oplus\mathscr{F}_{r_{2},s_{2},c}$
    contains $N\times N$ except possibly if there exist $c$-th roots $\sigma_{1}$ and $\sigma_{2}$ of $s_{1}$ and $s_{2}$ respectively such that
    $\gamma_{r_{2},\sigma_{2}}\in\gamma_{r_{1},\sigma_{1}}T.$
In the proof, we will denote by $\mathscr{G}$ the sheaf $\mathscr{F}_{0,1,c}=[u\mapsto u^{c}]^{*}\mathscr{F}$. Its complexity is bounded polynomially in terms of $c$ and the complexity of $\mathscr{F}$ by [qst, Th. 6.8] (or by elementary properties of the Fouvry–Michel–Kowalski complexity, in this case). We denote by $G$ the geometric monodromy group of $\mathscr{G}$; recall that it contains $N$.
We will distinguish the cases where $G$ is finite or infinite. The former is significantly more involved, and readers interested primarily in the infinite case (which is the one relevant for most current applications) can skip it.
Proof when $G$ is infinite.
In this case, recall that $G^{0}=N$. The arithmetic monodromy group $G^{a}$ of $\mathscr{G}$ normalizes $G$, and hence it normalizes its connected component $G^{0}$, which is a characteristic subgroup of $G$. It follows that the Lie algebra $\operatorname{Lie}(G^{0})$ is a subrepresentation of the adjoint representation of $G^{a}$. This representation of $G^{a}$ corresponds to an $\ell$-adic sheaf $\mathscr{H}$ on $\mathbf{A}^{1}_{k}$, and it is geometrically irreducible because $G^{0}$ is simple by assumption.
We define $T$ to be the group of affine linear transformations $\gamma\in\mathrm{Aff}(\bar{k})$ such that $\gamma^{*}\mathscr{H}$ is geometrically isomorphic to $\mathscr{H}$. We will check that this group has all the required properties whenever $q$ is large enough in terms of $c$ and the complexity of $\mathscr{F}$.
Since the adjoint representation of a closed subgroup of $\mathbf{GL}(V)$ is isomorphic to the tensor product of its “tautological” representation with its contragredient, the sheaf $\mathscr{H}$ is a summand of the sheaf $\mathscr{G}\otimes\mathscr{G}^{\vee}$. Its complexity is therefore bounded in terms of $c$ and (polynomially) of the complexity of $\mathscr{F}$ (see [qst, Th. 6.8, Prop. 6.14]).
With the notation of Proposition 2.3, applied to $\mathscr{H}$, we have $T=\mathbf{B}_{\mathscr{H}}$. Since $\mathscr{F}$ is assumed to be gallant, we note that $\mathscr{H}$ is geometrically irreducible (because $N=G^{0}$ acts irreducibly, by definition), so that the proposition applies.
The first possible conclusion of Proposition 2.3 does not hold, because the generic rank of $\mathscr{H}$ is the dimension of the Lie algebra of $G^{0}$, which is $\geqslant 2$ (again since $G^{0}=N$ is a simple algebraic group by definition). The third cannot be valid if $q$ is sufficiently large in terms of the complexity of $\mathscr{H}$. Hence the second holds, and it implies that $T$ is a finite cyclic group of order bounded polynomially in terms of the complexity of $\mathscr{H}$, hence also in terms of $c$ and the complexity of $\mathscr{F}$.
Consider now $(r_{1},s_{1})$ and $(r_{2},s_{2})$ in $\bar{k}\times\bar{k}^{\times}$, and assume that the geometric monodromy group $K$ of the sheaf $\mathscr{F}_{r_{1},s_{1},c}\oplus\mathscr{F}_{r_{2},s_{2},c}$ does not contain $N\times N$ (note that the connected component of the geometric monodromy group of each summand coincides with that of $\mathscr{G}$, which is $G$, so this potential inclusion makes sense).
Let $(H_{1},H_{2},\varphi)$ be a Goursat datum for the subgroup $K\subset G\times G$. At least one of $H_{1}$ or $H_{2}$ does not contain $N$, since otherwise we would have $N\times N\subset K$.
Since $H_{i}$ is a normal subgroup of $G$, its Lie algebra $\operatorname{Lie}(H_{i})$ is an ideal in $\operatorname{Lie}(G)=\operatorname{Lie}(G^{0})$. Since $G^{0}$ is a simple linear algebraic group, its Lie algebra is simple, so that $\operatorname{Lie}(H_{i})$ is either 0 or $\operatorname{Lie}(G)$, and similarly for $\operatorname{Lie}(G/H_{i})$. The isomorphism $\varphi\colon G/H_{1}\to G/H_{2}$ implies that the Lie algebras of $G/H_{1}$ and $G/H_{2}$ are isomorphic, so either both are zero, or both are equal to $\operatorname{Lie}(G)$. In the first case, it would follow that $G/H_{i}$ is finite, so that $H_{i}\supset G^{0}=N$ for $i=1$, 22, which we saw is not the case. Thus we have $\operatorname{Lie}(G/H_{i})=\operatorname{Lie}(G)$ for $i=1$, 22. The isomorphism $\varphi\colon G/H_{1}\to G/H_{2}$ then induces an automorphism
$\widetilde{\varphi}\colon\operatorname{Lie}(G)\to\operatorname{Lie}(G).$
Since $\widetilde{\varphi}$ is compatible with the $G$-action, this linear map is a (geometric) isomorphism of the corresponding $\ell$-adic representations of the Galois group of the function field. These correspond to the sheaves
$\mathscr{F}_{r_{1},s_{1},c}\otimes\mathscr{F}_{r_{1},s_{1},c}^{\vee}\quad\text{ and }\quad\mathscr{F}_{r_{2},s_{2},c}\otimes\mathscr{F}_{r_{2},s_{2},c}^{\vee},$
respectively. However, since $\mathscr{F}_{r_{i},s_{i},c}=\gamma_{r_{i},\sigma_{i}}^{*}\mathscr{G}$ whenever $\sigma_{i}^{c}=s_{i}$ (see (3.2)), and similarly for the contragredient, we conclude that that the isomorphism above is a geometric isomorphism
$\gamma_{r_{1},\sigma_{1}}^{*}\mathscr{G}\to\gamma_{r_{2},\sigma_{2}}^{*}\mathscr{G},$
and hence $\gamma_{r_{1},\sigma_{1}}\gamma_{r_{2},\sigma_{2}}^{-1}\in T$. This conclusion is what we desired. ∎
Before we treat the case when $G$ is finite, we prove two group-theoretical lemmas, one of which is classical and the second likely well-known.
Lemma 3.7.
Let $G$ be a finite group. Let $M$ and $N$ be normal subgroups of $G$. If $N$ is quasisimple and $M$ does not contain $N$, then $M$ commutes with $N$.
Proof.
Since $M$ and $N$ are both normal subgroups, the commutator subgroup $[M,N]$ (generated by commutators $m,n]=mnm^{-1}n^{-1}$ with $(m,n)\in M\times N$) is contained in $M\cap N$ (since $m,n]=m(nm^{-1}n^{-1})=(mnm^{-1})n$). If $M$ does not contain $N$, then $M\cap N$ is a proper normal subgroup of $N$, hence is contained in the center of $N$. Thus $N,M],N]=[[M,N],N]=\{1\}$ and then a classical lemma (see, e.g., [isaacs, Lemma 4.9]) implies that $N,N],M]=\{1\}$. Since $N$ is perfect, this means that $[N,M]=1$, which is the desired conclusion. ∎
Lemma 3.8.
Let $E$ be an algebraically closed field, $r\geqslant 0$ an integer and let $G\subset\mathbf{GL}_{r}(E)$ be a finite gallant group. The core subgroup $N$ in $G$ is a characteristic subgroup of $G$, i.e., it is invariant by all group automorphisms of $G$.
Proof.
Let $\varphi\in\operatorname{Aut}(G)$ be an automorphism of $G$ and let $M=\varphi(N)$. By Lemma 3.7, either $N\subset M$, in which case either $M=N$, since both groups have the same order, or $M$ would commute with $N$. But since $N$ acts irreducibly on $E^{r}$, it would then follow from Schur’s Lemma that $M$ is abelian, which is a impossible. ∎
Proof of Proposition 3.6 when $G$ is finite.
By definition of a gallant group, the finite group $G$ contains the quasisimple normal subgroup $N$, which acts irreducibly. We denote by $A$ the center of $N$ and by $S$ the quotient $S=N/A=N/Z(N)$, which is a non-abelian simple group. We view $G$ as a subgroup of $\mathbf{GL}(V)$, where $V$ is the space on which the corresponding representation $\varrho\colon\Gamma\to\mathbf{GL}(V)$ of the Galois group $\Gamma$ of the field $k(X)$ acts.
Since $N$ acts irreducibly on $V$, it follows from Schur’s Lemma that $A$ is a group of scalar matrices; each of these must have order dividing $|A|$, so this group is in fact the cyclic group $\boldsymbol{\mu}_{|A|}\mathrm{Id}$ of scalar matrices of order dividing $|A|$.
Since $A$ is central, the conjugation action of $G$ on $N$ induces a homomorphism
$\alpha\colon G\to\operatorname{Aut}(N/A)=\operatorname{Aut}(S).$
Let $(r,s)\in\bar{k}\times\bar{k}^{\times}$. The sheaf $\mathscr{F}_{r,s,c}$ determines a group homomorphism of the geometric Galois group $\Gamma^{g}$ of $\bar{k}(X)$ to $G$, and thus by composition with $\alpha$ a group homomorphism
$\varrho_{r,s,c}\colon\Gamma^{g}\to\operatorname{Aut}(S).$
Note that $\mathrm{Aff}(\bar{k})$ acts naturally on $\Gamma^{g}$, hence on the set $X_{S}$ of homomorphisms $\Gamma^{g}\to\operatorname{Aut}(S)$ modulo $\operatorname{Aut}(S)$-conjugacy. Recalling that we defined $\mathscr{G}=\mathscr{F}_{0,1,c}$, we denote by $T$ the stabilizer (for this action of $\mathrm{Aff}(\bar{k})$) of the class in $X_{S}$ of the homomorphism $\varrho_{0,1,c}$ associated to $\mathscr{G}$. In other words, $T$ is the set of affine transformations $\gamma$ such that the homomorphisms associated to $\mathscr{G}$ and $\gamma^{*}\mathscr{G}$ are conjugate by an element of $\operatorname{Aut}(S)$.
To conclude the proof, we will now prove that the subgroup $T$ has the desired properties.
Step 1. For $q$ large enough, the group $T$ is a finite cyclic group of size depending only on $c$ and the complexity of $\mathscr{F}$.
We note first that the definition of $T$ makes it clear that $T$ is a subgroup of $\mathrm{Aff}(\bar{k})$. Any subgroup of $\mathrm{Aff}(\bar{k})$ of order coprime to the characteristic $q$ of $k$ is cyclic, since its intersection with the kernel of the projection $\mathrm{Aff}(\bar{k})\to\bar{k}^{\times}$ is trivial and any finite subgroup of $\bar{k}^{\times}$ is cyclic. So it suffices to check that $T$ is finite of size depending only on $c$ and the complexity of $\mathscr{F}$.
By Lemma 3.8, the subgroup $N$ is a characteristic subgroup of $G$. In particular, the action by conjugation of the arithmetic monodromy group $G^{a}$ of $\mathscr{G}$ on its normal subgroup $G$ gives an action of $G^{a}$ on $N$; this action is trivial on $Z(N)$, since the latter consists of scalar matrices, hence we obtain an action of $G^{a}$ on $S$, or equivalently a morphism
$\beta\colon G^{a}\to\operatorname{Aut}(S).$
Since $N$ acts irreducibly on $V$, Schur’s Lemma implies that the kernel of this action is contained in the group of scalar matrices. It follows that the linear action of $G^{a}$ on $\operatorname{End}(V)$, which is also by conjugation, factors through the image of $\beta$.
Let $W=\overline{\mathbf{Q}}_{\ell}[\operatorname{Aut}(S)]$ be the regular representation of $\operatorname{Aut}(S)$. Let $\mathscr{H}$ be the sheaf on $\mathbf{A}^{1}_{k}$ associated to $W$. Since the monodromy group of $\mathscr{H}$ is finite, this sheaf geometrically semisimple. After replacing $k$ by a finite extension, if necessary, we may assume that $\mathscr{H}$ splits as a direct sum of geometrically irreducible sheaves. Let then $I$ be the finite set of geometrically irreducible components which arise in this decomposition.
Any $\gamma\in T$ gives a geometric isomorphism $\gamma^{*}\mathscr{H}\simeq\mathscr{H}$ arising from the action on $W$ of the element of $\operatorname{Aut}(S)$ that conjugates the homomorphism associated to $\gamma^{*}\mathscr{G}$ to the homomorphism associated to $\mathscr{G}$, and therefore permutes the set $I$. Since the size of $I$ is at most $|S|$, we will obtain the conclusion of Step 1 once we show that the stabilizer of some irreducible component $\mathscr{H}^{\prime}\in I$ is finite of size bounded in terms of $c$ and of the complexity of $\mathscr{F}$.
But the stabilizer in $\mathrm{Aff}(\bar{k})$ of the isomorphism class of any component $\mathscr{H}^{\prime}$ coincides by definition with the group $\mathbf{B}_{\mathscr{H}^{\prime}}$. Thus, we may argue using Proposition 2.3 again. Note that since $\mathscr{H}^{\prime}$ is a summand of the $k$th tensor power of $\mathscr{G}$ for $k$ bounded in terms of $S$ and thus bounded in terms of $\operatorname{rank}(\mathscr{G})$, its complexity is bounded in terms of $c$ and the complexity of $\mathscr{F}$. We obtain the desired conclusion for $q$ large enough in terms of the complexity of $\mathscr{F}$ provided $\mathscr{H}^{\prime}$ does not have generic rank $1$. Such a component $\mathscr{H}^{\prime}$ exists because $W$ is a non-trivial representation of $S$.
This concludes the proof of Step 1.
Step 2. Let $(r_{1},s_{1})$ and $(r_{2},s_{2})\in\bar{k}\times\bar{k}^{\times}$ be such that the geometric monodromy group of the sheaf $\mathscr{F}_{r_{1},s_{1},c}\oplus\mathscr{F}_{r_{2},s_{2},c}$ does not contain $N\times N$ as a subgroup; we need to check that we then have
$\gamma_{\sigma_{2},r_{2}}\in\gamma_{\sigma_{1},r_{1}}T,$
where $\sigma_{i}$ is an arbitrary element of $\bar{k}$ with $\sigma_{i}^{c}=s_{i}$, and this will conclude the proof of Proposition 3.6 when $G$ is finite.
Recall that the geometric monodromy groups of $\mathscr{F}_{r_{i},s_{i},c}$ are both identified with $G$. Let $K$ be the geometric monodromy group of the direct sum. Let $(H_{1},H_{2},\varphi)$ be a Goursat datum for $K\subset G\times G$ (Lemma 3.1). Since we assume that $K$ does not contain $N\times N$, we know that either $H_{1}$ or $H_{2}$ does not contain $N$.
Step 2.1. We claim that $H_{1}$ and $H_{2}$ are both contained in $Z(G)$.
Indeed, if $H_{1}$ does not contain $N$, then by Lemma 3.7 (applied to $H_{1}$ and $N$), we see that $H_{1}$ is contained in the centralizer of $N$ in $G$, and in particular is abelian by Schur’s Lemma. Thus, if we consider a composition series for $G$ built out of composition series for
$G/H_{1},\quad H_{1},$
we see that the only non-abelian factors are those in $G/H_{1}$. By uniqueness of Jordan–Hölder factors, a composition series built from
$G/H_{2},\quad H_{2}$
already provides all the non-abelian factors from $G/H_{2}$, since $\varphi\colon G/H_{1}\to G/H_{2}$ is an isomorphism. This implies that $H_{2}$ has an abelian composition series, i.e., that it is solvable. In particular, it is also true that $H_{2}$ does not contain $N$, and so is also contained in the centralizer of $N$.
Similary, exchanging the roles of $H_{1}$ and $H_{2}$, we conclude that both are always contained in the centralizer of $N$. We then recall that Schur’s Lemma shows that the latter is contained in $Z(G)$.
Step 2.2. We have $\varphi(NH_{1}/H_{1})=NH_{2}/H_{2}$.
Indeed, it suffices to prove that $\varphi(NH_{1}/H_{1})\supset NH_{2}/H_{2}$, and then to exchange the role of $H_{1}$ and $H_{2}$ to get the converse inclusion.
Let $p\colon G\to G/H_{2}$ be the projection. The subgroup $N^{*}=p^{-1}(\varphi(NH_{1}/H_{1}))$ is a normal subgroup of $G$, so by Lemma 3.7, we have $N\subset N^{*}$ (otherwise $N^{*}$ would commute with $N$, hence would be abelian by Schur’s Lemma, which is not the case). Thus $N\subset N^{*}$, which means that $NH_{2}/H_{2}\subset\varphi(NH_{1}/H_{1})$; this proves the claim.
Step 2.3. Conclusion.
Let $M_{i}=NH_{i}/H_{i}$; since $M_{i}$ is a non-trivial image of a quasi-simple group, it is quasi-simple and the quotient isomorphism $N/Z(N)\to S$ induces an isomorphism $M_{i}/Z(M_{i})\to S$. It follows that the isomorphism $\varphi\colon NH_{1}/H_{1}\to NH_{2}/H_{2}$ induces an automorphism $\widetilde{\varphi}\colon S\to S$. By construction, $\widetilde{\varphi}$ conjugates the actions of $G^{a}$ on $S$ associated to $\mathscr{F}_{r_{1},s_{1},c}$ and $\mathscr{F}_{r_{2},s_{2},c}$, which means that $\gamma_{r_{1},s_{1}}\gamma_{r_{2},s_{2}}^{-1}\in T$. ∎
3.5. The case of oxozonic sheaves
This section will only be used in the proof of Theorem 1.6. It provides a version of Proposition 3.6 adapted to oxozonic sheaves.
We use the same notation as in the previous section for the sheaves $\mathscr{F}_{r,s,c}$, etc.
Proposition 3.9.
Let $k$ be a finite field. Let $c$ be an odd integer invertible in $k$. Let $\mathscr{F}$ be an oxozonic $\ell$-adic sheaf over $k$.
If the characteristic $q$ of $k$ is large enough, in terms of $c$ and the complexity of $\mathscr{F}$, then there exists a finite cyclic subgroup $T$ of $\mathrm{Aff}(\bar{k})$ such that the following properties hold:
  1. (1)
    the order of $T$ is bounded polynomially in terms of $c$ and the complexity of $\mathscr{F}$ only;
  2. (2)
    for all $(r_{1},s_{1})$ and $(r_{2},s_{2})$ in $\bar{k}\times\bar{k}^{\times}$, the geometric monodromy group of the sheaf
    $\mathscr{F}_{r_{1},s_{1},c}\oplus\mathscr{F}_{r_{2},s_{2},c}$
    contains $\mathbf{SO}_{4}\times\mathbf{SO}_{4}$ except possibly if there exist $c$-th roots $\sigma_{1}$ and $\sigma_{2}$ of $s_{1}$ and $s_{2}$ respectively such that
    $\gamma_{r_{2},\sigma_{2}}\in\gamma_{r_{1},\sigma_{1}}T.$
Proof.
Since $c$ is odd, the geometric monodromy group $G$ of the sheaf $v\mapsto v^{c}]^{*}\mathscr{F}$ is still isomorphic to $\mathbf{O}_{4}$.
We denote by $\mathscr{G}$ the sheaf associated to the action of the arithmetic monodromy group on the Lie algebra $\operatorname{Lie}(\mathbf{SO}_{4})=\operatorname{Lie}(\mathbf{SL}_{2})\oplus\operatorname{Lie}(\mathbf{SL}_{2})$, and define $T$ to be the group of affine linear transformations $\gamma$ such that $\gamma^{*}\mathscr{G}$ is geometrically isomorphic to $\mathscr{G}$. As in the proof of the infinite case of Proposition 3.6, we can then show that $T$ is a finite cyclic group of order bounded polynomially in terms of $c$ and the complexity of $\mathscr{F}$.
Let $(r_{1},s_{1})$ and $(r_{2},s_{2})$ be pairs such that geometric monodromy group of the sheaf $\mathscr{F}_{r_{1},s_{1},c}\oplus\mathscr{F}_{r_{2},s_{2},c}$ fails to contain $\mathbf{SO}_{4}\times\mathbf{SO}_{4}$ as a subgroup. By Lemma 3.1, we obtain a Goursat datum $(H_{1},H_{2},\varphi)$ where $H_{i}$ is a normal subgroup of $G$ and $\varphi\colon G/H_{1}\to G/H_{2}$ is an isomorphism, such that $\mathbf{SO}_{4}$ is not contained in both $H_{1}$ and $H_{2}$.
If $\mathbf{SO}_{4}$ is not contained in $H_{1}$, then $H_{1}=\{1\}$ or $H_{1}=Z_{n}$ (Lemma 3.5, (1)), and the isomorphism $\varphi$ implies that either $H_{2}=H_{1}=\{1\}$ or $H_{2}=H_{1}=Z_{n}$. Mutatis mutandis, the same is checked in the case where $\mathbf{SO}_{4}$ is not contained in $H_{2}$.
Now, arguing as in the end of the proof of Corollary 3.4 we conclude that $\mathscr{F}_{r_{2},s_{2},c}$ is geometrically isomorphic to $\mathscr{F}_{r_{1},s_{1},c}$, up to a possible quadratic twist, and this means that for $c$-th roots $\sigma_{i}$ of $s_{i}$, we we have $\gamma_{r_{2},\sigma_{2}}\in\gamma_{r_{1},\sigma_{1}}T$. ∎
3.6. Diagonal estimates
Combining Proposition 3.6 and Proposition 3.3, we will ultimately seek to obtain estimates for sums of products of pullbacks of a gallant sheaf by the transformations $\gamma_{r,s,c}$. We prove here, in different situations, that “most” of the sheaves which arise in sums of products of a certain type have trivial topmost cohomology group, a property which translates in bounds for the corresponding sums after applying the Riemann Hypothesis over finite fields.
We will use repeatedly the following simple combinatorial lemma. By “graph”, we mean undirected graphs without multiple edges, but where loops are allowed.
Lemma 3.10.
Let $(V,E)$ be a finite graph, with maximal degree $C\geqslant 0$. Let $m\geqslant 1$ be an integer. Denote
[…]
We have
$|\Delta_{m}(V,E)|\ll|V|^{m},$
where the implied constant depends only on $m$ and $C$.
Proof.
We denote by $[2m]$ the set $\{1,\ldots,2m\}$.
For any $\text{${v}$}=(v_{1},\ldots,v_{2m})\in\Delta_{m}(V,E)$, define a graph $\Theta_{\text{${v}$}}$ with vertex set $[2m]$ and edges joining $i$ to $j$ if and only if $v_{i}$ is connected to $v_{j}$ in $(V,E)$.
Since the number of possibles graphs $\Theta_{\text{${v}$}}$ is bounded in terms of $m$ and $C$ only, we obtain the estimate
$|\Delta_{m}(V,E)|\ll\max_{\Theta}|\{\text{${v}$}\in V^{2m}\,\mid\,\Theta_{\text{${v}$}}=\Theta\}|,$
where the implied constant depends only on $l$ and $C$, and the maximum ranges over graphs $\Theta$ which arise as $\Theta_{\text{${w}$}}$ for some $\text{${w}$}\in\Delta_{m}(V,E)$.
[…]
Let $I\subset[2m]$ be one connected component of $\Theta$, and $i_{0}\in I$. If $\text{${v}$}\in V^{2m}$ satisfies $\Theta_{\text{${v}$}}=\Theta$, then all $v_{i}$ for $i\in I$ are contained in the ball $B(v_{i_{0}},|I|)$ of radius $|I|$ in $\Gamma$ around $v_{i_{0}}$. This ball has $\leqslant C^{|I|}$ elements, so it follows that there are at most
$\sum_{v\in V}|B(v;|I|)|\leqslant C^{|I|}|V|$
[…]
$C^{2m}|V|^{\pi_{0}}\leqslant C^{2m}|V|^{m}$
[…]
This concludes the proof. ∎
We now assume given the data $(k,\mathscr{F})$ of the previous section. For $m\geqslant 1$, and for $(\text{${r}$},\text{${s}$})\in k^{2m}\times(k^{\times})^{2m}$, we denote
(3.3) $\mathscr{F}_{\text{${r}$},\text{${s}$},c}=\bigotimes_{1\leqslant j\leqslant m}\mathscr{F}_{r_{j},s_{j},c}\otimes\mathscr{F}_{r_{j+m},s_{j+m},c}^{\vee}.$
The first diagonal statement is the following, which is enough to deal with Type I sums.
Proposition 3.11.
Let $\mathscr{F}$ be a gallant sheaf. Let $m\geqslant 1$ be an integer. Let
$\Delta_{m,c}(\mathscr{F};k)=\{(\text{${r}$},\text{${s}$})\in k^{2m}\times(k^{\times})^{2m}\,\mid\,H^{2}_{c}(\mathbf{A}^{1}_{\bar{k}},\mathscr{F}_{\text{${r}$},\text{${s}$},c})\not=0\}.$
Then we have
$|\Delta_{m,c}(\mathscr{F};k)|\ll|k|^{2m}$
where the implied constant depends on $m$ and on $c$ and the complexity of $\mathscr{F}$.
Proof.
Let $G$ denote the connected component of the identity of the geometric monodromy group of $\mathscr{F}$. Let $T\subset\mathrm{Aff}(\bar{k})$ denote the finite subgroup determined by Proposition 3.6.
Let $\varrho_{r,s,c}$ denote the Galois representation associated to $\mathscr{F}_{r,s,c}$, and let $G_{r,s,c}$ denote the geometric monodromy group of $\mathscr{F}_{r,s,c}$. Let $G_{\text{${r}$},\text{${s}$},c}$ be the geometric monodromy group of the sheaf
$\bigoplus_{1\leqslant j\leqslant 2m}\mathscr{F}_{r_{j},s_{j},c},$
which we view as a subgroup of
$\prod_{1\leqslant j\leqslant 2m}G_{r_{j},s_{j},c}.$
By the co-invariant formula for the top-degree cohomology of a sheaf, the condition
$H^{2}_{c}(\mathbf{A}^{1}_{\bar{k}},\mathscr{F}_{\text{${r}$},\text{${s}$},c})\not=0$
is equivalent to the vanishing of the co-invariant space of $G_{\text{${r}$},\text{${s}$},c}$ acting by the representation
$\bigotimes_{1\leqslant j\leqslant m}\varrho_{r_{j},s_{j},c}\otimes\varrho_{r_{j+m},s_{j+m},c}^{\vee}.$
If this space is non-zero, then Proposition 3.3 (applied with $N_{i}$ equal to the core subgroup of the geometric monodromy group of $\mathscr{F}$ for all $i$) proves, that for each $i\leqslant 2m$, there exists $j\not=i$ such that the image of $G$ in $G_{r_{i},s_{i},c}\times G_{r_{j},s_{j},c}$ does not contain $N\times N$. By Proposition 3.6, this implies that $\gamma_{r_{i},s_{i},c}\gamma_{r_{j},s_{j},c}^{-1}\in T$.
Define a graph $\Gamma=(V,E)$ with vertices $k\times k^{\times}$ and edges joining $(r_{1},s_{1})$ to $(r_{2},s_{2})$ if $\gamma_{r_{i},s_{i}}\gamma_{r_{j},s_{j}}^{-1}\in T$ (in particular, note that each vertex has a loop). The previous results then show that, with the notation of Lemma 3.10, the inclusion
$\Delta_{m,c}(\mathscr{F};k)\subset\Delta_{m}(\Gamma)$
holds, and since the degree of $\Gamma$ is $|T|$, which is bounded in terms of $c$ and the complexity of $\mathscr{F}$, Lemma 3.10 concludes the proof. ∎
The second statement will occur in the study of Type II sums.
Proposition 3.12.
Let $\mathscr{F}$ be a gallant sheaf. Let $m\geqslant 1$ be an integer. Let $d\geqslant 1$ be the order of the cyclic group $T$ from Proposition 3.6 applied to $\mathscr{F}$. Let
$Y_{d}=\{(s_{1},s_{2})\in{\mathbf{G}_{m}}\times{\mathbf{G}_{m}}\,\mid\,s^{d}_{1}\not=s^{d}_{2}\}.$
Let
$\Delta^{\prime}_{m,c}(\mathscr{F};k)=\{(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})\in k^{2m}\times Y_{d}(k)^{2m}\,\mid\,H^{2}_{c}(\mathbf{A}^{1}_{\bar{k}},\mathscr{F}_{\text{${r}$},\text{${s}$}_{1},c}\otimes\mathscr{F}_{\text{${r}$},\text{${s}$}_{2},c}^{\vee})\not=0\}.$
Then we have
$|\Delta^{\prime}_{m,c}(\mathscr{F};k)|\ll|k|^{3m}$
where the implied constant depends on $m$, $c$, $d$ and the complexity of $\mathscr{F}$.
Proof.
Since
$\mathscr{F}_{\text{${r}$},\text{${s}$}_{1},c}\otimes\mathscr{F}_{\text{${r}$},\text{${s}$}_{2},c}^{\vee}=\Bigl(\bigotimes_{1\leqslant j\leqslant m}\varrho_{r_{j},s_{1,j},c}\otimes\varrho_{r_{j+m},s_{1,j+m},c}^{\vee}\Bigr)\otimes\Bigl(\bigotimes_{1\leqslant j\leqslant m}\varrho_{r_{j},s_{2,j},c}^{\vee}\otimes\varrho_{r_{j+m},s_{2,j+m},c}\Bigr),$
we see as in the proof of the previous proposition (keeping its notation) that the non-vanishing condition is now equivalent with the non-vanishing of the coinvariant space of the representation ${\bigotimes_{1\leqslant j\leqslant 4m}\varrho_{j}}$ where
[…]
We denote by $\delta_{j}$ the element $\gamma_{r,s}\in\mathrm{Aff}(\bar{k})$ corresponding to each $\varrho_{j}$ (e.g., $\delta_{j}=\gamma_{r_{j},s_{2,j}}$ if $m+1\leqslant j\leqslant 2m$).
Suppose that the coinvariant space is non-zero. By Proposition 3.6, there exists for each positive integer $j\leqslant 4m$ an integer $i\not=j$ with $1\leqslant i\leqslant 4m$ and $\delta_{j}\delta_{i}^{-1}\in T$. In particular, since $T$ is of order $d$, it follows that the coefficients $\sigma_{j}$, $\sigma_{i}$ such that $\delta_{j}=\gamma_{\tau_{j},\sigma_{j}}$ and $\delta_{i}=\gamma_{\tau_{i},\sigma_{i}}$ (for some $\tau_{i}$, $\tau_{j}$) satisfy $\sigma_{i}^{d}=\sigma_{j}^{d}$. A first application of Lemma 3.10 establishes that the number of possible pairs $(\text{${s}$}_{1},\text{${s}$}_{2})\in(k^{\times})^{4m}$ with this property is $O(|k|^{2m})$.
Let $\text{${s}$}=(\text{${s}$}_{1},\text{${s}$}_{2})\in Y_{d}(k)^{2m}$. We will now bound (uniformly in terms of these tuples) the size of the set $\Delta_{\text{${s}$}_{1},\text{${s}$}_{2}}$ of those $\text{${r}$}\in k^{2m}$ such that $(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})\in\Delta^{\prime}_{m,c}(\mathscr{F};k)$.
[…]
$\Delta_{\text{${s}$}_{1},\text{${s}$}_{2}}\subset\Delta_{m}(V,E),$
in which case Lemma 3.10 proves that $|\Delta_{\text{${s}$}_{1},\text{${s}$}_{2}}|\ll|k|^{m}$, uniformly in terms of $(\text{${s}$}_{1},\text{${s}$}_{2})$, and therefore $|\Delta^{\prime}_{m,c}(\mathscr{F};k)|\ll|k|^{3m}$, as desired.
Let $\text{${r}$}=(r_{j})_{1\leqslant j\leqslant 2m}\in\Delta_{\text{${s}$}_{1},\text{${s}$}_{2}}$. We prove that the vertex $r_{1}$ of $\Gamma$ is connected to a vertex $r_{j}$ for some $j\not=1$. The same argument, up to notation, will apply to $r_{2}$, …, $r_{2m}$, and then conclude the proof.
We use the notation $(\delta_{j})_{1\leqslant j\leqslant 4m}$ introduced above; we have therefore $\delta_{1}=\gamma_{r_{1},s_{1,1}}$. Since $(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})$ is in $\Delta^{\prime}_{m,c}(\mathscr{F};k)$, we have seen above that there exists an integer $i$ with $2\leqslant i\leqslant 4m$ such that $\delta_{1}\delta_{i}^{-1}\in T$. If $\delta_{i}=\gamma_{r_{j},\sigma}$ for some $j\geqslant 2$, then we are done. But the only other possibility would be that $\delta_{i}=\gamma_{r_{1},s_{2,1}}$, and this is not the case since $s_{1,1}^{d}\not=s_{2,1}^{d}$, and a fortiori $s_{1,1}\not=s_{2,1}$. ∎
Finally, in the case of Type II sums, we will also require the following result, which is of a slightly different nature than the previous ones.
Proposition 3.13.
Let $\mathscr{F}$ be a gallant sheaf over $k$. Let $l\geqslant 1$ be an integer and let $c\geqslant 1$ be an integer invertible in $k$.
[…]
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i})^{c}s]^{*}\mathscr{F}\otimes\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i+l})^{c}s]^{*}\mathscr{F}^{\vee}\Bigr)\not=0.$
The constant $C$ and the degree of $\mathscr{V}$ depend only on $c$, $l$ and the complexity of $\mathscr{F}$.
Proof.
Let $G$ denote the connected component of the identity of the geometric monodromy group of $\mathscr{F}$ and $N$ its core subgroup. Let $T\subset\mathrm{Aff}(\bar{k})$ denote the finite subgroup determined by Proposition 3.6. It has order bounded in terms of $c$ and the complexity of $\mathscr{F}$.
For $u\in\mathbf{A}^{1}$, we denote here $\mathscr{G}_{u}=[s\mapsto us]^{*}\mathscr{F}$ (note that we do not exclude the possibility that $u=0$, in which case $\mathscr{G}_{u}$ is zero), so that we are considering the cohomology group
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}\mathscr{G}_{(r+v_{i})^{c}}\otimes\mathscr{G}_{(r+v_{i+l})^{c}}^{\vee}\Bigr).$
We first consider the case $r=0$, and $\text{${v}$}\in({k^{\times}})^{2l}$. Note that for $v\not=0$, we have $\mathscr{G}_{v^{c}}=\mathscr{F}_{0,v^{c}}$. Arguing as in the beginning of the proof of Proposition 3.11 and applying Proposition 3.6, we see that if
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}\mathscr{G}_{v_{i}^{c}}\otimes\mathscr{G}_{v_{i+l}^{c}}^{\vee}\Bigr)$
is non-zero, then for all $i$ with $1\leqslant i\leqslant 2l$, there exists $j\not=i$ such that $\gamma_{0,v_{i}^{c}}\gamma_{0,v_{j}^{c}}^{-1}\in T$.
[…]
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}[s\mapsto v_{i}^{c}s]^{*}\mathscr{F}\otimes\bigotimes_{i=1}^{l}[s\mapsto v_{i+l}^{c}s]^{*}\mathscr{F}^{\vee}\Bigr)$
[…]
Let $k_{n}$ be a finite extension of $k$. Let $\Gamma_{n}=(V_{n},E_{n})$ be the graph with vertex set $k_{n}^{\times}$ and with an edge from $v$ to $w$ if and only if $\gamma_{0,v^{c}}\gamma_{0,w^{c}}^{-1}\in T$. Applying Lemma 3.10, we deduce that
$|\mathscr{W}(k_{n})|\leqslant\Delta_{l}(\Gamma_{n})\ll|k_{n}|^{l},$
where the implied constant depends on $l$ and the size of $T$ only. If follows (by the Lang–Weil bound for instance) that $\dim(\mathscr{W})\leqslant l$.
We now define $\mathscr{V}$ to be the closure in $\mathbf{A}^{2l}$ of $\mathscr{W}$, and we claim that $\mathscr{V}$ has the desired properties. First, by construction, it remains of dimension $\leqslant l$ and with bounded degree.
Let $\text{${v}$}\in k_{n}^{2l}$. For $r\in k_{n}$, the condition
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i})^{c}s]^{*}\mathscr{F}\otimes\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i+l})^{c}s]^{*}\mathscr{F}^{\vee}\Bigr)\not=\{0\}$
implies that $r+\text{${v}$}\in\mathscr{V}(k_{n})$. The set of $r\in\mathbf{A}^{1}$ such that $r+\text{${v}$}\in\mathscr{V}$ is a closed subvariety of $\mathbf{A}^{1}$. If it is a proper subvariety, then it is a finite set of size bounded in terms of the degree of $\mathscr{V}$. Otherwise, taking $r=0$, we obtain $\text{${v}$}\in\mathscr{V}$. ∎
Remark 3.2.
A typical situation is when $\mathscr{W}$ is defined by a union of varieties defined by equations stating pairwise equalities like
$v_{1}=v_{2},\ \ldots,\ v_{2l-1}=v_{2l},$
with the added condition that the coordinates are invertible. Then $\mathscr{V}$ is defined by the same equations without invertibility conditions, and we see that the equations are invariant under translation by a common element $r$.
The following corollary illustrates in general the use of this type of results. It will be applied in the proof of Theorem 1.4, but is also of independent interest.
Corollary 3.14.
Let $q$ be a prime number and let $\mathscr{F}$ be a gallant and light sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$. Let $l\geqslant 1$ be an integer. Let $K$ be the trace function of $\mathscr{F}$.
There exists an algebraic subvariety $\mathscr{V}$ of $\mathbf{A}^{2l}_{\mathbf{F}_{q}}$ of dimension $\leqslant l$ and degree bounded in terms of $l$ and the complexity of $\mathscr{F}$ such that the estimate
(3.4) $\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}\ll q^{1/2}$
[…]
Proof.
Let $j\colon U\to\mathbf{A}^{1}$ be the open immersion of a dense open set where $\mathscr{F}$ is lisse, and let $\mathscr{F}^{*}$ be the middle-extension sheaf $j_{*}j^{*}\mathscr{F}$. It is a gallant and light sheaf whose trace function $K^{*}$ coincides with that of $K$ except for a set of values of size bounded in terms of the complexity of $\mathscr{F}$. Hence, for any $\text{${u}$}\in\mathbf{F}_{q}^{2l}$, the estimate
$\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}=\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K^{*}(u_{i}v)\overline{K^{*}(u_{i+l}v)}+O(1)$
holds, where the implied constant depends on $l$ and the complexity of $\mathscr{F}$. In particular, the bound (3.4) is equivalent to its analogue for $K^{*}$.
For given $\text{${u}$}\in\mathbf{F}_{q}^{2l}$, the function
$v\mapsto\prod_{i=1}^{l}K^{*}(u_{i}v)\overline{K^{*}(u_{i+l}v)}$
is the trace function of the sheaf
$\mathscr{G}_{\text{${u}$}}=\bigotimes_{i=1}^{l}[s\mapsto v_{i}s]^{*}\mathscr{F}^{*}\otimes\bigotimes_{i=1}^{l}[s\mapsto v_{i+l}s]^{*}(\mathscr{F}^{*})^{\vee}$
which is the sheaf appearing in Proposition 3.13 for the sheaf $\mathscr{F}^{*}$, with $c=1$ and $r=0$. The result then follows with the algebraic variety $\mathscr{V}$ given by the proposition (in view of the fact that the exceptional values $r$ in loc. cit. are all non-zero), by the Riemann Hypothesis applied to $\mathscr{G}_{\text{${u}$}}$ (see Theorem 2.2). ∎
Remark $3.$
Suppose that $\mathscr{F}$ is an oxozonic sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$. Using Proposition 3.9 instead of Proposition 3.6, we see that Propositions 3.11, 3.12 and 3.13 are also valid for $\mathscr{F}$.
4. General estimates for bilinear forms
4.1. Reduction to complete sums
The main result of this section is Proposition 4.1, which gives a general reduction of bounds for bilinear forms (type I or II) with kernel defined modulo a prime modulus $q$ to estimates for certain “complete” sums. It is not needed here to assume that $K$ is a trace function.
We use throughout the following notation. We denote by $q$ a prime number, and by $K\colon\mathbf{F}_{q}\to\mathbf{C}$ a function which we identify with a $q$-periodic function on $\mathbf{Z}$. We fix two non-zero integers $b$ and $c$.
We also fix two positive integers $M$ and $N$. We then denote
$B_{b,c}(\text{${\alpha}$},N;K)$ $=\sum_{m\sim M}\sum_{n\sim N}\alpha_{m}K(m^{b}n^{c}),$
$B_{b,c}(\text{${\alpha}$},\text{${\beta}$};K)$ $=\sum_{m\sim M}\sum_{n\sim N}\alpha_{m}\beta_{n}K(m^{b}n^{c}),$
for arbitrary families $\text{${\alpha}$}=(\alpha_{m})_{m\sim M}$ and $\text{${\beta}$}=(\beta_{n})_{n\sim N}$ of complex numbers. In fact, since $K$, $b$ and $c$ are fixed in this section, we will most often abbreviate and write
$B(\text{${\alpha}$},N)=B_{b,c}(\text{${\alpha}$},N;K),\quad B(\text{${\alpha}$},\text{${\beta}$})=B_{b,c}(\text{${\alpha}$},\text{${\beta}$};K)$
We denote
$\|\alpha\|_{2}=\Bigl(\sum_{m\sim M}|\alpha_{m}|^{2}\Bigr)^{1/2},$
and similarly for $\|\beta\|_{2}$.
Let $l\geqslant 2$ be integer. For $r$ and $s$ in $\mathbf{F}_{q}$ and for
$\text{${v}$}=(v_{1},\cdots,v_{2l})\in\mathbf{F}_{q}^{2l},$
we define
$\mathbf{K}_{c}(r,s,\text{${v}$})=\prod_{i=1}^{l}K(s(r+v_{i})^{c})\overline{K(s(r+v_{i+l})^{c})}.$
The reduction statement is the following.
Proposition 4.1.
We keep the notation above. Let $l\geqslant 1$ be an integer and let $M,N,V\geqslant 1$ be real numbers satisfying
(4.1) $V\leqslant\frac{N}{10},\quad M,N,\frac{N^{2}}{V}\leqslant q.$
  1. (1)
    For any $\varepsilon>0$, we have
    (4.2) $B(\text{${\alpha}$},N)\ll q^{\varepsilon}{\|\text{${\alpha}$}\|_{2}M^{1/2}N}\Bigl(\frac{1}{MN^{2}V^{2l-1}}\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$})|\Bigr)^{1/(2l)}$
    where
    $\Sigma_{I}(\text{${v}$})=\mathop{\sum\sum}\limits_{(r,s)\in\mathbf{F}_{q}\times\mathbf{F}^{\times}_{q}}\mathbf{K}_{c}(r,s,\text{${v}$})$
    and the implied constant depends only on $\varepsilon$ and $l$.
  2. (2)
    Let $d\geqslant 1$ be an integer. For any $\varepsilon>0$, we have
    (4.3) $B(\text{${\alpha}$},\text{${\beta}$})\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}(MN)^{1/2}\Bigl(\frac{1}{M}+\Bigl(\frac{1}{M^{2}N^{2}V^{2l-1}}\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\Bigr)^{1/(2l)}\Bigr)^{1/2},$
    where
    $\Sigma_{II}^{(d)}(\text{${v}$})=\sum_{r\in\mathbf{F}_{q}}\mathop{\sum\sum}\limits_{\begin{subarray}{c}s_{1},s_{2}\in\mathbf{F}^{\times}_{q}\\ s^{d}_{1}\not=s^{d}_{2}\end{subarray}}\mathbf{K}_{c}(r,s_{1},\text{${v}$})\overline{\mathbf{K}_{c}(r,s_{2},\text{${v}$})}$
    and the implied constant depends only on $\varepsilon$, $\|K\|_{\infty}$, $d$ and $l$.
Remark 4.1.
It is easy to check that the dependency on $\|K\|_{\infty}$ of the implied constant in (4.3) can be removed if one replaces the term $M^{-1}$ on the right-hand side with $\|K\|_{\infty}^{2}M^{-1}$, either by examining the proof or by an amplification argument. However, the quantity $\|K\|_{\infty}$ will be uniformly bounded in our applications.
Remark 4.2.
We recall how such bounds can lead to results like Theorem 1.3. We assume that $K$ is bounded by an absolute constant (in practice, the complexity of the underlying object of which $K$ is the trace function). Since the factors
$\|\text{${\alpha}$}\|_{2}M^{1/2}N,\quad\quad\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}(MN)^{1/2}$
appearing in the estimates represent the “trivial” bounds for $B(\text{${\alpha}$},N)$ and $B(\text{${\alpha}$},\text{${\beta}$})$, the subsequent factors are the potential savings. How much one does save depends on our success in bounding the sums
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$})|,\quad\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{d)}(\text{${v}$})|,$
depending on the unspecified parameters $V$ and (for the second sum) $d$.
[…]
$\Sigma_{I}(\text{${v}$})\ll q$
[…]
$v_{i}=v_{i+l},\quad\quad i=1,\cdots,l$
for which only the trivial bound holds. If there are not many more diagonal cases, we can therefore hope to obtain
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$})|\ll V^{l}q^{2}+V^{2l}q,$
which is $\ll q^{3}$ when choosing $V=q^{1/l}$. This bound turns out to imply the estimates for Type I sums.
For the type II sums, interpolating similarly between the trivial bound $\Sigma_{II}^{(d)}(\text{${v}$})\ll q^{3}$ along a diagonal set and a generic square-root cancellation bound $\Sigma_{II}^{(d)}(\text{${v}$})\ll q^{3/2}$, we can hope to obtain
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\ll V^{l}q^{3}+V^{2l}q^{3/2}$
which is $\ll q^{9/2}$ when choosing $V=q^{3/(2l)}$. This leads to the bounds for Type II sums.
Although this intuitive explanation is not too far from the truth, we will see that, as in previous works, the actual details are more complicated and additional “stratification” will be used.
4.2. Proof of Proposition 4.1
We first prove the bound for the Type I sums.
Let $V\geqslant 1$ such that (4.1) holds. We define $U=N/(10V)$, so that $U\geqslant 1$.
Let $f_{N}$ denote the piecewise linear function on $\mathbf{R}$ which is equal to $1$ in the interval $[N,2N]$ and is zero in $\mathopen{]}-\infty,N-1]\cup[2N+1,\infty\mathclose{[}$. Let $\widehat{f}_{N}$ be the Fourier transform of $f_{N}$, defined by
$\widehat{f}_{N}(t)=\int_{\mathbf{R}}f_{N}(s)e(-st)ds$
for $t\in\mathbf{R}$. The estimate
$|\widehat{f}_{N}(t)|\ll\Xi_{N}(t)$
holds, where $\Xi_{N}(t)=\min(\log N,|t|^{-1},|t|^{-2})$.
By elementary changes of variable, we have
$B(\text{${\alpha}$},N)$ $=\frac{1}{UV}\sum_{(u,v)\sim U\times V}\mathop{\sum\sum}\limits_{m\sim M,n+uv\sim N}\alpha_{m}K(m^{b}(n+uv)^{c})$
$=\frac{1}{UV}\sum_{(u,v)\sim U\times V}\mathop{\sum\sum}\limits_{m\sim M,n\in[N/2,2N]}\alpha_{m}f_{N}(n+uv)K(m^{b}(n+uv)^{c}).$
Applying the Fourier inversion formula,
$f_{N}(s)=\int_{\mathbf{R}}\widehat{f}_{N}(t)e(ts)dt,$
we deduce that
$B(\text{${\alpha}$},N)$ $=\frac{1}{UV}\int_{\mathbf{R}}\sum_{(u,v)\sim U\times V}\mathop{\sum\sum}\limits_{m\sim M,n\in[N/2,2N]}\alpha_{m}e(t(n+uv))K(m^{b}(n+uv)^{c})\widehat{f_{N}}(t)dt$
$\leqslant\frac{1}{UV}\int_{\mathbf{R}}\mathop{\sum\sum}\limits_{m\sim M,n\in[N/2,2N]}|\alpha_{m}|\sum_{u\sim U}\Bigl|\sum_{v\sim V}e(tv)K(n^{c}m^{b}(\overline{u}n+v)^{c})\Bigr|\frac{1}{u}\Bigl|\widehat{f_{N}}\Bigl(\frac{t}{u}\Bigr)\Bigr|dt$
$\ll\frac{1}{UV}\int_{\mathbf{R}}\mathop{\sum\sum}\limits_{m\sim M,n\in[N/2,2N]}|\alpha_{m}|\sum_{u\sim U}\Bigl|\sum_{v\sim V}e(tv)K(n^{c}m^{b}(\overline{u}n+v)^{c})\Bigr|\frac{1}{U}\Xi_{N}\Bigl(\frac{t}{U}\Bigr)dt$
$\ll\frac{\log N}{UV}\mathop{\sum\sum}\limits_{m\sim M,n\in[N/2,2N]}|\alpha_{m}|\sum_{u\sim U}\Bigl|\sum_{v\sim V}\varepsilon(v)K(n^{c}m^{b}(\overline{u}n+v)^{c})\Bigr|$
where $\varepsilon(v)=e(t_{0}v)$ for some $t_{0}\in\mathbf{R}$. In the last step, we have also used the fact that
$\int_{\mathbf{R}}\frac{1}{U}\Xi_{N}\Bigl(\frac{t}{U}\Bigr)dt=O(\log N).$
Given $(r,s)\in\mathbf{F}^{\times}_{q}\times\mathbf{F}^{\times}_{q}$ we define
$\nu(r,s)=\mathop{\sum\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle u^{c}m^{b}\equiv s\,(\mathrm{mod}\,{q})}}}{{{\scriptstyle\overline{u}n\equiv r\,(\mathrm{mod}\,{q})}}}}|\alpha_{m}|,$
where the variables are integers restricted to satisfy the size conditions
(4.4) $m\sim M,\quad n\sim N,\quad u\sim U,\quad v\sim V.$
Applying Hölder’s inequality, we deduce from the previous inequality that the estimate
$B(\text{${\alpha}$},N)\ll\frac{\log N}{UV}\Bigl(\sum_{r,s}|\nu(r,s)|\Bigr)^{1-\frac{1}{l}}\Bigl(\sum_{r,s}|\nu(r,s)|^{2}\Bigr)^{\frac{1}{2l}}\Bigl(\sum_{r,s}\Bigl|\sum_{v\sim V}\varepsilon(v)K(s(r+v)^{c})\Bigr|^{2l}\Bigr)^{\frac{1}{2l}}$
holds.
Lemma 4.2.
We have
(4.5) $\sum_{r,s}|\nu(r,s)|\ll UN\|\text{${\alpha}$}\|_{1}\leqslant UNM^{1/2}\|\text{${\alpha}$}\|_{2}$
and
(4.6) $\sum_{r,s}|\nu(r,s)|^{2}\ll q^{\varepsilon}\|\alpha\|_{2}^{2}UN\Bigl(1+\frac{UN}{q}\Bigr)\Bigl(1+\frac{M}{q}\Bigr)$
for any $\varepsilon>0$, where the implied constant depends on $\varepsilon$ and $b$.
Proof.
The bound (4.5) follows immediately from (4.4).
For (4.6), we write
$\sum_{r,s}|\nu(r,s)|^{2}=\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle u_{1},u_{2},n_{1},n_{2}}}}{{{\scriptstyle u_{2}n_{1}\equiv u_{1}n_{2}}}}}\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle m_{1},m_{2}}}}{{{\scriptstyle u_{1}^{c}m_{1}^{a}\equiv u_{2}^{c}m_{2}^{a}}}}}\alpha_{m_{1}}\overline{\alpha_{m_{2}}}\ll\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle u_{1},u_{2},n_{1},n_{2}}}}{{{\scriptstyle u_{2}n_{1}\equiv u_{1}n_{2}}}}}\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle m_{1},m_{2}}}}{{{\scriptstyle u_{1}^{c}m_{1}^{a}\equiv u_{2}^{c}m_{2}^{a}}}}}|\alpha_{m_{1}}|^{2}$
since $|\alpha_{m_{1}}\overline{\alpha_{m_{2}}}|\leqslant|\alpha_{m_{1}}|^{2}+|\alpha_{m_{2}}|^{2}$.
We express the congruence $u_{2}n_{1}\equiv u_{1}n_{2}\,(\mathrm{mod}\,{q})$ as an equality $u_{1}n_{2}=u_{2}n_{1}+qk$ for some integer $k$ such that $|k|\ll(UN/q+1)$ according to the size conditions above. Fixing $u_{2}$, $n_{1}$ and $k$, there are at most $(UN)^{\varepsilon}\ll q^{\varepsilon}$ solutions $(u_{1},n_{2})$ of these equations for any $\varepsilon>0$. Moreover, for any given $(u_{1},u_{2},m_{1})$, the number of $m_{2}$ satisfying the congruence
$u_{1}^{c}m_{1}^{b}\equiv u_{2}^{c}m_{2}^{b}\,(\mathrm{mod}\,{q})$
is $\ll M/q+1\ll 1$, where the implied constant depends on $b$. Combining these bounds concludes the proof of the lemma. ∎
Applying Lemma 4.2 and the bound $|\varepsilon(v)|\leqslant 1$ for all $v$, we deduce the estimate
$B(\text{${\alpha}$},N)\ll q^{\varepsilon}\frac{\|\text{${\alpha}$}\|_{2}^{1-1/l}M^{1/2}\|\text{${\alpha}$}\|_{2}^{1/l}N}{V(UMN)^{1/(2l)}}\Bigl(\sum_{\text{${v}$}\sim[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$};q)|\Bigr)^{1/(2l)}$
for any $\varepsilon>0$, where the implied constant depends on $\varepsilon$ and $b$.
We now prove the Type II bound, and we will be brief. We denote again $U=N/(10V)$.
Applying the Cauchy–Schwarz inequality and following the argument of [Pisa]*§4.1, we deduce the estimate
$B(\text{${\alpha}$},\text{${\beta}$})\ll\|\text{${\beta}$}\|_{2}(\|\text{${\alpha}$}\|^{2}_{2}N+S^{\not\equiv})^{1/2}$
where the implied constant is absolute and
$S^{\not\equiv}=\mathop{\sum\sum}\limits_{m_{1}^{bd}\not\equiv m_{2}^{bd}}\alpha_{m_{1}}\overline{\alpha_{m_{2}}}\sum_{n\sim N}K(m_{1}^{b}n^{c})\overline{K}(m_{2}^{b}n^{c})\\ \ll\frac{\log q}{UV}\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle r,s_{1},s_{2}}}}{{{\scriptstyle s_{1}^{d}\neq s_{2}^{d}}}}}\nu(r,s_{1},s_{2})\Bigl|\sum_{v\sim V}\varepsilon(v)K(s_{1}(r+v)^{c})\overline{K}(s_{1}(r+v)^{c})\Bigr|$
where $\varepsilon(v)=e(t_{0}v)$ for some $t_{0}\in\mathbf{R}$ and we define
$\nu(r,s_{1},s_{2})=\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle u,n,\overline{u}n\equiv r}}}{{{\scriptstyle u^{c}m^{b}_{i}\equiv s_{i},\ i=1,2}}}}|\alpha_{m_{1}}{\alpha_{m_{2}}}|,$
for $(r,s_{1},s_{2})\in(\mathbf{F}^{\times}_{q})^{3}$, the variables being constrained by the size conditions
$u\sim U,\quad m_{1},m_{2}\sim M,\quad n\sim N.$
Lemma 4.3.
Assume that $M\leqslant q$ and $UN\leqslant q$. The estimates
$\sum_{r,s_{1},s_{2}}\nu(r,s_{1},s_{2})\ll UMN\|\text{${\alpha}$}\|^{2}_{2},$
with an absolute implied constant, and
$\sum_{r,s_{1},s_{2}}\nu(r,s_{1},s_{2})^{2}\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}^{4}UN,$
for any $\varepsilon>0$, hold, where the implied constants depend only on $b$ and $c$ and $\varepsilon$.
Proof.
The first bound follows from the elementary estimates
$\sum_{r,s_{1},s_{2}}\nu(r,s_{1},s_{2})\ll\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle u,n}}}{{{\scriptstyle m_{1},m_{2}}}}}|\alpha_{m_{1}}\alpha_{m_{2}}|\ll UN\|\text{${\alpha}$}\|_{1}^{2}\ll UMN\|\text{${\alpha}$}\|_{2}^{2},$
where the implied constant is absolute.
To prove the second bound, we first note that
$\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu(r,s_{1},s_{2})^{2}\ll\mathop{\sum\sum}\limits_{u_{1},n_{1},m_{1},m_{2}}|\alpha_{m_{1}}{\alpha_{m_{2}}}|\mathop{\sum\sum}\limits_{\begin{subarray}{c}u_{2},n_{2},\mu_{1},\mu_{2}\\ \overline{u}_{2}n_{2}\equiv\overline{u}_{1}n_{1}\\ u_{2}^{c}\mu_{i}^{b}\equiv u_{1}^{c}\mu_{i}^{b},\ i=1,2\end{subarray}}|\alpha_{\mu_{1}}{\alpha_{\mu_{2}}}|.$
We express the congruence $\overline{u}_{2}n_{2}\equiv\overline{u}_{1}n_{1}$ as an equation $u_{2}n_{1}=u_{1}n_{2}+qk$ for some integer $k\ll UN/q$ (by the size conditions). Once $u$, $n_{2}$ and $k$ are fixed, there are $\ll(UN)^{\varepsilon}$ solutions $(u_{2},n)$ of the equations, for any $\varepsilon>0$, where the implied constant depends only on $\varepsilon$. Writing
$|\alpha_{m_{1}}{\alpha_{m_{2}}}|\leqslant|\alpha_{m_{1}}|^{2}+|{\alpha_{m_{2}}}|^{2}$
and using the estimate
$\mathop{\sum\sum}\limits_{u_{2}^{c}\mu_{i}^{b}\equiv u_{1}^{c}m_{i}^{b}}|\alpha_{m_{i}}|^{2}\ll\|\text{${\alpha}$}\|_{2}^{2}\Bigl(\frac{M}{q}+1\Bigr)$
for $i=1$ and $2$, where the implied constant depends on $b$, we see that
$\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu(r,s_{1},s_{2})^{2}\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}^{4}\Bigl(\frac{M}{q}+1\Bigr)^{2}UN\Bigl(1+\frac{UN}{q}\Bigr)\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}^{4}UN,$
which concludes the proof. ∎
Applying Hölder’s inequality twice, as in [Pisa, § 4.1], we deduce the bound
$\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu^{1-\frac{1}{l}+\frac{1}{l}}\Bigl|\sum_{v\sim V}\cdots\Bigr|$ $\leqslant\Bigl(\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu\Bigr)^{1-\frac{1}{l}}\Bigl(\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu\Bigl|\sum_{v\sim V}\cdots\Bigr|^{l}\Bigr)^{\frac{1}{l}}$
$\leqslant\Bigl(\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu\Bigr)^{1-\frac{1}{l}}\Bigl(\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\nu^{2}\Bigr)^{\frac{1}{2l}}\Bigl(\mathop{\sum\sum}\limits_{r,s_{1},s_{2}}\Bigl|\sum_{v\sim V}\cdots\Bigr|^{2l}\Bigr)^{\frac{1}{2l}}$
and using Lemma 4.3 (under the assumption $M,UN\leqslant q$), we conclude that
$S^{\not\equiv}$ $\ll\frac{q^{\varepsilon}}{UV}\|\text{${\alpha}$}\|_{2}^{2}(UMN)^{1-1/l}(UN)^{1/(2l)}\Bigl(\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}(K;\text{${v}$})|\Bigr)^{1/(2l)}$
$\ll\frac{q^{\varepsilon}}{N}\|\text{${\alpha}$}\|_{2}^{2}(MN^{2}/V)^{1-1/l}(N^{2}/V)^{1/(2l)}\Bigl(\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}(K;\text{${v}$})|\Bigr)^{1/(2l)}$
$\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}^{2}MN\Bigl(\frac{1}{M^{2}N^{2}V^{2l-1}}\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}(K;\text{${v}$})|\Bigr)^{1/(2l)},$
for any $\varepsilon>0$.
This finishes the proof of Proposition 4.1.
4.3. Reduction to a stratification statement
We now indicate the precise step which we will use to implement Proposition 4.1 in the case where $K$ is a trace function. However, we do not yet need to make this assumption.
We use the notation from the previous section. Recall our convention for the degree of an algebraic variety which is not necessarily irreducible.
Proposition 4.4.
Let $q$ be a prime number, let $K\colon\mathbf{F}_{q}\to\mathbf{C}$ be a function, and let $l\geqslant 2$ be an integer. Let $C\geqslant 1$ be a real number.
  1. (1)
    Suppose that there exist algebraic varieties $\mathscr{V}_{1}$ and $\mathscr{V}_{1}^{\Delta}$ over $\mathbf{F}_{q}$ such that
    $\mathbf{A}^{2l}_{\mathbf{F}_{q}}\supset\mathscr{V}_{1}\supset\mathscr{V}_{1}^{\Delta}$
    $\dim(\mathscr{V}_{1})\leqslant l+\Bigl\lceil\frac{l}{2}\Bigr\rceil,\quad\dim(\mathscr{V}^{\Delta}_{1})\leqslant l,$
    $\max(\deg(\mathscr{V}_{1}),\deg(\mathscr{V}_{1}^{\Delta}))\leqslant C$
    and such that the estimates
    (4.7) $|\Sigma_{I}(\text{${v}$})|\leqslant\begin{cases}Cq&\text{ for }\text{${v}$}\not\in\mathscr{V}_{1}(\mathbf{F}_{q}),\\ Cq^{3/2}&\text{ for }\text{${v}$}\not\in\mathscr{V}^{\Delta}_{1}(\mathbf{F}_{q}),\\ Cq^{2}&\text{ for }\text{${v}$}\in\mathscr{V}_{1}^{\Delta}(\mathbf{F}_{q}),\end{cases}$
    hold. We then have
    (4.8) $\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$})|\ll q^{3+1/(2l)},$
    for $V=q^{1/l}$, where the implied constant depends only on $C$.
  2. (2)
    Suppose that there exist algebraic varieties $\mathscr{V}_{2}$ and $\mathscr{V}_{2}^{\Delta}$ over $\mathbf{F}_{q}$ such that
    $\mathbf{A}^{2l}_{\mathbf{F}_{q}}\supset\mathscr{V}_{2}\supset\mathscr{V}_{2}^{\Delta}$
    $\dim(\mathscr{V}_{2})\leqslant l+2\Bigl\lceil\frac{l}{3}\Bigr\rceil,\quad\dim(\mathscr{V}^{\Delta}_{2})\leqslant l,$
    $\max(\deg(\mathscr{V}_{2}),\deg(\mathscr{V}_{2}^{\Delta}))\leqslant C,$
    and such that the estimates
    (4.9) $|\Sigma_{II}^{(d)}(\text{${v}$})|\leqslant\begin{cases}Cq^{3/2}&\text{ for }\text{${v}$}\not\in\mathscr{V}_{2}(\mathbf{F}_{q}),\\ Cq^{2}&\text{ for }\text{${v}$}\not\in\mathscr{V}^{\Delta}_{2}(\mathbf{F}_{q}),\\ Cq^{3}&\text{ for }\text{${v}$}\in\mathscr{V}_{2}^{\Delta}(\mathbf{F}_{q}),\end{cases}$
    hold. We then have
    (4.10) $\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\ll q^{9/2+2/l},$
    for $V=q^{3/(2l)}$, where the implied constant depends only on $C$.
Proof.
(1) Applying Lemma 2.1, the assumption implies that
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{I}(\text{${v}$})|$ $\ll V^{l}q^{2}+V^{\dim(\mathscr{V}_{1})}q^{3/2}+V^{2l}q$
$\ll V^{l}q^{2}+V^{l+\lceil\frac{l}{2}\rceil}q^{3/2}+V^{2l}q\leqslant V^{l}q^{2}+V^{{3l}/{2}+1/2}q^{3/2}+V^{2l}q,$
where the implied constant depends on $C$. Picking $V=q^{1/l}$, we obtain the conclusion.
(2) Similarly, we get
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\ll V^{l}q^{3}+V^{l+2\lceil\frac{l}{3}\rceil}q^{2}+V^{2l}q^{3/2}\leqslant V^{l}q^{3}+V^{5{l}/{3}+4/3}q^{2}+V^{2l}q^{3/2},$
and conclude after taking $V=q^{3/(2l)}$. ∎
Remark 4.3.
If $l$ is divisible by 22 (resp. by 33) the exponent $1/(2l)$ in (4.8) (resp. $2/l$ in (4.10)) can be replaced by 0.
Remark 4.4.
Note a crucial difference with the similar argument in [Pisa, Th. 4.5]: we do not require that the subvarieties $\mathscr{V}_{i}$ and $\mathscr{V}_{i}^{\Delta}$ be defined over $\mathbf{Z}$. However, this means that we require a bound on the degree of these varieties. In our application, this will be provided by an application of Quantitative Sheaf Theory.
5. Moment estimates for gallant sheaves
This and the next section contains the core of the proof of Theorem 1.3. We will establish that the assumptions of Proposition 4.4 are satisfied if $K$ is the trace function of a gallant sheaf $\mathscr{F}$ modulo $q$, with the constant $C$ depending only on the complexity of $\mathscr{F}$. This will be deduced from Xu’s idea (Theorem 2.4), which we implement in the next section.
Let $q$ be a prime number and $\mathscr{F}$ an $\ell$-adic sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$ for some prime $\ell\not=q$. We assume in this section that $\mathscr{F}$ is gallant and light and is a middle-extension sheaf. We suppose also given non-zero integers $b$ and $c$ coprime to $q$, an integer $l\geqslant 1$, and an integer $d\geqslant 1$ coprime to $q$.
We fix an algebraic closure $\overline{\mathbf{F}}_{q}$ of $\mathbf{F}_{q}$. For any finite extension $k/\mathbf{F}_{q}$ contained in $\overline{\mathbf{F}}_{q}$ and $x\in k$, we denote
$K(x;k)=t_{\mathscr{F}}(x;k),$
the values of the trace function of $\mathscr{F}$ over $k$. For $(r,s)\in k\times k^{\times}$ and $\text{${v}$}\in k^{2l}$, we define
$\mathbf{K}_{c}(r,s,\text{${v}$};k)=\prod_{i=1}^{l}K(s(r+v_{i})^{c};k)\overline{K(s(r+v_{i+l})^{c};k)}$
and
$\Sigma_{I}(\text{${v}$};k)=\mathop{\sum\sum}\limits_{(r,s)\in k\times k^{\times}}\mathbf{K}_{c}(r,s,\text{${v}$};k),$
$\Sigma_{II}^{(d)}(\text{${v}$};k)=\sum_{r\in k}\mathop{\sum\sum}\limits_{\begin{subarray}{c}s_{1},s_{2}\in k^{\times}\\ s^{d}_{1}\not=s^{d}_{2}\end{subarray}}\mathbf{K}_{c}(r,s_{1},\text{${v}$};k)\overline{\mathbf{K}_{c}(r,s_{2},\text{${v}$};k)}.$
Thus, for $k=\mathbf{F}_{q}$, these sums are equal to the sums $\Sigma_{I}(\text{${v}$})$ and $\Sigma_{II}^{(d)}(\text{${v}$})$ of the previous section.
Proposition 5.1.
Assume that $q$ is large enough so that Proposition 3.6 applies to $\mathscr{F}$ over $\mathbf{F}_{q}$, and let $d$ be the order of the finite cyclic group $T$ of loc. cit. for $\mathscr{F}$. Let $m\geqslant 0$ be an integer. With notation as above, the estimates
(5.1) $\sum_{\text{${v}$}\in k^{2l}}|\Sigma_{I}(\text{${v}$};k)|^{2m}$ $\ll\left\lvert k\right\rvert^{2m+2l}+\left\lvert k\right\rvert^{4m+l},$
(5.2) $\sum_{\text{${v}$}\in k^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$};k)|^{2m}$ $\ll\left\lvert k\right\rvert^{3m+2l}+\left\lvert k\right\rvert^{6m+l}$
hold for all finite extensions $k/\mathbf{F}_{q}$.
The first step of the proof is to express these moments in a different way in terms of one variable sums. We use for this purpose a simple lemma.
Lemma 5.2.
Let $X$ and $Y$ be finite sets and $\alpha\colon X\times Y\to\mathbf{C}$ a function on $X\times Y$. For integers $l\geqslant 1$ and $m\geqslant 1$, the equality
$\sum_{\text{${x}$}\in X^{2l}}\Bigl|\sum_{y\in Y}\beta(\text{${x}$},y)\Bigr|^{2m}=\sum_{\text{${y}$}\in Y^{2m}}\Bigl|\sum_{x\in X}\gamma(x,\text{${y}$})\Bigr|^{2l}$
holds, where
$\beta(\text{${x}$},y)=\prod_{i=1}^{l}\alpha(x_{i},y)\overline{\alpha(x_{i+l},y)},$
$\gamma(x,\text{${y}$})=\prod_{j=1}^{m}\alpha(x,y_{j})\overline{\alpha(x,y_{j+l})}.$
Proof.
Opening out fully either side of the equality, we see that the resulting sums over $\text{${x}$}\in X^{2l}$ and $\text{${y}$}\in Y^{2m}$ have exactly the same terms. ∎
Applied with suitable choices, this lemma implies the formula
(5.3) $\sum_{\text{${v}$}\in k^{2l}}\left\lvert\Sigma_{I}(\text{${v}$};k)\right\rvert^{2m}=\sum_{(\text{${r}$},\text{${s}$})\in(k\times{k^{\times}})^{2m}}\Bigl|\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${r}$},\text{${s}$},v;k)\Bigr|^{2l}$
where
$\mathbf{K}^{\prime}_{c}(\text{${r}$},\text{${s}$},v;k)=\prod_{j=1}^{m}K(s_{j}(v+r_{j})^{c};k)\overline{K(s_{j+m}(v+r_{j+m})^{c};k)}.$
Similarly, defining
$Y_{d}=\{(s_{1},s_{2})\in{\mathbf{G}_{m}}\times{\mathbf{G}_{m}}\,\mid\,s^{d}_{1}\not=s^{d}_{2}\},$
we obtain the identity
(5.4) $\sum_{\text{${v}$}\in k^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$};k)|^{2m}=\sum_{\text{${r}$}\in k^{2m}}\sum_{(\text{${s}$}_{1},\text{${s}$}_{2})\in Y_{d}(k)^{2m}}\Bigl|\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${r}$},\text{${s}$}_{1},v;k)\overline{\mathbf{K}^{\prime}_{c}(\text{${r}$},\text{${s}$}_{2},v;k)}\Bigr|^{2l}.$
(Note that in this formula, the fact that $(\text{${s}$}_{1},\text{${s}$}_{2})\in Y_{d}(k)^{2m}$ means that
$\text{${s}$}_{1}=(s_{1,j})_{1\leqslant j\leqslant 2m},\quad\quad\text{${s}$}_{2}=(s_{2,j})_{1\leqslant j\leqslant 2m},$
with $s_{1,j}^{d}\not=s_{2,j}^{d}$ for all $j$.)
Thus we are reduced to estimating one variable sums, namely
$\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$},\text{${r}$},v;k)\quad\text{ and }\quad\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$}_{1},\text{${r}$},v;k)\overline{\mathbf{K}^{\prime}_{c}(\text{${s}$}_{2},\text{${r}$},v;k)}.$
These are expressions of “sums of products” type, hence we can use the Goursat–Kolchin–Ribet machinery, and in particular the new results of Section 3, to approach them.
Proposition 5.3.
With notation and assumptions as above, the following hold:
  1. (1)
    For all finite extensions $k$ of $\mathbf{F}_{q}$, the bound
    (5.5) $\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$},\text{${r}$},v;k)\ll\left\lvert k\right\rvert^{1/2}$
    holds for all but $O(\left\lvert k\right\rvert^{2m})$ values of $(\text{${r}$},\text{${s}$})\in k^{2m}\times(k^{\times})^{2m}$.
  2. (2)
    For all finite extensions $k$ of $\mathbf{F}_{q}$, the bound
    (5.6) $\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$}_{1},\text{${r}$},v;k)\overline{\mathbf{K}^{\prime}_{c}(\text{${s}$}_{2},\text{${r}$},v;k)}\ll\left\lvert k\right\rvert^{1/2}$
    holds for all but $O(\left\lvert k\right\rvert^{3m})$ values of $(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})\in k^{2m}\times Y_{d}(k)^{2m}$.
In both cases, the implied constants depend on $m$, $c$ and the complexity of $\mathscr{F}$, and depends also on $d$ in the second estimate.
Proof.
(1) The case of the estimate (5.5) follows immediately from Proposition 3.11 in view of the Grothendieck–Lefschetz trace formula and the Riemann Hypothesis (Theorem 2.2), since the function $v\mapsto K^{\prime}_{c}(\text{${r}$},\text{${s}$},v;k)$ is the trace function over $k$ of the sheaf $\mathscr{F}_{\text{${r}$},\text{${s}$},c}$ of (3.3); note that we use here our assumption that $\mathscr{F}$ is a middle-extension sheaf to ensure that the trace function of $\mathscr{F}^{\vee}$ is the complex conjugate of that of $\mathscr{F}$.
(2) Similarly, the sum
$\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$}_{1},\text{${r}$},v;k)\overline{\mathbf{K}^{\prime}_{c}(\text{${s}$}_{2},\text{${r}$},v;k)}.$
is the sum of the trace function of the sheaf
$\mathscr{F}_{\text{${s}$}_{1},\text{${r}$},c}\otimes\mathscr{F}_{\text{${s}$}_{2},\text{${r}$},c}^{\vee},$
and hence the estimate (5.6) follows similarly from Proposition 3.12 using the Grothendieck–Lefschetz trace formula and the Riemann Hypothesis (Theorem 2.2). ∎
We can now complete the proof of Proposition 5.1. Recall that the trivial bound
$\mathbf{K}^{\prime}_{c}(\text{${s}$},\text{${r}$},v;k)\ll 1$
also holds, where the implied constant depends on $m$ and the complexity of $\mathscr{F}$ (simply because the trace function of $\mathscr{F}$, which is mixed of weights $\leqslant 0$, is bounded in terms of the complexity of $\mathscr{F}$ only). Thus the one variable sums in Proposition 5.3 are always $\ll|k|$.
For (5.1), using (5.3), we obtain
$\sum_{\text{${v}$}\in k^{2l}}|\Sigma_{I}(\text{${v}$};k)|^{2m}$ $=\sum_{(\text{${r}$},\text{${s}$})\in(k\times{k^{\times}})^{2m}}\Bigl|\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$},\text{${r}$},v;k)\Bigr|^{2l}$
$\ll|k|^{2m+2l}+|k|^{4m+l},$
(the first term accounting for the “diagonal cases” and the second for the generic bound). Similarly for (5.2), using (5.4), we obtain
$\sum_{\text{${v}$}\in k^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$};k)|^{2m}$ $=\sum_{\text{${r}$}\in k^{2m}}\sum_{(\text{${s}$}_{1},\text{${s}$}_{2})\in Y_{d}(k)^{2m}}\Bigl|\sum_{v\in k}\mathbf{K}^{\prime}_{c}(\text{${s}$}_{1},\text{${r}$},v;k)\overline{\mathbf{K}^{\prime}_{c}(\text{${s}$}_{2},\text{${r}$},v;k)}\Bigr|^{2l}$
$\ll|k|^{3m+2l}+|k|^{6m+l},$
as claimed, with implied constants depending only on $m$ and the complexity of $\mathscr{F}$.
6. Stratification for gallant sheaves and conclusion of the proof
We will now complete the proof of Theorem 1.3. We first remark that it is enough to prove it when $K$ is the trace function of a sheaf $\mathscr{F}$ which is a gallant and light middle-extension sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$. This follows from the simple lemma below applied to $\mathscr{F}$ and $\mathscr{F}^{*}=j_{*}j^{*}\mathscr{F}$ for some open immersion $j\colon U\to\mathbf{A}^{1}$ of a dense open subset on which $\mathscr{F}$ is lisse.
Lemma 6.1.
Let $\mathscr{F}_{1}$ and $\mathscr{F}_{2}$ be gallant and light sheaves on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$ whose trace functions $K_{1}$ and $K_{2}$ coincide outside of a set $S\subset\mathbf{F}_{q}$. Let $b$ and $c$ be positive integers coprime to $q$. We then have
$\sum_{m}\sum_{n}\alpha_{m}\beta_{n}K_{1}(m^{b}n^{c})-\sum_{m}\sum_{n}\alpha_{m}\beta_{n}K_{2}(m^{b}n^{c})\ll\|\alpha\|_{2}\,\|\beta\|_{2},$
where the implied constant depends only on $b$, $c$, the complexities of $\mathscr{F}_{1}$ and $\mathscr{F}_{2}$, and the size of $S$.
Proof.
The difference is bounded by
$(\|K_{1}\|_{\infty}+\|K_{2}\|_{\infty})\sum_{s\in S}\sum_{\begin{subarray}{c}m\sim M,n\sim N\\ m^{b}n^{c}=s\end{subarray}}|\alpha_{m}|\,|\beta_{n}|.$
We consider the sum over $m$ and $n$ for each $s\in S$ separately. For each element $x$ of $\mathbf{F}_{q}$ which is a $c$-th power, we fix arbitrarily a $c$-th root $\sqrt[c]{x}$. Then this sum is equal to
$\sum_{\zeta^{c}=1}\sum_{\begin{subarray}{c}m\sim M\\ s\overline{m}^{b}\in(\mathbf{F}^{\times}_{q})^{c}\end{subarray}}|\alpha_{m}|\,|\beta_{\zeta\sqrt[c]{s\overline{m}^{b}}}|\leqslant\sum_{\zeta^{c}=1}\Bigl(\sum_{\begin{subarray}{c}m\sim M\\ s\overline{m}^{b}\in(\mathbf{F}^{\times}_{q})^{c}\end{subarray}}|\alpha_{m}|^{2}\Bigr)^{1/2}\Bigl(\sum_{\begin{subarray}{c}m\sim M\\ s\overline{m}^{b}\in(\mathbf{F}^{\times}_{q})^{c}\end{subarray}}|\beta_{\zeta\sqrt[c]{s\overline{m}^{b}}}|^{2}\Bigr)^{1/2},$
where we extend $\beta$ by zero outside of the values $n\sim N$.
By positivity, the first sum over $m$ is $\leqslant\|\alpha\|_{2}^{2}$. For the second sum, we note that for each $n\sim N$, there are at most $b$ values of $m$ such that $\zeta\sqrt[c]{s\overline{m}^{b}}=n$, hence this sum is $\leqslant b\|\beta\|_{2}^{2}$. The result follows. ∎
We now check that the stratification statement of Proposition 4.4 holds for the trace function of a gallant and light middle-extension sheaf.
Proposition 6.2.
Let $q$ be a prime number and let $\mathscr{F}$ be a gallant and light middle-extension sheaf on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$. Let $K\colon\mathbf{F}_{q}\to\mathbf{C}$ be the trace function of $\mathscr{F}$. Let $d\geqslant 1$ be the order of the finite group $T$ of Proposition 3.6 applied to $\mathscr{F}$.
Let $l\geqslant 1$ be an integer, and define $\Sigma_{I}(\text{${v}$})$ and $\Sigma_{II}^{(d)}(\text{${v}$})$ for $\text{${v}$}\in\mathbf{F}_{q}^{2l}$ as in Section 4.1.
There exists an integer $C\geqslant 1$, depending only on $l$, $c$ and the complexity of $\mathscr{F}$ such that the following properties hold if $q$ is large enough:
  1. (1)
    There exist algebraic varieties $\mathscr{V}_{1}$ and $\mathscr{V}_{1}^{\Delta}$ over $\mathbf{F}_{q}$ such that
    $\mathbf{A}^{2l}_{\mathbf{F}_{q}}\supset\mathscr{V}_{1}\supset\mathscr{V}_{1}^{\Delta}$
    $\dim(\mathscr{V}_{1})\leqslant l+\Bigl\lceil\frac{l}{2}\Bigr\rceil,\quad\dim(\mathscr{V}^{\Delta}_{1})\leqslant l,$
    $\max(\deg(\mathscr{V}_{1}),\deg(\mathscr{V}_{1}^{\Delta}))\leqslant C$
    and
    $|\Sigma_{I}(\text{${v}$})|\leqslant\begin{cases}Cq&\text{ for }\text{${v}$}\not\in\mathscr{V}_{1}(\mathbf{F}_{q}),\\ Cq^{3/2}&\text{ for }\text{${v}$}\not\in\mathscr{V}^{\Delta}_{1}(\mathbf{F}_{q}),\\ Cq^{2}&\text{ for }\text{${v}$}\in\mathscr{V}_{1}^{\Delta}(\mathbf{F}_{q}).\end{cases}$
  2. (2)
    There exist algebraic varieties $\mathscr{V}_{2}$ and $\mathscr{V}_{2}^{\Delta}$ over $\mathbf{F}_{q}$ such that
    $\mathbf{A}^{2l}_{\mathbf{F}_{q}}\supset\mathscr{V}_{2}\supset\mathscr{V}_{2}^{\Delta}$
    $\dim(\mathscr{V}_{2})\leqslant l+2\Bigl\lceil\frac{l}{3}\Bigr\rceil,\quad\dim(\mathscr{V}^{\Delta}_{2})\leqslant l,$
    $\max(\deg(\mathscr{V}_{2}),\deg(\mathscr{V}_{2}^{\Delta}))\leqslant C,$
    and
    $|\Sigma_{II}^{(d)}(\text{${v}$})|\leqslant\begin{cases}Cq^{3/2}&\text{ for }\text{${v}$}\not\in\mathscr{V}_{2}(\mathbf{F}_{q}),\\ Cq^{2}&\text{ for }\text{${v}$}\not\in\mathscr{V}^{\Delta}_{2}(\mathbf{F}_{q}),\\ Cq^{3}&\text{ for }\text{${v}$}\in\mathscr{V}_{2}^{\Delta}(\mathbf{F}_{q}).\end{cases}$
Once this Proposition is proven we can conclude the proof of Theorem 1.3. Indeed, we already reduced to the case of a middle-extension sheaf. Proposition 6.2 shows that the trace function of a gallant and light middle-extension sheaf $\mathscr{F}$ satisfies the conditions of Proposition 4.4. In the case of Type II sums, we then combine the resulting estimate
$\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\ll q^{9/2+2/l}$
for $V=q^{3/(2l)}$ ((4.1) is satisfied by (1.2)) with the bound (4.3) from Proposition 4.1, applied with $d$ the order of the subgroup $T$, to deduce
$\sum_{m\sim M}\sum_{n\sim N}\alpha_{n}\beta_{n}K(m^{b}n^{c})$ $\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}(MN)^{{\textstyle{\frac{1}{2}}}}\Bigl(\frac{1}{M}+\Bigl(\frac{1}{M^{2}N^{2}V^{2l-1}}\sum_{\text{${v}$}\in[V,2V]^{2l}}|\Sigma_{II}^{(d)}(\text{${v}$})|\Bigr)^{\tfrac{1}{2l}}\Bigr)^{{\textstyle{\frac{1}{2}}}}$
$\ll q^{\varepsilon}\|\text{${\alpha}$}\|_{2}\|\text{${\beta}$}\|_{2}(MN)^{{\textstyle{\frac{1}{2}}}}\Bigl(\frac{1}{M}+\Bigl(\frac{1}{MN}\Bigr)^{\tfrac{1}{l}}\Bigl(q^{\tfrac{9}{2}+\tfrac{2}{l}-\tfrac{3(2l-1)}{2l}}\Bigr)^{\tfrac{1}{2l}}\Bigr)^{{\textstyle{\frac{1}{2}}}},$
which is (1.3). A similar argument deduces (1.1) from (4.2).
Proof.
(1) Let $p_{12}\colon(\mathbf{A}^{1}\times\mathbf{G}_{m}\times\mathbf{A}^{2l})_{\mathbf{F}_{q}}\to\mathbf{A}^{2l}_{\mathbf{F}_{q}}$ be the projection $(r,s,\text{${v}$})\mapsto\text{${v}$}$. For $1\leqslant i\leqslant 2l$, let
$g_{i}\colon(\mathbf{A}^{1}\times\mathbf{G}_{m}\times\mathbf{A}^{2l})_{\mathbf{F}_{q}}\to\mathbf{A}^{1},$
be the morphism defined by
$g_{i}(r,s,\text{${v}$})=s(r+v_{i})^{c}.$
Let $M$ be the mixed complex on $\mathbf{A}^{2l}_{\mathbf{F}_{q}}$ defined by
$M=Rp_{12!}\Bigl(\bigotimes_{i=1}^{l}g_{i}^{*}\mathscr{F}\otimes g_{i+l}^{*}\mathscr{F}^{\vee}\Bigr),$
(where $\mathscr{F}$ and $\mathscr{F}^{\vee}$ are viewed as complexes in degree 0). By the Grothendieck–Lefschetz trace formula, the complex $M$ satisfies
$t_{M}(\text{${v}$};k)=\Sigma_{I}(\text{${v}$};k)$
for any finite extension $k$ of $\mathbf{F}_{q}$ and any $\text{${v}$}\in k^{2l}$. By the Riemann Hypothesis (in the form of [WeilII, Th. I]), the complex $M$ is mixed of integral weights $\leqslant 0$ since we assumed that $\mathscr{F}$ itself is mixed of integral weights $\leqslant 0$.
We apply Theorem 2.4 to $X=\mathbf{A}^{2l}$ and to the complex $M$. We let $m=\lceil\frac{l}{2}\rceil$ and put $A=4\lceil\frac{l}{2}\rceil+l$. Proposition 5.1 gives
$\sum_{\text{${v}$}\in k^{2l}}\left\lvert\Sigma_{I}(\text{${v}$};k)\right\rvert^{2m}\ll\left\lvert k\right\rvert^{2\lceil\frac{l}{2}\rceil+2l}+\left\lvert k\right\rvert^{4\lceil\frac{l}{2}\rceil+l}\ll 2\left\lvert k\right\rvert^{A}$
for any finite extension $k$ of $\mathbf{F}_{q}$. Theorem 2.4 implies that there exist closed subschemes $X^{(3)}\supset X^{(4)}$ with
$\dim(X^{(3)})\leqslant A-3m=\Bigl\lceil\frac{l}{2}\Bigr\rceil+l,\quad\quad\dim(X^{(4)})\leqslant l,$
such that
$\Sigma_{I}(\text{${v}$};k)$ $\ll\left\lvert k\right\rvert^{\frac{3-1}{2}}=|k|\text{ for }\text{${v}$}\not\in X^{(3)}(k),$
$\Sigma_{I}(\text{${v}$};k)$ $\ll\left\lvert k\right\rvert^{3/2}\text{ for }\text{${v}$}\not\in X^{(4)}(k).$
This gives the desired stratification with $\mathscr{V}_{1}^{\Delta}=X^{(4)}$ and $\mathscr{V}_{1}=X^{(3)}$ .
(2) We recall that $Y_{d}$ is the subvariety of $\mathbf{G}_{m}\times\mathbf{G}_{m}$ defined by the equation $s_{1}^{d}\not=s_{2}^{d}$. Let $p_{123}\colon(\mathbf{A}^{1}\times Y_{d}\times\mathbf{A}^{2l})_{\mathbf{F}_{q}}\to\mathbf{A}^{2l}_{\mathbf{F}_{q}}$ be the projection $(r,s_{1},s_{2},\text{${v}$})\to\text{${v}$}$. For $1\leqslant i\leqslant 2l$, let $g_{i}$ and $h_{i}$ be the morphisms
$(\mathbf{A}^{1}\times Y_{d}\times\mathbf{A}^{2l})_{\mathbf{F}_{q}}\to\mathbf{A}^{1},$
defined by
$g_{i}(r,s_{1},s_{2},\text{${v}$})=s_{1}(r+v_{i})^{c},\quad\quad h_{i}(r,s_{1},s_{2},\text{${v}$})=s_{2}(r+v_{i})^{c}.$
Let $M$ be the mixed complex on $\mathbf{A}^{2l}_{\mathbf{F}_{q}}$ defined by
$M=Rp_{123!}\Bigl(\Bigl(\bigotimes_{i=1}^{l}g_{i}^{*}\mathscr{F}\otimes g_{i+l}^{*}\mathscr{F}^{\vee}\Bigr)\otimes\Bigl(\bigotimes_{i=1}^{l}h_{i}^{*}\mathscr{F}^{\vee}\otimes h_{i+l}^{*}\mathscr{F}\Bigr)\Bigr).$
By the Lefschetz trace formula, the complex $M$ satisfies $t_{M}(\text{${v}$};k)=\Sigma_{II}(\text{${v}$};k)$ for all $\text{${v}$}\in k^{2l}$, and by the Riemann Hypothesis [WeilII, Th. I], it is mixed of integral weights $\leqslant 0$.
We apply Theorem 2.4 to $X=\mathbf{A}^{2l}$ and to $M$. For $m=\lceil\frac{l}{3}\rceil$ and $A=6\lceil\frac{l}{3}\rceil+l$, we have
$\sum_{\text{${v}$}\in\mathbf{A}^{2l}(k)}\left\lvert\Sigma_{II}(K,\text{${v}$};k)\right\rvert^{2m}\ll\left\lvert k\right\rvert^{3m+2l}+\left\lvert k\right\rvert^{6m+l}\ll\left\lvert k\right\rvert^{A}$
for any finite extension $k/\mathbf{F}_{q}$ by Proposition 5.1. Hence we obtain closed subschemes
$X^{(3)}\supset X^{(4)}\supset X^{(5)}\supset X^{(6)}$
such that
$\dim(X^{(w)})$ $\leqslant A-mw=(6-w)\Bigl\lceil\frac{l}{3}\Bigr\rceil+l,$
$\Sigma_{II}(\text{${v}$};k)$ $\ll\left\lvert k\right\rvert^{\frac{w-1}{2}}\quad\text{ for }\text{${v}$}\not\in X^{(w)}(k).$
We define $\mathscr{V}_{2}=X^{(4)}$, so that $\dim(\mathscr{V}_{2})\leqslant l+2\lceil\frac{l}{3}\rceil$ and
$\Sigma_{II}(\text{${v}$};k)\ll\left\lvert k\right\rvert^{3/2}\quad\text{ for }\text{${v}$}\not\in\mathscr{V}_{2}(k),$
which corresponds to the first part of our claim.
However, neither $X^{(5)}$ nor $X^{(6)}$ is suitable as a choice of $\mathscr{V}_{2}^{\Delta}$ (the former’s codimension is too large, and the bound given outside of the latter is of size $|k|^{5/2}$ instead of our goal of $|k|^{2}$). To define $\mathscr{V}_{2}^{\Delta}$, we use instead a different trick.
Let $k$ be a finite extension of $\mathbf{F}_{q}$. We observe that
$\Sigma_{II}(\text{${v}$};k)$ $=\sum_{r\in k}\Bigl|\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\Bigr|^{2}-\sum_{r\in k}\sum_{\begin{subarray}{c}s_{1},s_{2}\in k^{\times}\\ s_{1}^{d}=s_{2}^{d}\end{subarray}}\mathbf{K}(r,s_{1},\text{${v}$};k)\overline{\mathbf{K}(r,s_{2},\text{${v}$};k)}$
$=\sum_{r\in k}\Bigl|\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\Bigr|^{2}-\sum_{\begin{subarray}{c}\xi\in k^{\times}\\ \xi^{d}=1\end{subarray}}\sum_{r\in k}\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\overline{\mathbf{K}(r,\xi s,\text{${v}$};k)}.$
Since $\mathscr{F}$ is mixed of weights $\leqslant 0$, we have
$\sum_{\xi^{d}=1}\sum_{r\in k}\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\overline{\mathbf{K}(r,\xi s,\text{${v}$};k)}\ll|k|^{2}$
for all $\text{${v}$}\in k^{2l}$, and it is therefore enough to prove the existence of a subvariety $\mathscr{V}_{2}^{\Delta}$ over $\mathbf{F}_{q}$ of dimension $\leqslant l$ and bounded degree such that
$\sum_{r\in k}\Bigl|\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\Bigr|^{2}\ll|k|^{2}$
for $\text{${v}$}\notin\mathscr{V}_{2}^{\Delta}(k)$. This follows from Proposition 3.13. Indeed, let $\mathscr{V}^{\Delta}_{2}$ be the algebraic variety provided by this proposition applied to $\mathscr{F}$. Let $\text{${v}$}\notin\mathscr{V}^{\Delta}_{2}(k)$. For all but a bounded number of $r\in k$, we have
$H^{2}_{c}\Bigl(\mathbf{A}^{1}_{\bar{k}},\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i})^{c}s]^{*}\mathscr{F}\otimes\bigotimes_{i=1}^{l}[s\mapsto(r+v_{i+l})^{c}s]^{*}\mathscr{F}^{\vee}\Bigr)=0$
by the proposition, hence
$\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\ll|k|^{1/2}$
by the Riemann Hypothesis (Theorem 2.2). For the possible exceptional values of $r$, we have
$\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\ll|k|$
by the trivial bound. Hence, we deduce that
$\sum_{r\in k}\Bigl|\sum_{s\in k^{\times}}\mathbf{K}(r,s,\text{${v}$};k)\Bigr|^{2}\ll|k|^{2}$
for all $\text{${v}$}\notin\mathscr{V}^{\Delta}_{2}(k)$, as desired. ∎
7. Bounds for trilinear sums with monomial arguments
In this section, we prove Theorem 1.4, and we use the notation from that statement. As in the proof of Theorem 1.3 in the previous section, we check first that we can assume that $\mathscr{F}$ is also a middle-extension sheaf. Viewing the integers $a$, $b$ and $c$, as well as the trace function $K$ of the gallant sheaf $\mathscr{F}$ as fixed, we denote simply
$T(\text{${\alpha}$},\text{${\beta}$},\text{${\gamma}$})=\mathop{\sum\sum\sum}\limits_{j\sim J,m\sim M,n\sim N}\alpha_{j}\beta_{m}\gamma_{n}K(j^{a}m^{b}n^{c})$
for families $\text{${\alpha}$}=(\alpha_{j})_{j\sim J}$, $\text{${\beta}$}=(\beta_{m})_{m\sim M}$ and $\text{${\gamma}$}=(\gamma_{n})_{n\sim N}$ of complex numbers. We assume that
$|\alpha_{j}|\leqslant 1,\quad|\beta_{m}|\leqslant 1,\quad|\gamma_{n}|\leqslant 1$
for all $j$, $m$ and $n$.
We write
$T(\text{${\alpha}$},\text{${\beta}$},\text{${\gamma}$})=\mathop{\sum\sum}\limits_{u,v\in\mathbf{F}^{\times}_{q}}\xi_{u}\zeta_{v}K(uv)$
where
$\xi_{u}=\sum_{\begin{subarray}{c}j\sim J\\ j^{a}\equiv u\,(\mathrm{mod}\,{q})\end{subarray}}\alpha_{j},\quad\quad\zeta_{v}=\mathop{\sum\sum}\limits_{\stackrel{{\scriptstyle{\scriptstyle m\sim M,n\sim N}}}{{{\scriptstyle m^{b}n^{c}\equiv v\,(\mathrm{mod}\,{q})}}}}\beta_{m}\gamma_{n}.$
We have the following simple estimates for the size of these coefficients.
Lemma 7.1.
We have
$\|\text{${\xi}$}\|_{1}\ll J,\quad\quad\|\text{${\xi}$}\|^{2}_{2}\ll J\Bigl(\frac{J}{q}+1\Bigr),$
$\|\text{${\zeta}$}\|_{1}\ll MN,\quad\quad\|\text{${\zeta}$}\|_{2}^{2}\ll\frac{(MN)^{2}}{q}+MN(\log q)^{2},$
where the implied constants depend on $a$, $b$ and $c$.
Proof.
The bounds for $\|\text{${\xi}$}\|_{1}$ and $\|\text{${\zeta}$}\|_{1}$ are clear.
Next, using $|\alpha_{j}|\leqslant 1$, we have
$\|\text{${\xi}$}\|_{2}^{2}=\sum_{u\in\mathbf{F}^{\times}_{q}}|\xi_{u}|^{2}\leqslant\sum_{\stackrel{{\scriptstyle{\scriptstyle j_{1}\sim J,\ j_{2}\sim J}}}{{{\scriptstyle j_{1}^{a}=j_{2}^{a}\,(\mathrm{mod}\,{q})}}}}1\ll L(L/q+1).$
Finally, we use multiplicative characters for the last estimate. We have
$\|\text{${\zeta}$}\|^{2}_{2}=\sum_{v}|\zeta_{v}|^{2}$ $\leqslant\sum_{\begin{subarray}{c}m_{1},m_{2},n_{1},n_{2}\\ m_{i}\sim M,\,n_{i}\sim N\\ m_{1}^{b}n_{1}^{c}\equiv m_{2}^{b}n_{2}^{c}\,(\mathrm{mod}\,{q})\end{subarray}}1$
$=\frac{1}{q-1}\sum_{\chi\,(\mathrm{mod}\,{q})}\Bigl|\sum_{m\sim M,n\sim N}\chi^{b}(m)\chi^{c}(n)\Bigr|^{2}$
$\ll\frac{(MN)^{2}}{q}+\frac{1}{q}\sum_{\begin{subarray}{c}\chi\\ \chi^{b},\chi^{c}\not=1\end{subarray}}\Bigl|\sum_{m,n}\chi^{b}(m)\chi^{c}(n)\Bigr|^{2},$
where the sums over $\chi$ are over characters of $\mathbf{F}_{q}^{\times}$.
By the Cauchy-Schwarz inequality, we deduce
$\|\text{${\zeta}$}\|^{2}_{2}\ll\frac{(MN)^{2}}{q}+\Bigl(\frac{1}{q-1}\sum_{\chi^{b}\not=1}|\sum_{m\sim M}\chi^{b}(m)|^{4}\Bigr)^{1/2}\Bigl(\frac{1}{q-1}\sum_{\chi^{c}\not=1}|\sum_{n\sim N}\chi^{c}(n)|^{4}\Bigr)^{1/2}.$
By positivity, note that
$\frac{1}{q-1}\sum_{\chi^{b}\not=1}\Bigl|\sum_{m\sim M}\chi^{b}(m)\Bigr|^{4}\leqslant\frac{1}{q-1}\sum_{\chi\not=1}\Bigl|\sum_{m\sim M}\chi(m)\Bigr|^{4}\ll M^{2}(\log q)^{2},$
where the last inequality is a result of Ayyad, Cochrane and Zheng [Ayyad]*Thm. 2. Arguing similarly for the sum over characters with $\chi^{c}\not=1$, we conclude that
$\|\text{${\zeta}$}\|^{2}_{2}\ll\frac{(MN)^{2}}{q}+MN(\log q)^{2},$
which finishes the proof. ∎
Let $l\geqslant 2$ be an integer. We write
$|\zeta_{v}|=|\zeta_{v}|^{(l-1)/l}|\zeta_{v}|^{1/l}$
and apply Hölder’s inequality followed by the Cauchy-Schwarz inequality, leading to the inequality
$\Bigl|\mathop{\sum\sum}\limits_{u,v\in\mathbf{F}^{\times}_{q}}\xi_{u}\zeta_{v}K(uv)\Bigr|$ $\leqslant\|\text{${\zeta}$}\|_{1}^{(l-1)/l}\Bigl(\sum_{v\in\mathbf{F}^{\times}_{q}}|\zeta_{v}|\,\Bigl|\sum_{u\in\mathbf{F}^{\times}_{q}}\xi_{u}K(uv)\Bigr|^{l}\Bigr)^{1/l}$
(7.1) $\leqslant\|\text{${\zeta}$}\|_{1}^{1-1/l}\|\text{${\zeta}$}\|_{2}^{1/l}\Bigl(\sum_{v\in\mathbf{F}^{\times}_{q}}\Bigl|\sum_{u\in\mathbf{F}^{\times}_{q}}\xi_{u}K(uv)\Bigl|^{2l}\Bigr)^{1/(2l)}.$
Combined with Lemma 7.1, this implies
$\Bigl|\mathop{\sum\sum}\limits_{u,v\in\mathbf{F}^{\times}_{q}}\xi_{u}\zeta_{v}K(uv)\Bigr|\ll(\log q)^{2/l}(MN)^{1-1/(2l)}\Bigl(\sum_{v\in\mathbf{F}^{\times}_{q}}\Bigl|\sum_{u\in\mathbf{F}^{\times}_{q}}\xi_{u}K(uv)\Bigr|^{2l}\Bigr)^{1/(2l)}.$
We then expand the $(2l)$-th power to write
$\sum_{v\in\mathbf{F}^{\times}_{q}}\Bigl|\sum_{u\in\mathbf{F}^{\times}_{q}}\xi_{u}K(uv)\Bigr|^{2l}=\sum_{\text{${u}$}\in(\mathbf{F}^{\times}_{q})^{2l}}\xi_{\text{${u}$}}\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}$
where we denote
$\text{${u}$}=(u_{1},\cdots,u_{2l}),\quad\quad\xi_{\text{${u}$}}=\prod_{i=1}^{l}\xi_{u_{i}}\overline{\xi_{u_{i+l}}}$
for $\text{${u}$}\in(\mathbf{F}^{\times}_{q})^{2l}$.
Since $\mathscr{F}$ is mixed of weights $\leqslant 0$, we have the bound
$\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}\ll q$
for all $\text{${u}$}\in(\mathbf{F}_{q}^{\times})^{2l}$. Moreover, since $\mathscr{F}$ is gallant and light, and a middle-extension, it follows from Corollary 3.14 that there exists an algebraic subvariety $\mathscr{V}$ of $\mathbf{A}^{2k}_{\mathbf{F}_{q}}$ of dimension $\leqslant l$ and degree bounded in terms of $l$ and the complexity of $\mathscr{F}$ such that the estimate
$\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}\ll q^{1/2}$
[…]
We next observe that
$\sum_{\text{${u}$}\in\mathscr{V}(\mathbf{F}_{q})}|\xi_{\text{${u}$}}|=\sum_{\text{${u}$}\in\mathscr{V}(\mathbf{F}_{q})}\Bigl|\prod_{i=1}^{l}\xi_{u_{i}}\overline{\xi_{u_{i+l}}}\Bigr|\leqslant\sum_{\text{${u}$}\in\mathscr{V}(\mathbf{F}_{q})}\prod_{i=1}^{2l}\sum_{\begin{subarray}{c}j\sim J\\ j^{a}=u_{i}\end{subarray}}1\leqslant a^{2l}\sum_{\begin{subarray}{c}\text{${j}$}\in[J,2J]^{2l}\\ f(\text{${j}$})\in\mathscr{V}(\mathbf{F}_{q})\end{subarray}}1,$
where we define the morphism $f\colon\mathbf{G}_{m}^{2l}\to\mathbf{G}_{m}^{2l}$ by
$f(x_{1},\ldots,x_{2l})=(x_{1}^{a},\ldots,x_{2l}^{a}).$
Applying Lemma 2.1 to $f^{-1}(\mathscr{V})$, which also has dimension $\leqslant l$, we deduce that
$\sum_{\text{${u}$}\in\mathscr{V}(\mathbf{F}_{q})}|\xi_{\text{${u}$}}|\ll J^{l},$
and hence
$\sum_{\text{${u}$}\in(\mathbf{F}^{\times}_{q})^{2l}}\xi_{\text{${u}$}}\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}$ $\ll q\sum_{\text{${u}$}\in\mathscr{V}(\mathbf{F}_{q})}|\xi_{\text{${u}$}}|+\sum_{\text{${u}$}\not\in\mathscr{V}(\mathbf{F}_{q})}\Bigl|\sum_{v\in\mathbf{F}^{\times}_{q}}\prod_{i=1}^{l}K(u_{i}v)\overline{K(u_{i+l}v)}\Bigr|$
$\ll J^{l}q+J^{2l}q^{1/2}.$
Combining the previous estimates, we obtain the inequality
$T(\text{${\alpha}$},\text{${\beta}$},\text{${\gamma}$})\ll(\log q)^{1/l}(MN)^{1-1/(2l)}(J^{2l}q^{1/2}+J^{l}q)^{1/(2l)}\ll q^{\varepsilon}JMN\Bigl(\frac{q^{1/2}}{MN}+\frac{q}{J^{l}MN}\Bigr)^{1/(2l)},$
for any $\varepsilon>0$, which concludes the proof of Theorem 1.4.
8. The oxozonic case
In this short section we explain how to adapt the previous proofs to establish Theorem 1.6. Thus let $a$, $b$, $c$ be non-zero integers, let $q$ be a prime and let $\mathscr{F}$ be an oxozonic sheaf, mixed of weights $\leqslant 0$, on $\mathbf{A}^{1}_{\mathbf{F}_{q}}$.
First, assuming that $c$ is odd (which implies that the geometric monodromy of $x\mapsto x^{c}]^{*}\mathscr{F}$ is still $\mathbf{O}_{4}$), we can follow the proof of Theorem 1.3 using simply Proposition 3.9 (see also Remark 3.3) in place of Proposition 3.6 to establish the analogue of Proposition 5.1. Similarlly, the proof of Theorem 1.4 also applies verbatim, in all cases here since only the sheaf $\mathscr{F}$ appears in this argument.
9. Ubiquity of gallant sheaves
This section provides a large sample of examples of gallant sheaves; this will illustrate that this class is much wider, and much more flexible than the restricted types of sheaves allowed in the previous papers such as [KMSAnn] and [Pisa]. In particular, Theorem 1.3 should have many applications in the future.555 We will not discuss systematically the weights of the sheaves which appear; checking that they satisfy the additional weight conditions is usually much more straightforward, using the Riemann Hypothesis.
There are two basic principles involved. The first one is that Katz has computed the geometric monodromy group of a very wide variety of families of exponential sums. These give many examples of trace functions, and an “experimental” fact is that in many cases the (connected component of the identity of this) group is a simple algebraic group, which implies by definition that the corresponding sheaf is gallant. The second general principle is that the connected component of the geometric monodromy group of a sheaf is a “robust” invariant, in the sense that if a sheaf is transformed in certain ways, then this group remains the same. Thus, any such operation will transform a gallant sheaf into another one.
We begin by describing some of these transformations. Throughout, we denote by $k$ a finite field and by $\mathscr{F}$ a constructible $\ell$-adic sheaf on $\mathbf{A}^{1}$ over $k$, for some prime $\ell$ invertible in $k$. We denote by $G$ the geometric monodromy group of $\mathscr{F}$, and by $G^{0}$ its connected component of the identity.
9.1. Bountiful sheaves
In the paper [sumproducts], Fouvry, Kowalski and Michel presented in a concrete way the application of the Goursat–Kolchin–Ribet criterion of Katz in the context of estimating sums of products of trace functions of the form
$\sum_{x\in\mathbf{F}_{q}}K(\gamma_{1}\cdot x)\cdots K(\gamma_{r}\cdot x)e\Bigl(\frac{hx}{q}\Bigr)$
for some trace function $K$ and $\gamma_{i}\in\mathbf{GL}_{2}(\mathbf{F}_{q})$ acting as fractional linear transformations. To encapsulate this method, they defined bountiful sheaves (see [sumproducts, Def. 1.2]). In particular, these sheaves have geometric monodromy group equal to either $\mathbf{SL}_{r}$ or $\mathbf{Sp}_{r}$ for some integer $r\geqslant 2$; since these are simple algebraic groups, and the definition also requires that $\mathscr{F}$ is mixed of weights $\leqslant 0$ and pure of weight 0 on a dense open set, it follows immediately that any bountiful sheaf is gallant and light.
9.2. Birationality
Since the definition of the geometric monodromy group of a sheaf only depends on the restriction of this sheaf to an arbitrary open dense subset of $\mathbf{A}^{1}_{k}$, it follows that for any open immersion $j\colon U\to\mathbf{A}^{1}_{k}$ with $U$ not empty, the sheaves $j_{!}j^{*}\mathscr{F}$ and $j_{*}j^{*}\mathscr{F}$ are gallant if and only if $\mathscr{F}$ is gallant.
9.3. Forms of exponential sums
By definition, whether $\mathscr{F}$ is gallant or not only depends on its geometric monodromy group, and hence only depends on the base change of $\mathscr{F}$ to an algebraic closure $\bar{k}$ of the base field $k$. Using standard terminology from algebraic geometry, two objects defined over $k$ which become isomorphic over $\bar{k}$ are called forms of each other, and in the context of trace functions, one could speak of forms of exponential sums.
Here is an example to illustrate the fact that this notion can be far from trivial. Let $r\geqslant 1$ be an integer and let $\text{${a}$}=(a_{1},\ldots,a_{r})$ be a family of positive integers. For any prime number $q$ and any $v\in\mathbf{F}_{q}^{\times}$, we may define the exponential sums
(9.1) $\widetilde{H}_{\text{${a}$}}(v;q)=\frac{1}{q^{(r-1)/2}}\sum_{\begin{subarray}{c}x\in(\mathbf{F}_{q}^{\times})^{r}\\ x_{1}^{a_{1}}\cdots x_{r}^{a_{r}}=v\end{subarray}}e\Bigl(\frac{x_{1}+\cdots+x_{r}}{q}\Bigr).$
We claim that this is a “form” of the hyper-Kloosterman sums with characters defined as follows: we let
$a=a_{1}+\cdots+a_{r},$
and we let
$\text{${\chi}$}=(\chi_{1},\ldots,\chi_{a})$
be an arbitrary ordering of the $\ell$-adic characters $\chi$ of $\mathbf{F}^{\times}_{q}$ such that $\chi^{a_{i}}=1$ for some $i$, repeated with multiplicity (so that there are indeed $a$ such characters). Then the corresponding hyper-Kloosterman sums are defined by
(9.2) $\operatorname{\mathrm{Kl}}_{a}(u,\text{${\chi}$};q)=\frac{1}{q^{(a-1)/2}}\sum_{\begin{subarray}{c}y_{1},\ldots,y_{a}\in\mathbf{F}_{q}^{\times}\\ y_{1}\cdots y_{a}=u\end{subarray}}\prod_{j=1}^{a}\chi_{j}(y_{j})e\Bigl(\frac{y_{1}+\cdots+y_{a}}{q}\Bigr)$
for $u\in\mathbf{F}^{\times}_{q}$.
Precisely, the claim is that the natural hypergeometric complexes over $\mathbf{F}_{q}$ (in the sense of Katz) with trace functions (9.1) and (9.2)666 Actually, a multiplicative translate of the latter. are geometrically isomorphic. This is a non-trivial fact, which ultimately depends on a version of the Hasse–Davenport relations (as in [GKM, p. 84]). One could then deduce a version of Theorem 1.3 for the sums $\widetilde{H}_{\text{${a}$}}$ from the result of [Pisa], which apply to hyper-Kloosterman sums. However, we will not elaborate on this, since we will see in Section 9.6 that “most” hypergeometric sheaves are gallant.
9.4. Change of variable
Let $j\colon\mathbf{A}^{1}_{k}\to\mathbf{P}^{1}_{k}$ denote the open immersion. It is essentially obvious that for any automorphism $\varphi$ of $\mathbf{P}^{1}_{k}$, the pullback sheaves $\varphi^{*}(j_{*}\mathscr{F}$) and $\varphi^{*}(j_{!}\mathscr{F})$ are gallant if and only if $\mathscr{F}$ itself is gallant.
Concretely, this means that if Theorem 1.1 applies to a trace function $K$ modulo a prime $q$, then it also applies to the function
$x\mapsto K\Bigl(\frac{ax+b}{cx+d}\Bigr),$
for any $\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ in $\mathbf{GL}_{2}(\mathbf{F}_{q})$, with the convention that $K(\infty)=0$. This is already quite significant: even the case of a translation applied to a Kloosterman sheaf, i.e., the case of the trace function
$a\mapsto\frac{1}{p^{(r-1)/2}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{r}\in\mathbf{F}_{q}\\ x_{1}\cdots x_{r}=a+h\end{subarray}}e\Bigl(\frac{x_{1}+\cdots+x_{r}}{q}\Bigr),\quad\quad h\in\mathbf{F}_{q},$
was not previously known for (fixed) $h\not=0$.
More generally, if the geometric monodromy group $G$ of $\mathscr{F}$ is infinite and $\mathscr{F}$ is gallant, then for any non-constant morphism $f\colon\mathbf{P}^{1}_{k}\to\mathbf{P}^{1}_{k}$, the sheaves $j^{*}f^{*}j_{*}\mathscr{F}$ or $j^{*}f^{*}j_{!}\mathscr{F}$ are gallant. Indeed, it is known in general that the connected component of the geometric monodromy group of these will be finite-index subgroups of $G^{0}$, so must coincide with it since the latter is simple by definition.
This corresponds concretely to a change of variable of the form
$x\mapsto K(f(x)),\quad\quad K(\infty)=0,$
for an arbitrary non-constant $f\in\mathbf{F}_{q}(X)$.
If $\mathscr{F}$ is gallant but has finite geometric monodromy group (the second case of Definition 1.2), then a restricted version of this principle applies: if $f\colon\mathbf{P}^{1}_{k}\to\mathbf{P}^{1}_{k}$ is a cyclic cover, then $j^{*}f^{*}j_{*}\mathscr{F}$ and $j^{*}f^{*}j_{!}\mathscr{F}$ are also gallant, because their geometric monodromy group will still contain (as a normal subgroup) a copy of the perfect subgroup $N$ of the definition.
In particular, if $\mathscr{F}$ is gallant then in any case, the sheaf $x\mapsto x^{c}]^{*}\mathscr{F}$ is gallant if $c$ is a non-zero integer. For trace functions, this corresponds to $x\mapsto K(x^{c})$.
9.5. Tannakian operations
Let $j\colon U\to\mathbf{A}^{1}_{k}$ be the open immersion of some open dense subset of $\mathbf{A}^{1}_{k}$ on which $\mathscr{F}$ is lisse. Let $G^{a}$ denote the arithmetic fundamental group of $j^{*}\mathscr{F}$. Let
$\varrho\colon G^{a}\to\mathbf{GL}(V)$
be a non-trivial finite-dimensional irreducible (continuous) representation of $G^{a}$ on a $\overline{\mathbf{Q}}_{\ell}$-vector space $V$. We can then form the sheaf $\varrho(j^{*}\mathscr{F})$ corresponding to the representation obtained by composing with $\varrho$ the representation associated to $j^{*}\mathscr{F}$.
Assume that $\mathscr{F}$ is gallant. If $G^{0}$ is infinite and the restriction of $\varrho$ to the subgroup $G^{0}$ is irreducible, or if $G$ is finite but the restriction to the normal subgroup $N$ of Definition 1.2 is irreducible, then the sheaves $j_{*}\varrho(j^{*}\mathscr{F})$ and $j_{!}\varrho(j^{*}\mathscr{F})$ are gallant. Indeed, their geometric monodromy groups coincide with the image of $G^{0}$ (resp. of $N$) under $\varrho$, but the restriction of $\varrho$ to $G^{0}$ is faithful in the infinite case, and otherwise the non-trivial image of the quasisimple group $N$ is quasisimple (see, e.g., [isaacs, Lemma 9.2]).
A concrete example here is the following. Consider the Kloosterman sheaf $\mathscr{F}$ modulo $q$ with trace function
$a\mapsto\operatorname{\mathrm{Kl}}_{2}(a;q)=\frac{1}{\sqrt{q}}\sum_{x\in\mathbf{F}_{q}^{\times}}e\Bigl(\frac{ax+\bar{x}}{q}\Bigr),$
for $a\in\mathbf{F}^{\times}_{q}$. Write $\operatorname{\mathrm{Kl}}_{2}(a;q)=2\cos(\theta(a;q))$ for some unique $\theta(a;q)\in[0,\pi]$. Then, because the group $G$ coincides with $G^{0}$ in that case and is the simple group $\mathbf{SL}_{2}$, we can apply any non-trivial irreducible representation of $\mathbf{SL}_{2}$. These representations are the symmetric powers of the standard representation, and this means that our results apply, for any integer $d\geqslant 1$, to the function $K^{(d)}$ defined by $K^{(d)}(0)=0$ and
$K^{(d)}(a)=\frac{\sin((d+1)\theta(a;q))}{\sin(\theta(a;q))}$
for $a\in\mathbf{F}_{q}^{\times}$. It would be interesting to see if such examples have concrete analytic applications.
9.6. Hypergeometric sheaves
In this section and the next, we discuss one of the most important class of examples of trace functions in the context of applications to analytic number theory, especially automorphic forms and $L$-functions. These are hypergeometric sums and their associated sheaves, as defined by Katz (see [GKM] and [ESDE]). It turns out, as we will explain, that many such sheaves are gallant. This can be derived from the basic facts established in [ESDE], but more recent work of Guralnick, Katz, Rojas–Léon and Tiep give more definitive statements, and we will quote from them.
Let $k$ be a finite field and $\ell$ a prime invertible in $k$. Let $\psi$ be a non-trivial additive $\ell$-adic character of $k$. Let $r\geqslant 0$ and $t\geqslant 0$ be integers which are not both 0. The basic data is given by two multisets
$\text{${\chi}$}=\{\chi_{1},\ldots,\chi_{r}\},\quad\quad\text{${\varrho}$}=\{\varrho_{1},\ldots,\varrho_{t}\}$
of $\ell$-adic characters of $k^{\times}$}, which are assumed to be disjoint (no character has multiplicity $\geqslant 1$ in both of them).
Remark 9.1.
In other words, taking $\chi$} for example, this is the data for each character $\eta$ of $k^{\times}$} of a non-negative integers $n_{\eta}=n_{\eta}(\text{${\chi}$})$, representing its multiplicity in $\chi$}, distinct elements of the tuple being exactly those characters with $n_{\eta}\geqslant 1$, with the condition that
$\sum_{\chi}n_{\chi}=n,$
with obvious notation. Another interpretation would be an integer-valued non-negative measure on the set of $\ell$-adic characters of $k^{\times}$}.
The hypergeometric sums associated to the pair $(\text{${\chi}$},\text{${\varrho}$})$ are defined for $u\in{k^{\times}}$ by
$\operatorname{Hyp}(u;\text{${\chi}$},\text{${\varrho}$};\psi)=\frac{1}{|k|^{(r+t-1)/2}}\mathop{\sum\cdots\sum}\limits_{\begin{subarray}{c}x_{1},\ldots,x_{r},y_{1},\ldots,y_{t}\in k\\ \frac{x_{1}\cdots x_{r}}{y_{1}\cdots y_{t}}=u\end{subarray}}\psi\Bigl(\sum_{i=1}^{r}{x_{i}}-\sum_{j=1}^{t}y_{j}\Bigr)\prod_{i=1}^{r}\chi_{i}(x_{i})\prod_{j=1}^{t}\overline{\varrho_{j}(y_{j})}.$
Katz constructed in [ESDE, Th 8.4.2] a middle-extension sheaf $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$ on $\mathbf{G}_{m,k}$ (which is denoted $\operatorname{\mathscr{H}}_{1}(!,\psi;\text{${\chi}$},\text{${\varrho}$})$ in loc. cit.) with trace function given by these hypergeometric sums. These occur frequently in analytic number theory, and we are interested in determining which of these are gallant.
Remark 9.2.
If $t=0$, then $\text{${\varrho}$}=\emptyset$ and $\operatorname{\mathscr{H}}(\text{${\chi}$},\emptyset;\psi)$ is a Kloosterman sheaf, also denoted $\operatorname{\mathscr{K}\!\ell}(\text{${\chi}$};\psi)$. The trace function is then the generalized Kloosterman sums
$\frac{1}{|k|^{(r-1)/2}}\mathop{\sum\cdots\sum}\limits_{\begin{subarray}{c}x_{1},\ldots,x_{r}\in k\\ x_{1}\cdots x_{r}=u\end{subarray}}\psi(x_{1}+\cdots+x_{r})\prod_{i=1}^{r}\chi_{i}(x_{i}).$
Since exchanging $\chi$} and $\varrho$} amounts to applying the inversion map $x\mapsto 1/x$ and applying complex conjugation to the multisets (see [ESDE, 8.2.4]), we may also assume that $r\geqslant t$ (this relies on the fact that being gallant is invariant under such a transformation, as explained in Section 9.4). We then denote
$n=r=\max(r,t)\geqslant 1,$
which is the rank of the hypergeometric sheaf.
We recall the following definitions:
  • An $n$-element multiset777 Meaning that the sum of the multiplicities of all characters is $n$. $\chi$} is called Kummer-induced if there exists a divisor $d>1$ of $n$ such that for some (or equivalently for any) character $\eta$ of order $d$ of $k^{\times}$}, the equality $\eta\cdot\text{${\chi}$}=\text{${\chi}$}$ holds, with obvious notation for multiplication by a fixed character. We then also say that $\chi$} is $d$-Kummer-induced. By convention, $\text{${\chi}$}=\emptyset$ is $d$-Kummer-induced for every $d>1$.
  • The pair $(\text{${\chi}$},\text{${\varrho}$})$ is Kummer-induced (see [ESDE, 8.9.3]) if there exists a common divisor $d>1$ of $r$ and $t$ such that $\chi$} and $\varrho$} are both $d$-Kummer induced in the sense of the previous item.
  • The pair $(\text{${\chi}$},\text{${\varrho}$})$ is Belyi-induced (see [ESDE, § 8.10.1]) if $r=t$ and if there exists a partition $n=d+e$ of $n$ in positive integers, and characters $\alpha$, $\beta$ of $k^{\times}$} such that
    • $\beta\not=1$;
    • $\chi$} is the union of the multiset of characters $\eta$ such that $\eta^{d}=\alpha$ and the multiset of characters $\eta$ such that $\eta^{e}=\beta$ (with multiplicity);
    • $\varrho$} is the multiset of all characters $\eta$ such that $\eta^{r}=\alpha\beta$.
    Note that if $t=0$, then $(\text{${\chi}$},\text{${\varrho}$})$ cannot be Belyi-induced.
  • The pair $(\text{${\chi}$},\text{${\varrho}$})$ is primitive if it is neither Kummer-induced nor Belyi-induced.
We first consider the case when $r\not=t$.
Theorem 9.1.
Suppose that $r>t$ and $r\geqslant 2$. Let $j\colon\mathbf{G}_{m}\to\mathbf{A}^{1}$ denote the open immersion. Let
[…]
Suppose that $(\text{${\chi}$},\text{${\varrho}$})$ is primitive. Suppose that q>2(rt)+1q>2(r-t)+1 and that $r\notin\{4,8,9\}$.
Then the geometric monodromy group of $\mathscr{F}$, which is the same as that of $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$, is infinite and $\mathscr{F}$ is gallant.
Proof.
This follows from the recent work of Katz and Tiep [katz-tiep]. More precisely, let $G$ denote the geometric monodromy group of $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$.
First, the primitivity assumption on $(\text{${\chi}$},\text{${\varrho}$})$ implies that the representation of $G$ corresponding to $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$ is primitive: this follows from [KatzDMJ, Lem. 11&12] if $t=0$, and from [KRLT, Prop. 1.2] in general.
Combined with the condition q>2(rt)+1q>2(r-t)+1, this implies by [katz-tiep, Th. 2.4.4] that $G$ is infinite. Then by [katz-tiep, Th. 5.2.9], the group $G$ satisfies condition (S+); by [katz-tiep, Lemma 1.1.3, (1)] (which goes back to work of Guralnick and Tiep), this implies that the connected component of $G$ is a simple linear algebraic group acting irreducibly. ∎
The results of Katz and Tiep are in fact much more precise, both determining exactly the geometric monodromy groups in many cases, and handling small characteristic situations. However, we are essentially interested in applications where $q$ will be large (tending to infinity), so we did not attempt to summarize all known results.
Nevertheless, we discuss briefly the exceptional cases $r\in\{4,8,9\}$ since these may occur naturally. From the combination of several results of Katz [ESDE], namely Theorem 8.8.1, Theorem 8.8.2, Theorem 8.11.2 and Corollary 8.11.2.1, Theorem 8.11.3, Theorem 10.1.3, Theorem 10.7.1, Theorem 10.8.1 and Theorem 10.9.1 from loc. cit., one obtains the following:
Theorem 9.2.
Suppose that $r\in\{4,8,9\}$, that $r>t$ and that $(\text{${\chi}$},\text{${\varrho}$})$ is primitive. Let $\mathscr{F}$ be as in Theorem 9.1.
If $q$ is large enough compared with $r$, then the geometric monodromy group of $\mathscr{F}$ is infinite and is gallant unless $r\equiv t\,(\mathrm{mod}\,{2})$ and one of the following holds:
  • $r=4$, there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
    $\eta^{r-t}\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=\chi_{1/2}$
    where $\chi_{1/2}$ is the unique character of exact order $2$. In that case we have $G^{0}=\mathbf{SO}_{4}$ acting on $\mathrm{St}_{4}$ and the geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ equals and
    $G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$
  • $(r,t)=(8,2)$, $|k|\equiv 1\,(\mathrm{mod}\,{4})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ of not exact order $4$, such that
    $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/4},\chi_{3/4}\},$
    where $\chi_{1/4},\chi_{3/4}$ denote the two characters of $k^{\times}$} of exact order $4$.
    In that case one has $G^{0}=\mathbf{SL}_{2}\times\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ acting on ${\mathrm{Std}_{2}}^{\otimes 3}$.
  • $(r,t)=(9,3)$, $|k|\equiv 1\,(\mathrm{mod}\,{3})$ and there exists a character $\eta$, three characters $\chi,\varrho,\xi$ (not of order dividing $3$) satisfying $\chi.\varrho.\xi=1$, such that
    $\eta\cdot\text{${\chi}$}=\{\chi,\varrho,\xi\}\cup\{\hbox{$2$-nd roots of }\overline{\chi},\overline{\varrho},\overline{\xi}\},\ \eta\cdot\text{${\varrho}$}=\{1,\chi_{1/3},\chi_{2/3}\}$
    where $\chi_{1/3},\chi_{2/3}$ denote the two characters of $k^{\times}$} of exact order $3$.
    In that case one has $G^{0}=\mathbf{SL}_{3}\times\mathbf{SL}_{3}$ acting on ${\mathrm{Std}_{3}}^{\otimes 2}$.
Remark 9.3.
The first case above provides examples of sulfatic and oxozonic sheaves in the sense of Section 1.3.
In the $r=t$ case, the monodromy group could be finite. Under the assumption that it is infinite, the results of Katz give the following result.
Theorem 9.3.
Suppose that $r=t\geqslant 2$ and $q>r$. Assume also that $(\text{${\chi}$},\text{${\varrho}$})$ is primitive.
Let $j\colon\mathbf{G}_{m}\to\mathbf{A}^{1}$ denote the open immersion. Let
[…]
If $q$ is large enough compared with $r$ and $G$ is infinite, then $\mathscr{F}$ is gallant unless $r=4$ and there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
$\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=\chi_{1/2}$
where $\chi_{1/2}$ is the unique character of exact order $2$. In that case we have $G^{0}=\mathbf{SO}_{4}$ acting on $\mathrm{Std}_{4}$ and the geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ satisfies
$G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$
9.7. Hypergeometric sheaves with finite monodromy
To be effective, the previous result requires a criterion to determine when $G$ is finite. A fundamental result of Katz shows that this is the case if and only if the corresponding hypergeometric differential equation has finite differential Galois group; one can then apply the classification of such equations by Beukers and Heckman to determine which of these have finite geometric monodromy groups, and when this is the case, one can determine which of these are still gallant.
More precisely, let $\gamma$ be a generator of $k^{\times}$}. Using $\gamma$, we can identify the group of characters of $\mathbf{F}^{\times}_{q}$ with the subgroup $\frac{1}{|k|-1}\mathbf{Z}/\mathbf{Z}\subset\mathbf{Q}/\mathbf{Z}$, which we identify further set-theoretically with $\mathbf{Q}\cap[0,1\mathclose{[}$. In particular, the data of $\chi$} (resp. $\varrho$}) corresponds to a multiset $\text{${x}$}=\{x_{1},\ldots,x_{r}\}$ in $\mathbf{Q}/\mathbf{Z}$ (resp. a multiset $\text{${y}$}=\{y_{1},\ldots,y_{t}\}$ in $\mathbf{Q}/\mathbf{Z}$). Viewing these as multisets in $0,1\mathclose{[}$, we then have the associated hypergeometric differential equation $D_{\text{${x}$},\text{${y}$}}f=0$, with
$D_{\text{${x}$},\text{${y}$}}=\prod_{i=1}^{r}(\partial-x_{i})-x\prod_{j=1}^{t}(\partial-y_{j}),\quad\quad\partial=x\partial_{x}$
(see [ESDE, § 3.1] and [BH, § 2]). Katz [ESDE, Th. 8.17.12, Cor. 8.17.15] proves that the geometric monodromy group of $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$ is finite if and only if the differential Galois group of this differential equation is finite, and that these two groups are then isomorphic.
[…]
In addition, Beukers and Heckman [BH, Th. 7.1] have given the list of all primitive hypergeometric differential equations with $r\geqslant 3$, the case $r=2$ being a classical result of Schwarz (see the notes of Matsuda [27] for a detailed modern account). The corresponding finite groups are all finite complex reflection groups, which have been classified by Shephard and Todd. It is elementary for each of them to determine (with the help, e.g., of Magma [9]) if they are gallant. We obtain this way the following list of gallant groups, where $ST_{k}$ refers to the Shephard–Todd classification:
Group Other name $r$
$ST_{1}$ $S_{n}$ n1n-1, $n\geqslant 6$
$ST_{23}$ $W(H_{3})$ 33
$ST_{24}$ $W(J_{3}(4))$ 33
$ST_{27}$ $W(J_{3}(5))$ 33
$ST_{32}$ $W(L_{4})$ 44
$ST_{33}$ $W(K_{5})$ 55
$ST_{34}$ $W(K_{6})$ 66
$ST_{35}$ $W(E_{6})$ 66
$ST_{36}$ $W(E_{7})$ 77
$ST_{37}$ $W(E_{8})$ 88
As a concrete example, assume that $|k|\equiv 1\,(\mathrm{mod}\,{120})$. Let $\chi_{1/30}$ be a generator of the group of characters of order $30$ and $\eta_{1/8}$ a generator of the group of characters of order $8$. The group $W(E_{8})$ is the monodromy group of the hypergeometric sheaf $\operatorname{\mathscr{H}}(\text{${\chi}$},\text{${\varrho}$};\psi)$ where
$\text{${\chi}$}=\{\chi_{1/30},\chi^{7}_{1/30},\chi^{11}_{1/30},\chi^{13}_{1/30},\chi^{17}_{1/30},\chi^{19}_{1/30},\chi^{23}_{1/30},\chi^{29}_{1/30}\}$
$\text{${\varrho}$}=\{1,\eta_{1/8},\eta^{2}_{1/8},\eta^{3}_{1/8},\eta^{4}_{1/8},\eta^{5}_{1/8},\eta^{6}_{1/8},\eta^{7}_{1/8}\}$
9.8. Other examples with finite monodromy
As we mentioned in the introduction, one new feature of this paper, relying on the definition of gallant sheaves, is that we can handle some sheaves with finite monodromy group. An interesting family of examples is given by sheaves of the form
$\mathscr{F}=f_{*}\overline{\mathbf{Q}}_{\ell}/\overline{\mathbf{Q}}_{\ell},$
where $f\in k[X]$ is a non-constant polynomial, viewed as a morphism $f\colon\mathbf{A}^{1}_{k}\to\mathbf{A}^{1}_{k}$. This sheaf has trace function
$K(x)=\sum_{\begin{subarray}{c}y\in k\\ f(y)=x\end{subarray}}-1.$
The computation of the geometric monodromy group of such sheaves is a classical question, often phrased in the context of the Galois group of the polynomial equation
f(Y)X=0f(Y)-X=0
in $k(X)[Y]$. A particulary simple case arise when $f$ is a “supermorse function” (in the terminology of [ESDE, (7.10.2.2)]), which means that $\deg(f)$ is strictly less than the characteristic of $k$, that zeros of $f^{\prime}$ are simple, and that $f$ separates the zeros of $f^{\prime}$. In this case, it is known (see, e.g., [ESDE, proof of Lemma 7.10.2.3]) that the geometric monodromy group of $\mathscr{F}$ is isomorphic to the symmetric group $S_{\deg(f)-1}$ (acting by the “standard” irreducible representation of dimension $\deg(f)-1$). If $\deg(f)\geqslant 6$, the symmetric group $S_{n}$ is gallant, and hence we have here a large variety of examples where our result applies. To give a very concrete example, for an integer $d\geqslant 6$, we can take
$f=X^{d}-adX$
for some $a\in\mathbf{F}^{\times}_{q}$, where $q>d$ (see [ESDE, Th. 7.10.5]). The corresponding bilinear forms (for $b=c=1$ for simplicity) are given by the formula
$\sum_{m\sim M}\sum_{n\sim N}\alpha_{m}\beta_{n}\Bigl(\sum_{\begin{subarray}{c}y\in\mathbf{F}_{q}\\ y^{d}-ady=mn\end{subarray}}1-1\Bigr).$
It would be interesting to see concrete arithmetic applications of such bilinear forms.
9.9. Further examples
We include one further class of gallant sheaves, with infinite monodromy group, taken again from the work of Katz [ESDE, Ch. 7], and chosen because it has no obvious relation to hypergeometric sheaves. We consider again $f\in k[X]$, but instead of the sheaf of the previous section, we consider its (unitarily normalized) Fourier transform, with respect to some non-trivial additive character $\psi$ of $k$. The trace function of this Fourier transform is
$K(x)=\frac{1}{\sqrt{|k|}}\sum_{y\in k}\psi(xf(y))$
for $x\in{k^{\times}}$.
Assume that $f$ is a supermorse polynomial, and furthermore that the set $C\subset\bar{k}$ of critical values of $f$ (the set of values $f(y)$ for $y\in\bar{k}$ a root of $f^{\prime}$) is a Sidon set, i.e., the equation
$s_{1}+s_{2}=s_{3}+s_{4}$
with $s_{i}\in C$ has only solutions with $s_{1}\in\{s_{3},s_{4}\}$. Then Katz proved (combine [ESDE, Th. 7.9.6] with [ESDE, Lemma 7.10.2.3]) that, provided the characteristic of $k$ is $2\deg(f)-1$, the Fourier transform sheaf has geometric monodromy group with connected component equal to $\mathbf{SL}_{\deg(f)-1}$. Thus, this sheaf is gallant as soon as $\deg(f)\geqslant 3$.
Here also, it would be interesting to see applications of the corresponding bilinear forms, such as
$\sum_{m\sim M}\sum_{n\sim N}\alpha_{m}\beta_{n}\sum_{y\in\mathbf{F}_{q}}e\Bigl(\frac{mn(y^{d}-ady)}{q}\Bigr).$
Remark 9.4.
The long list of computations of monodromy groups in [ESDE, Ch. 7] provides a wide variety of additional examples of sheaves which are known to be gallant.
9.10. The rank one case
The very definition of a gallant sheaf implies that its rank is at least $2$. Our general method does in fact allow us to handle trace functions of rank $1$ sheaves, such as
$K(x)=\chi(f(x))\psi(g(x)),$
where $\chi$ (resp. $\psi$) is a non-trivial multiplicative (resp. additive) character of $\mathbf{F}_{q}$, and $f$, $g$ are rational functions.
Studying bilinear forms with kernel $K(m^{b}n^{c})$ reduces (with our approach) to bounding the two families of sums discussed in Proposition 5.3. These are one-variable sums of trace functions of rank one sheaves, so cancellation amounts to deciding for which value of the parameters $(\text{${r}$},\text{${s}$})$ or $(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})$ the corresponding sheaf is geometrically trivial or not.
For instance, in the case above, the first of the relevant exponential sums is
$\sum_{v\in k}\chi\Bigl(\prod_{j=1}^{m}\frac{f(s_{j}(v+r_{j})^{c})}{f(s_{j+m}(v+r_{j+m})^{c})};k\Bigr)\psi\Bigl(\sum_{j=1}^{m}{(g(s_{j}(v+r_{j})^{c})-g(s_{j+m}(v+r_{j+m})^{c}))};k\Bigr),$
for finite extensions $k$ of $\mathbf{F}_{q}$, and what is required is to determine for which $(\text{${r}$},\text{${s}$})$ or $(\text{${r}$},\text{${s}$}_{1},\text{${s}$}_{2})$, the rational fractions derived from $f(X)$ and $g(X)$ are constant or not, and especially to show that they most often are not. (Here we denote $\chi(x;k)=\chi(N_{k/\mathbf{F}_{q}}(x))$ and $\psi(x;k)=\psi(\operatorname{Tr}_{k/\mathbf{F}_{q}}(x))$.)
An example where the argument will go through is when if $g$ has at least one pole and $c\geqslant 1$, since the cancellation of the pole will only be achieved by “combinatorial” restrictions on the parameters, from which the desired estimates on the size of the diagonal sets will follow. An example is $g=1/X$, where the condition is that
$\sum_{j=1}^{m}{\Bigl(\frac{1}{s_{j}(X+r_{j})^{c}}-\frac{1}{s_{j+m}(X+r_{j+m})^{c}}\Bigr)}$
should be constant, which in turns requires that the $(r_{j+m})$ be a permutation of the $(r_{j})$, with the corresponding $s_{j}$ and $s_{j+m}$ also equal.
On the other hand, if $f=1$, $c\geqslant 1$ and $g$ is a polynomial, then there will be many more diagonal configurations. For instance, with $g=X^{2}$, $c=1$, the condition is whether the polynomial
$\sum_{j=1}^{m}{(s_{j}^{2}(X+r_{j})^{2}-s_{j+m}^{2}(X+r_{j+m})^{2})}$
is constant or not, and this will hold whenever
$\sum_{j=1}^{m}s_{j}^{2}=\sum_{j=m+1}^{2m}s_{j}^{2}$
$\sum_{j=1}^{m}r_{j}s_{j}^{2}=\sum_{j=m+1}^{2m}r_{j}s_{j}^{2},$
which defines a subvariety of codimension at most $2$, which will have many more solutions than desired for our method to work.
We leave further discussion to readers interested in specific applications.

### v3

@@
 The key achievement of the present work is to obtain estimates as good as those of these papers, but in much greater generality.
-More precisely, whereas it was previously always assumed that $K$ was of a specific type (essentially some exponential sum of specific type, such as hyper-Kloosterman sums), we are now able to succeed with qualitative conditions on $K$, and these conditions are, on the one hand, known to hold for a very wide variety of trace functions and, on the other hand, extremely robust under a variety of transformations.
+More precisely, whereas it was previously always assumed that $K$ was of a specific type (essentially some exponential sum of specific type, such as hyper-Kloosterman sums), we are now able to succeed with qualitative conditions on $K$, and these conditions are, on the one hand, known to hold for a wide variety of trace functions and, on the other hand, extremely robust under a variety of natural transformations.
 In order to state a simplified version of the main result, we recall that a trace function $K$ modulo $q$ is associated to a certain algebraic object.
@@
 After permuting $(a,b,c)$ and applying complex conjugation with $L(1/2,\overline{\chi})=\overline{L(1/2,\chi)}$, we see that is enough to consider the moments $M_{a,b,c}(q)$ and $M_{a,b,-c}(q)$ when $a$, $b$ and $c$ are positive integers, which we will always assume from now on in this section.
-It is also useful to decompose $M_{a,b,c}(q)$ along the even and odd characters, so we write
-$M_{a,b,c}(q)=M^{+1}_{a,b,c}(q)+M^{-1}_{a,b,c}(q)$
-where
-$M^{\varepsilon}_{a,b,c}(q)=\frac{1}{q-1}\sum_{\stackrel{{\scriptstyle{\scriptstyle\chi\,(\mathrm{mod}\,{q})}}}{{{\scriptstyle\chi(-1)=\varepsilon}}}}L(1/2,\chi^{a})L(1/2,\chi^{b})L(1/2,\chi^{c}),\quad\quad\varepsilon=+1,\ -1.$
 Define exponential sums modulo $q$ by
@@
 Assume that $\mathscr{K}_{a,b,\varepsilon c}$ is either gallant or oxozonic in the sense of Definition 1.2 and of Section 1.3.
-There exists $\eta>0$ and there exist polynomials $P_{a,b,c}=1$ and $P_{a,b,-c}$ such that
-$\deg(P_{a,b,-c})$ $=\delta_{a,c}=\begin{cases}1&\text{ if }a=c,\\ 0&\text{ if }a\not=c,\end{cases}$
-$P_{a,b,-c}$ $\text{ has positive leading term},$
-such that the estimates
-$M^{+1}_{a,b,\varepsilon c}(q)$ $\geqslant P_{a,b,\varepsilon c}(\log q)+O(q^{-\eta}),$
-$M_{a,a,\varepsilon c}(q)$ $=P_{a,a,\varepsilon c}(\log q)+O(q^{-\eta}),$
+There exists $\eta=\eta(a,b,c)>0$ and $D_{a,b,\varepsilon c}\geqslant 1$ such that the estimates
+$M_{a,b,\varepsilon c}(q)$ $\geqslant D_{a,b,\varepsilon c}+O(q^{-\eta}),$
+$M_{a,a,\varepsilon c}(q)$ $=D_{a,a,\varepsilon c}+O(q^{-\eta}),$
 hold for all primes $q$, where the implied constant depends only on $(a,b,c)$.
-In order for this result to be applicable, we will also prove that almost all sheaves $\mathscr{K}_{a,b,c}$ are either gallant or oxozonic, using the work of Katz.
+In order for this result to be applicable, we will also prove that almost all sheaves $\mathscr{K}_{a,b,c}$ are either gallant or oxozonic, using the works of Katz and Katz–Tiep.
 Theorem 1.8.
@@
 […]
-$|\{\chi\,(\mathrm{mod}\,{q}),\ \chi\not=1,\ L(1/2,\chi^{a})L(1/2,\chi^{b}),\ \varepsilon(\chi)\in I\}|\geqslant c(q-2),$
+$|\{\chi\,(\mathrm{mod}\,{q}),\ \chi\not=1,\ L(1/2,\chi^{a})L(1/2,\chi^{b})\not=0,\ \varepsilon(\chi)\in I\}|\geqslant c(q-2),$
 holds, where $\varepsilon(\chi)$ denotes the unitarily normalized Gauss sum of $\chi$.
@@
 3.1. Goursat’s Lemma
-We recall a version of the classical lemma of Goursat.
+We recall a version of the classical lemma of Goursat (see also the version by Serre in [SerreBook, Prop. 1.6]).
 Lemma 3.1 (Goursat’s Lemma).
@@
 There exist normal subgroups $H_{i}\subset G_{i}$ and an isomorphism $\varphi\colon G_{1}/H_{1}\to G_{2}/H_{2}$ such that
-$H_{1}\times H_{2}\subset G\subset\{(x,y)\in G_{1}\times G_{2}\,\mid\,\varphi(xH_{1})=yH_{2}\}.$
+$H_{1}\times H_{2}\subset G=\{(x,y)\in G_{1}\times G_{2}\,\mid\,\varphi(xH_{1})=yH_{2}\}.$
 We recall the proof for completeness.
 Proof.
-The kernel of the restriction of $p_{1}$ (resp.
-$p_{2}$) to $G$ is identified with a subgroup $H_{2}$ of $G_{2}$ (resp.
-a subgroup $H_{1}$ of $G_{1}$).
-One then checks that the image of the natural morphism
+[…]
 $G\to G_{1}/H_{1}\times G_{2}/H_{2}$
@@
 $\varrho_{j}([x^{\prime},y^{\prime}])=[\varrho_{j}(x^{\prime}),\varrho_{j}(y^{\prime})]=1\quad\text{ for }2\leqslant j\leqslant l,$
-(since either $\varrho_{j}(x^{\prime})$ or $\varrho_{j}(y^{\prime})$ equals $1$); this completes the induction (compare with the proof of [31, Lemma 3.3], due to Serre).
+(since either $\varrho_{j}(x^{\prime})$ or $\varrho_{j}(y^{\prime})$ equals $1$); this completes the induction (compare with the proof of [32, Lemma 3.3], due to Serre).
 The restriction $\bar{\varrho}$ of $\varrho=\varrho_{1}\otimes\cdots\otimes\varrho_{m}$ to the subgroup $K_{m}$ is isomorphic to the restriction of $\varrho_{1}\otimes 1\otimes\cdots\otimes 1$ to $K_{m}$; the space of $G$-coinvariants of $\varrho$ is contained in the space of $K_{m}$-coinvariants of $\bar{\varrho}$, and the latter is contained in the space of $N_{1}$-coinvariants of $\varrho_{1}$ (since $\varrho_{1}(K_{m})$ contains $N_{1}$), which vanishes by assumption.
@@
 Let $\mathscr{H}$ be the sheaf on $\mathbf{A}^{1}_{k}$ associated to $W$.
-Since the monodromy group of $\mathscr{H}$ is finite, this sheaf geometrically semisimple.
+Since the monodromy group of $\mathscr{H}$ is finite, this sheaf is geometrically semisimple.
 After replacing $k$ by a finite extension, if necessary, we may assume that $\mathscr{H}$ splits as a direct sum of geometrically irreducible sheaves.
@@
 We prove here, in different situations, that “most” of the sheaves which arise in sums of products of a certain type have trivial topmost cohomology group, a property which translates in bounds for the corresponding sums after applying the Riemann Hypothesis over finite fields.
-We will use repeatedly the following simple combinatorial lemma. By “graph”, we mean undirected graphs without multiple edges, but where loops are allowed.
+We will use repeatedly the following simple combinatorial lemma. By “graph”, we mean an undirected graph without multiple edges, but where loops are allowed.
 Lemma 3.10.
@@
 Since the definition of the geometric monodromy group of a sheaf only depends on the restriction of this sheaf to an arbitrary open dense subset of $\mathbf{A}^{1}_{k}$, it follows that for any open immersion $j\colon U\to\mathbf{A}^{1}_{k}$ with $U$ not empty, the sheaves $j_{!}j^{*}\mathscr{F}$ and $j_{*}j^{*}\mathscr{F}$ are gallant if and only if $\mathscr{F}$ is gallant.
+9.3. Twists
+If $\mathscr{F}$ is gallant, and if $\mathscr{L}$ is a constructible $\ell$-adic sheaf of generic rank $1$ on $\mathbf{A}^{1}$ over $k$, lisse on an open dense subset $U$, then the twisted sheaf $\mathscr{L}\otimes\mathscr{F}$, which has trace function
+$x\mapsto t_{\mathscr{L}}(x;k)t_{\mathscr{F}}(x;k)$
+is also gallant.
+This fact is immediate from the definition when $G^{0}$ is non-trivial, because the geometric monodromy group of $\mathscr{L}\otimes\mathscr{F}$ is the same as that of $\mathscr{F}$.
+If $G^{0}$ is trivial, we can argue as follows.
+Let $N$ be the core subgroup of $G$, and $H$ the preimage of $N$ in the Galois group of $k(T)$ for the homomorphism corresponding to $\mathscr{F}$.
+We then have, by restriction, a surjective homomorphism $\varrho\colon H\to N$ corresponding to $\mathscr{F}$ and a character $\chi\colon H\to\overline{\mathbf{Q}}_{\ell}^{\times}$ corresponding to $\mathscr{L}$.
+The subgroup $\varrho(\ker(\chi))$ is a normal subgroup of $N$ since $\varrho$ is surjective, so it is either central or equal to $N$, since $N$ is quasisimple (see [isaacs, Lemma 9.2]).
+If it were central, we would obtain a surjective homomorphism
+$H/\ker(\chi)\to H/\varrho^{-1}(Z(N))\stackrel{{\scriptstyle\varrho}}{{\longrightarrow}}N/Z(N),$
+which is impossible since $H/\ker(\chi)$ is abelian and $N/Z(N)$ is a non-abelian simple group.
+So we deduce that $\varrho(\ker(\chi))=N$, and this implies that the image of the Galois representation associated to $\mathscr{L}\otimes\mathscr{F}$ contains $N$, so that this sheaf is gallant.
+This means, for instance, that if we can apply our results to a function $K$ on $\mathbf{F}_{q}$, then they are also applicable to functions like
+$K(x)\chi(g(x))e(f(x)/q)$
+for any multiplicative character $\chi$ modulo $q$, and any rational functions $f$ and $g$ modulo $q$ which can be written as ratios of polynomials with degree uniformly bounded as $q$ varies (the last requirement ensuring that the complexity of the corresponding twisted sheaf remains bounded).
 9.4. Forms of exponential sums
@@
 $K^{(d)}(a)=\frac{\sin((d+1)\theta(a;q))}{\sin(\theta(a;q))}$
-for $a\in\mathbf{F}_{q}^{\times}$. It would be interesting to see if such examples have concrete analytic applications.
+for $a\in\mathbf{F}_{q}^{\times}$. See [MiInv, FMAnnals, XiInv, XiIMRN] for situations where the problem of bounding such kind of bilinear sums might be useful.
 9.7. Hypergeometric sheaves
@@
 Nevertheless, we discuss briefly the exceptional cases $r\in\{4,8,9\}$ since these may occur naturally.
-From the combination of several results of Katz [ESDE], namely Theorem 8.8.1, Theorem 8.8.2, Theorem 8.11.2 and Corollary 8.11.2.1, Theorem 8.11.3, Theorem 10.1.3, Theorem 10.7.1, Theorem 10.8.1 and Theorem 10.9.1 from loc.
+From the combination of several results of Katz [ESDE], namely Theorem 8.8.1, Theorem 8.8.2, Theorem 8.11.2 and Corollary 8.11.2.1, Theorem 8.11.3, Lemma 8.11.6, Theorem 10.8.1 and Theorem 10.9.1 from loc.
 cit., one obtains the following:
@@
 If $q$ is large enough compared with $r$, then the geometric monodromy group of $\mathscr{F}$ is infinite and is gallant unless $r\equiv t\,(\mathrm{mod}\,{2})$ and one of the following holds:
-  • $r=4$, there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
+  • $r=4$, and there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
     $\eta^{r-t}\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=\chi_{1/2}$
     where $\chi_{1/2}$ is the unique character of exact order $2$.
-In that case we have $G^{0}=\mathbf{SO}_{4}$ acting on $\mathrm{St}_{4}$ and the geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ equals and
-    $G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$
-  • $(r,t)=(8,2)$, $|k|\equiv 1\,(\mathrm{mod}\,{4})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ of not exact order $4$, such that
-    $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/4},\chi_{3/4}\},$
+In that case we have $G^{0}=\mathbf{SO}_{4}$ acting by the standard representation $\mathrm{St}_{4}$.
+The geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is given by the formula
+    $G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta\cdot\chi_{i}=1\\ \mathbf{O}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta\cdot\chi_{i}=\chi_{1/2},\end{cases}$
+    where $\chi_{1/2}$ is the unique character of order $2$.
+  • $(r,t)=(8,2)$ and one of the two holds
+    1. (1)
+      $|k|\equiv 1\,(\mathrm{mod}\,{4})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ not of exact order $4$, such that
+      (9.3) $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/4},\chi_{3/4}\},$
       where $\chi_{1/4},\chi_{3/4}$ denote the two characters of $k^{\times}$} of exact order $4$.
-    In that case one has $G^{0}=\mathbf{SL}_{2}\times\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ acting on ${\mathrm{Std}_{2}}^{\otimes 3}$.
+    2. (2)
+      $|k|\equiv 1\,(\mathrm{mod}\,{3})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ not of exact order $4$, such that
+      (9.4) $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/3},\chi_{2/3}\},$
+      where $\chi_{1/3},\chi_{1/3}$ denote the two characters of $k^{\times}$} of exact order $3$.
+    In these cases one has $G^{0}=\mathbf{SL}_{2}\times\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ acting on ${\mathrm{Std}_{2}}^{\otimes 3}$.
   • $(r,t)=(9,3)$, $|k|\equiv 1\,(\mathrm{mod}\,{3})$ and there exists a character $\eta$, three characters $\chi,\varrho,\xi$ (not of order dividing $3$) satisfying $\chi.\varrho.\xi=1$, such that
-    $\eta\cdot\text{${\chi}$}=\{\chi,\varrho,\xi\}\cup\{\hbox{$2$-nd roots of }\overline{\chi},\overline{\varrho},\overline{\xi}\},\ \eta\cdot\text{${\varrho}$}=\{1,\chi_{1/3},\chi_{2/3}\}$
+    (9.5) $\eta\cdot\text{${\chi}$}=\{\chi,\varrho,\xi\}\cup\{\hbox{$2$-nd roots of }\overline{\chi},\overline{\varrho},\overline{\xi}\},\ \eta\cdot\text{${\varrho}$}=\{1,\chi_{1/3},\chi_{2/3}\}$
     where $\chi_{1/3},\chi_{2/3}$ denote the two characters of $k^{\times}$} of exact order $3$.
     In that case one has $G^{0}=\mathbf{SL}_{3}\times\mathbf{SL}_{3}$ acting on ${\mathrm{Std}_{3}}^{\otimes 2}$.
+Proof.
+Suppose $r=4>t$. By [ESDE]*Thm 8.11.3 (1) (2), we have $G^{0}=G^{0,der}$, and the sheaf is gallant unless rtr-t is even.
+Suppose rtr-t is even.
+Then, by [ESDE]*Thm 8.11.3 (3), we have $G^{0}=\mathbf{SL}_{4},\ \mathbf{SO}_{4}$ or $\mathbf{Sp}_{4}$, depending on whether or not some $\mathscr{F}$ twisted by some $\mathscr{L}_{\eta}$ is self-dual or not, and on the type (symmetric or alternating) of the duality.
+To be precise, by [ESDE]*Thm 8.8.1, a twist $\mathscr{L}_{\eta}\otimes\mathscr{F}$ is self-dual if and only if $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$.
+If this is not the case, then $G^{0}=\mathbf{SL}_{4}$.
+Otherwise, we have
+$\eta^{r-t}\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=1\text{ or }\chi_{1/2},$
+and by [ESDE]*Thm 8.8.2, the duality is alternating in the first case (and then $G^{0}=\mathbf{Sp}_{4}$) and symmetric in the second case (and then $G^{0}=\mathbf{SO}_{4}$).
+Suppose we are in the symmetric case.
+The determinant of $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is given by [ESDE]*Lemma 8.11.6; since $r-t\geqslant 2$, it is equal to $\mathscr{L}_{\Lambda_{\eta}}$ where
+$\Lambda_{\eta}=\eta^{4}\prod_{i=1}^{4}\chi_{i}.$
+This character is either $1$ or $\chi_{1/2}$, and the geometric monodromy group of $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is correspondingly either $\mathbf{SO}_{4}$ or $\mathbf{O}_{4}$.
+If $r=8$ or $9$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) $G^{0}=G^{0,der}$ and the sheaf is gallant unless rt=6r-t=6.
+If $(r,t)=(8,2)$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) and [ESDE]*Theorem 10.8.1 the sheaf is gallant unless $(\eta.\text{${\chi}$},\eta.\text{${\varrho}$})$ is of the shape (9.3) or (9.4).
+If $(r,t)=(9,3)$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) and [ESDE]*Theorem 10.9.1 the sheaf is gallant unless $(\eta.\text{${\chi}$},\eta.\text{${\varrho}$})$ is of the shape (9.5). ∎
 Remark 9.3.
 The first case above provides examples of sulfatic and oxozonic sheaves in the sense of Section 1.3.
-In the $r=t$ case, the monodromy group could be finite. Under the assumption that it is infinite, the results of Katz give the following result.
+In the $r=t$ case, the monodromy group could be finite. Under the assumption that it is infinite, the work of Katz provides the following:
 Theorem 9.3.
@@
 If $q$ is large enough compared with $r$ and $G$ is infinite, then $\mathscr{F}$ is gallant unless $r=4$ and there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
-$\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=\chi_{1/2}$
+$\prod_{i=1}^{4}\chi_{i}\times\prod_{j=1}^{4}\varrho_{j}^{-1}=\chi_{1/2}$
 where $\chi_{1/2}$ is the unique character of exact order $2$.
 In that case we have $G^{0}=\mathbf{SO}_{4}$ acting on $\mathrm{Std}_{4}$ and the geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ satisfies
-$G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$
+$G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta.\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta.\chi_{i}=\chi_{1/2}.\end{cases}$
 9.8. Hypergeometric sheaves with finite monodromy